### v1

# Adaptive Local Combining with Decentralized Decoding for Distributed Massive MIMO

Abstract
A major bottleneck in uplink distributed massive multiple-input multiple-output networks is the sub-optimal performance of local combining schemes, coupled with high fronthaul load and computational cost inherent in centralized large scale fading decoding (LSFD) architectures. This paper introduces a decentralized decoding architecture that fundamentally breaks from the conventional LSFD, by allowing each AP calculates interference-suppressing local weights independently and applies them to its data estimates before transmission. Furthermore, two generalized local zero-forcing (ZF) framework, generalized partial full-pilot ZF (G-PFZF) and generalized protected weak PFZF (G-PWPFZF), are introduced, where each access point (AP) adaptively and independently determines its combining strategy through a local sum spectral efficiency optimization that classifies user equipments (UEs) as strong or weak using only local information, eliminating the fixed thresholds used in PFZF and PWPFZF. To further enhance scalability, pilot-dependent combining vectors instead of user-dependent ones are introduced and are shared among users with the same pilot. The corresponding closed-form spectral efficiency expressions are derived. Numerical results show that the proposed generalized schemes consistently outperform fixed-threshold counterparts, while the introduction of local weights yields lower overhead and computation costs with minimal performance penalty compared to them.
I Introduction
Distributed massive MIMO (D-mMIMO) networks are a promising framework for next-generation wireless systems, as they can enhance weak users’ signal and guarantee consistent service quality across wide coverage areas by enabling numerous distributed access points (APs) to coherently serve user equipments (UEs) [1, 2]. Unlike conventional cellular architectures, such distributed systems enhance coverage and reliability through cooperative processing and coherent transmission/reception [3]. To fully use the advantages of D-mMIMO, while ensuring scalability, distributed signal processing at the APs is essential [4]. A major challenge in this regard is the development of effective uplink combining schemes that can alleviate inter-user interference within practical limitations.
Centralized zero-forcing (ZF) combining, as studied in [5, 6, 7], requires substantial fronthaul signaling and imposes prohibitively high computational costs. Similarly, centralized minimum mean square error (MMSE) combining [8] suffers from both high computation cost and fronthaul overhead, which makes it impractical for large-scale deployments. To alleviate this, partial MMSE combining is proposed in [9, 3, 10], where the interference from only a subset of UEs is considered when designing the combining vector for each UE. Along similar lines, partial centralized approaches have been explored in [11, 12], where joint ZF is applied across a subset of APs based on stronger channels. These partial centralized ZF schemes reduce the fronthaul load compared to fully centralized ZF, but the overhead still scales with the number of UEs, as for each UE, a subset of APs need to send the channel estimates to the CPU, which limits their suitability for ultra-dense network scenarios. These centralized and partial centralized combining schemes follow the the fully centralized architecture as shown in Fig. 1, which has very high computational cost and fronthaul overhead, as for all UEs, channel need to be estimated at the CPU with the cooperation of all or a subset of APs.
Several distributed combining approaches have been proposed in the literature. Maximum ratio (MR) combining, introduced in [1], is computationally simple and fronthaul-efficient, but its performance is fundamentally limited by its inability to effectively mitigate inter-user interference. To improve upon MR, a local partial MMSE scheme is proposed in [3, 10], which achieves superior interference suppression and performance. However, a key drawback of this approach is that closed-form expressions for the spectral efficiency (SE) cannot be derived. As a result, resource allocation and system optimization must rely on computationally expensive Monte Carlo simulations, and the lack of tractable expressions makes theoretical performance analysis and gaining insights into the system behavior difficult to obtain. In [1, 3], local combining schemes follow an architecture of simple decoding as shown in Fig. 1, thus reducing the overhead apart from local data estimates, yet it suffers from sub-optimality.
(a) FullyCentralized(b) Local Combining and LSFD(c) Local Combining and Simple decoding (d) Proposed Decentralized decoding APs CPU Received SignalsChannel Estimates Global Data Estimates $\hat{x}_{t}$ APs Local Combining $\mathbf{v}_{mt}^{H}\mathbf{y}_{m}\rightarrow\hat{x}_{mt}$ CPU Local Data Estimates $\hat{x}_{mt}$Local Channel Estimates Calculates LSFD Weights $\mathbf{a}_{t}$ Final Estimates $\hat{x}_{t}=\mathbf{a}_{t}^{H}\hat{\mathbf{x}}_{t}$ APs Local Combining $\mathbf{v}_{mt}^{H}\mathbf{y}_{m}\rightarrow\hat{x}_{mt}$ CPU Local Data Estimates $\hat{x}_{mt}$ Final Estimates $\hat{x}_{t}=\sum_{m}\hat{x}_{mt}$ APs Local Combining $\mathbf{v}_{mt}^{H}\mathbf{y}_{m}\rightarrow\hat{x}_{mt}$ Local Weighting $a_{mt}\cdot\hat{x}_{mt}$ CPU Pre-weighted DataEstimates $a_{mt}\hat{x}_{mt}$ Final Data Estimates $\hat{x}_{t}=\sum_{m}a_{mt}\hat{x}_{mt}$
Figure 1: Comparison of uplink processing architectures in D-mMIMO: (a) Fully centralized processing is computationally prohibitive. (b) Local combining and LSFD requires sharing soft estimates and global CSI to compute optimal weights at the CPU. (c) Local combining and Non-coherent decoding is simple but sub-optimal. (d) The proposed decentralized decoding architecture: each AP performs local combining and weighting independently, reducing the CPU to a simple aggregator and eliminating fronthaul overhead for coordination.
The local combining schemes such as full-pilot ZF (FZF), partial FZF (PFZF), and protected weak PFZF (PWPFZF) are proposed in [13, 14], offering tractable closed-form spectral efficiency (SE) expressions. These techniques construct combining vectors locally at each AP using channel statistics to suppress interference within designated user groups. In traditional FZF, PFZF, and PWPFZF, each AP generates a single combining vector per orthogonal pilot sequence and assigns scaled versions of this vector to individual UEs based on their estimated channel variances [13]. The fundamental limitations of PFZF and PWPFZF stem from two factors. First, the rigid, threshold-based user grouping may misclassify UEs, leading to suboptimal interference suppression and inefficient use of spatial resources. Second, all APs are constrained to follow the same combining scheme, which prevents each AP from independently adapting its strategy to local channel conditions and interference patterns, further limiting overall system performance.
The large-scale fading decoding (LSFD) as shown in Fig. 1, is employed at the central processing unit (CPU) to enhance uplink performance in D-mMIMO networks [8, 15, 13, 14]. In the optimal LSFD (o-LSFD) approach [13], each AP must transmit both channel estimates and data estimates for all served UEs to the CPU, effectively doubling the fronthaul load while imposing computational costs that scale with the number of users [16]. To alleviate this burden, partial LSFD (p-LSFD) is proposed in [15], which reduces computational cost and scalability requirements by considering only partial interference in the combining process. However, p-LSFD still requires matrix inversions of the same dimension as o-LSFD and the transmission of channel estimates to the CPU. As network size and UE density increase, these operations become increasingly demanding with respect to computation and overhead, significantly limiting the scalability and practicality of LSFD-based architectures in ultra-dense D-mMIMO deployments.
I-A Contributions
To overcome the above mentioned issues, this work introduces a novel decentralized decoding framework for D-mMIMO that fundamentally enhances scalability, adaptability, and analytical tractability. The principal contributions are summarized as follows:
  • Decentralized Decoding Architecture with Local Weighting: We propose a novel uplink processing architecture as shown in Fig. 1, that completely eliminates the need for centralized LSFD and its associated fronthaul overhead. Central to this architecture is the design of interference-suppressing local weights computed independently by each AP using only its local channel information. These weights are applied to data estimates before transmission, enabling the central processor to operate as a simple aggregator rather than a coordinator. This eliminates the need for any inter-AP coordination or channel information sharing with the CPU, reducing it to a low-complexity aggregator. This transition from a coordinated to a truly distributed architecture is the key enabler for the practical, large-scale deployment of cell-free networks.
  • Adaptive Combining Strategy: We introduce generalized PFZF (G-PFZF) and generalized PWPFZF (G-PWPFZF). Unlike conventional PFZF and PWPFZF, where all APs must employ the same scheme, the generalized framework allows each AP to independently adapt its combining strategy. Specifically, under G-PFZF an AP may locally switch among PFZF, FZF, or MR; under G-PWPFZF, it may switch among PWPFZF, FZF, or MR. Thereby eliminating rigid threshold-based grouping and unlocking per-AP adaptability for improved scalability and performance.
  • Novel Distributed Local Optimization for Pilot Partitioning: We also propose novel distributed optimization framework where each AP independently partitions pilots (and consequently, their associated UEs) into strong or weak groups based on a local sum spectral efficiency metric. This pilot-level optimization is solved using projected gradient ascent optimization and therefore maintaining low computational cost while maintaining functional equivalence with user-level partitioning.
  • Closed-Form Spectral Efficiency Expressions: We derive closed-form expressions for the achievable spectral efficiency of the proposed system. This analysis is essential due to the new statistical properties introduced by pilot-dependent processing and local weighting, and it provides critical insights for system design without the need for Monte Carlo simulation as in MMSE based schemes.
  • Performance Analysis: Extensive numerical results demonstrate that our decentralized framework and adaptive grouping strategy achieve good performance with reduced computation cost with state-of-art baselines. The adaptive UE partitioning outperforms traditional fixed-threshold schemes, while the overall design significantly reduces fronthaul and computational costs in comparison to LSFD framewrok and fixed threshold partitioning, validating the approach as a practical solution for large-scale deployment.
Organization: The remainder of the paper is organized as follows. Section II describes the system model, along with the channel estimation and data detection procedures. Section III presents the proposed combining schemes and derives their corresponding closed-form SE expressions. Section IV provides numerical simulations to evaluate performance. Finally, Section V concludes the paper and outlines potential directions for future work.
Notation: Scalars are denoted in italics (e.g., $x$), vectors by bold lowercase letters (e.g., $\mathbf{x}$), and matrices by bold uppercase letters (e.g., $\mathbf{X}$). The transpose and Hermitian transpose are denoted by $(\cdot)^{T}$ and $(\cdot)^{H}$, respectively, while the complex conjugate of a scalar is written as $(\cdot)^{*}$. The sets of real and complex numbers are denoted by $\mathbb{R}$ and $\mathbb{C}$. The cardinality of a set $\mathcal{S}$ is denoted by $|\mathcal{S}|$. The expectation operator is denoted by $\mathbb{E}[\cdot]$. The identity matrix of size $N$ is denoted by $\mathbf{I}_{N}$. A complex Gaussian random vector $\mathbf{x}$ with mean $\bm{\mu}$ and covariance matrix $\mathbf{K}$ is denoted as $\mathbf{x}\sim\mathcal{CN}(\bm{\mu},\mathbf{K})$. Also, $_{l}$ represent the $l$-th element of the vector $[.]$.
II System Model
The system consist of $T$ UEs with a single antenna and $M$ APs, where each AP has $A$ antennas, ensuring that $T\ll MA$. Within the coverage area of interest, there is a uniform distribution of both APs and UEs, and APs work together to service the UEs using the same frequency and time resources. Every AP is linked to the CPU by fronthaul link, so that the network can operate together and process signals. Using uplink pilot transmissions, the system estimates the AP-UE channels under the time-division duplexing (TDD) regime. The wireless channel is represented using a block fading model, characterized by a coherence block of length $L_{c}$ symbols of which $L_{p}$ pilot symbols are designated for uplink pilot training. The small-scale fading is represented by a Rayleigh fading vector $\textbf{h}_{mt}\in\mathbb{C}^{A\times 1}$, and the large-scale fading coefficient (LSFC), encompassing both path-loss and shadowing effects, is denoted by $\beta_{mt}$. The comprehensive channel vector $\textbf{g}_{mt}\in\mathbb{C}^{A\times 1}$ between the AP $m$ and the UE $t$, encompassing both small- and large-scale fading, is defined as $\mathbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}$. We suppose that $\textbf{h}_{mt}\sim\mathcal{N}_{\mathbb{C}}(0,\textbf{I}_{A})$, for all $m$ and $t$, is an independent and identically distributed complex Gaussian random vector, thus $\mathbf{g}_{mt}\sim\mathcal{N}_{\mathbb{C}}(0,\beta_{mt}\textbf{I}_{A})$.
II-A Channel Estimation
A pilot sequence $\sqrt{L_{p}}\bm{\psi}_{i_{t}}\in\mathbb{C}^{L_{p}\times 1}$ is transmitted by each UE $t$ during the uplink training stage, with $|\bm{\psi}_{i_{t}}|^{2}=1$, where $i_{t}$ is the pilot index of the UE $t$. At the $m$-th AP, the received signal $\textbf{y}^{\text{pilot}}_{m}\in\mathbb{C}^{A\times L_{p}}$ is provided by:
$\textbf{y}^{pilot}_{m}=\sum_{t=1}^{T}\sqrt{p^{p}_{t}L_{p}}\textbf{g}_{mt}\bm{\psi}_{i_{t}}^{{H}}+\textbf{N}_{m},$
where $\textbf{N}_{m}\in\mathbb{C}^{A\times L_{p}}$ represents the additive white Gaussian noise (AWGN) matrix characterized by independent and identically distributed complex Gaussian entries. Additionally, $p^{p}_{t}$ represents the normalized signal-to-noise power ratio for the UE $t$ during pilot transmission. The minimum mean square error (MMSE) estimate $\hat{\mathbf{g}}_{mt}\in\mathbb{C}^{A\times 1}$ of the true channel vector $\mathbf{g}_{mt}$ is provided by [1]:
$\hat{\textbf{g}}_{mt}=c_{mt}\textbf{y}^{pilot}_{m}\bm{\psi}_{i_{t}},$
where
${c}_{mt}=\frac{\sqrt{p^{p}_{t}L_{p}}\beta_{mt}}{\sum_{k=1}^{T}p^{p}_{k}L_{p}\beta_{mk}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+1}.$
The estimate $\hat{\mathbf{g}}_{mt}$ and estimated error $\tilde{\mathbf{g}}_{mt}={\mathbf{g}}_{mt}-\hat{\mathbf{g}}_{mt}$ are independent Gaussian with distributions $\hat{\mathbf{g}}_{mt}\sim\mathcal{N}_{\mathbb{C}}(0,\gamma_{mt}\textbf{I}_{A})$ and $\tilde{\mathbf{g}}_{mt}\sim\mathcal{N}_{\mathbb{C}}(0,(\beta_{mt}-\gamma_{mt})\textbf{I}_{A})$, where
$\gamma_{mt}=\mathbb{E}\big\{|[\hat{\mathbf{g}}_{mt}]_{l}|^{2}\big\}=\left(\frac{p^{p}_{t}L_{p}\beta_{mt}^{2}}{\sum_{k=1}^{T}p^{p}_{k}L_{p}\beta_{mk}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+1}\right),$ (1)
The matrix of estimated channels at the AP $m$ is denoted by $\mathbf{\hat{G}}_{m}=[\hat{\mathbf{g}}_{m1},\hat{\mathbf{g}}_{m2},\ldots,\hat{\mathbf{g}}_{mT}]\in\mathbb{C}^{A\times T}$. Owing to pilot reuse among the UEs, the estimated channel vectors $\hat{\mathbf{g}}_{mt}$ and $\hat{\mathbf{g}}_{mk}$ corresponding to two UEs $t$ and $k$ that share the same pilot sequence become linearly dependent. As a result, the columns of $\mathbf{\hat{G}}_{m}$ exhibit linear dependence, rendering the matrix $\mathbf{\hat{G}}_{m}$ rank-deficient. The full rank matrix is designed by removing the linear dependent columns and by including only those columns that contain linearly independent channel estimates, representing one vector for each pilot sequence. Thus, the full rank matrix $\mathbf{\bar{G}}_{m}\in\mathbb{C}^{A\times L_{p}}$ is designed as:
$\mathbf{\bar{G}}_{m}=\textbf{y}^{pilot}_{m}\bm{\Psi},$
where $\bm{\Psi}=[\bm{\psi}_{1},\bm{\psi}_{2},\ldots,\bm{\psi}_{L_{p}}]\in\mathbb{C}^{L_{p}\times L_{p}}$. The channel estimate $\hat{\mathbf{g}}_{mt}$ is rewritten as:
$\hat{\mathbf{g}}_{mt}=c_{mt}\mathbf{\bar{G}}_{m}\mathbf{e}_{i_{t}},$
where $\mathbf{e}_{i_{t}}$ is the $i_{t}$-th column of identity matrix $\mathbf{I}_{L_{p}}$.
II-B Uplink Transmission and Spectral Efficiency
Let $x_{t}$ denote the unit power uplink data signal transmitted by the UE $t$, satisfying $\mathbb{E}{|x_{t}|^{2}}=1$. The uplink normalised signal-to-noise power ratio corresponding to the UE $t$ and additive white Gaussian noise are denoted by $p^{u}_{t}$ and $\textbf{n}_{m}\sim\mathcal{N}_{\mathbb{C}}(0,\textbf{I}_{A})$, respectively. The uplink signal $\textbf{y}^{u}_{m}\in\mathbb{C}^{A\times 1}$ received at the $m$-th AP is expressed as:
$\textbf{y}^{u}_{m}=\sum_{t\in\mathcal{T}}\textbf{g}_{mt}\sqrt{p^{u}_{t}}x_{t}+\textbf{n}_{m}.$
In decentralized data detection, unlike centralized approach where each AP sends received signal and channel estimates to CPU, each AP performs data detection for its served UEs using a local linear received combining vector. Following the proposed decentralized decoding architecture, each AP computes its own interference-suppressing local weight based on the local signal-to-interference-plus-noise ratio (SINR). These weights are applied to the locally detected data symbols before being sent to the CPU. This differs from LSFD, where local data estimates and channel estimates are transmitted to the CPU for centralized processing.
The locally weighted data symbol $\hat{x}_{mt}$ for the UE $t$ at the AP $m$ is given by
$\hat{x}_{mt}=a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{y}^{u}_{m},$ (2)
where $\mathbf{v}_{m{i_{t}}}\in\mathbb{C}^{A\times 1}$ denotes the local combining vector for the UE $t$ at the AP $m$, and $a_{mt}$ is the corresponding local weight. These weighted local estimates are forwarded to the CPU, which aggregates them to obtain the final estimate of $x_{t}$:
$\hat{x}_{t}=\sum_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{y}^{u}_{m}.$ (3)
After expanding $\textbf{y}^{u}_{m}$, the (3) can also be rewritten as
$\hat{x}_{t}=\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}+\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\tilde{\textbf{g}}_{mt}+\sum_{k=1,k\neq t}^{T}\sum_{m=1}^{M}\sqrt{p^{u}_{k}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk}x_{k}+\sum_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m}.$ (4)
Here, the first term is the desired signal for the UE $t$, the second term is interference term due to imperfect channel estimation, the third term is interference term due to all other UEs and last term the noise. Based on the decomposition in (4), the achievable uplink spectral efficiency is obtained using the bounding technique and the Shannon capacity lower bound [17, 18, 16] is given by Theorem 1.
Theorem 1.
[16, 13] : A lower bound on the uplink ergodic capacity for UE $t$ is given by
$\textsc{SE}^{u}_{t}=L_{u}\log_{2}\big(1+\textsc{SINR}_{t}\big),$ (5)
where $L_{u}$ is $\Big(\frac{1-\frac{L_{p}}{L_{c}}}{2}\Big)$ and the SINR for the UE $t$ is defined by:
[…]
where,
$\begin{split}&\textsc{DS}_{t}=\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbb{E}\left\{\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}\right\}\\ &\textsc{BU}_{t}=\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}(\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mt}-\mathbb{E}\left\{\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mt}\right\})\\ &\textsc{UI}_{tk}=\textsc{PC}_{tk}=\sum_{m=1}^{M}\sqrt{p^{u}_{k}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk}\\ &\textsc{GN}_{t}=\sum_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m},\end{split}$
where $\mathcal{P}_{i_{t}}$ represents the subset of UEs sharing the pilot $i_{t}$. Also, the SE lower bound, in (5), is valid regardless of the combining scheme used.
III Adaptive Combining and Local Weights Analysis
In this section, we analyze the proposed generalized combining schemes, G-PFZF and G-PWPFZF, from both performance and computational cost perspectives. We first derive closed-form expressions for the uplink SE and local combining weights for both schemes. These expressions provide analytical insights into how the proposed framework improves performance while maintain scalability. Furthermore, we evaluate the computational and fronthauling costs associated with each scheme, highlighting the significant reductions achieved through decentralized decoding system design.
In [13], the combining vectors for any UE $t$ is dependent on $c_{mt}$, which is unique to each UE. As a result, generating a distinct ZF vector for every UE entails a number of multiplications that scales linearly with the total number of UEs. To reduce computational cost, we seek an alternative formulation in which the number of multiplications scales with the number of orthogonal pilot sequences, rather than with the number of UEs.
III-A Generalized Partial Full-Pilot Zero Forcing Combining Scheme
In the proposed G-PFZF combining scheme, each AP independently designs a combining vector for each of its pilots and pilot-sharing UEs use the same pilot, as APs are not able to distinguish among individual UEs sharing the same pilot. The design of combining vectors depends on whether a pilot is strong or weak. Thus, each AP categorizes the pilots into two groups, strong and weak based on the local sum SE optimization. Unlike traditional PFZF, G-PFZF does not forces the APs to select atleast one pilot as strong and allows the APs to choose the combining schemes among PFZF or MR or FZF.
For the AP $m$, a set of strong pilots is denoted by $\mathcal{S}_{m}$ and a complementary set of weak UEs as $\mathcal{W}_{m}$. Let $L_{\mathcal{S}_{m}}$ denote the number of distinct strong pilots used in the set $\mathcal{S}_{m}$, and define the corresponding set of pilot indices as $\mathcal{R}_{\mathcal{S}_{m}}=\{r_{m,1},r_{m,2},\dots,r_{m,L_{\mathcal{S}_{m}}}\}$. The matrix $\mathbf{\bar{G}}_{m}\in\mathbb{C}^{A\times L_{p}}$ contains the estimated channel vectors corresponding to all $L_{p}$ orthogonal pilot sequences, whereas the strong pilots utilize only $L_{\mathcal{S}_{m}}$ of them. To isolate the relevant columns of $\mathbf{\bar{G}}_{m}$, we define the selection matrix $\mathbf{E}_{\mathcal{S}_{m}}=[\mathbf{e}_{r_{m,1}},\dots,\mathbf{e}_{r_{m,L_{\mathcal{S}_{m}}}}]\in\mathbb{C}^{L_{p}\times L_{\mathcal{S}_{m}}}$, where $\mathbf{e}_{r}$ denotes the $r$-th column of the identity matrix $\mathbf{I}_{L_{p}}$. The resulting product $\mathbf{\bar{G}}_{m}\mathbf{E}_{\mathcal{S}_{m}}$ yields a full-rank matrix containing only the $L_{\mathcal{S}_{m}}$ linearly independent columns corresponding to the strong pilots. For a given pilot $i\in\mathcal{S}_{m}$, let $j_{mi}\in\{1,2,\dots,L_{\mathcal{S}_{m}}\}$ denote the index of the pilot $i$ within $\mathcal{R}_{\mathcal{S}_{m}}$, and define the vector $\bm{\varepsilon}_{j_{mi}}\in\mathbb{C}^{L_{\mathcal{S}_{m}}}$ as the $j_{mi}$-th column of the identity matrix $\mathbf{I}_{L_{\mathcal{S}_{m}}}$, such that $\mathbf{E}_{\mathcal{S}_{m}}\bm{\varepsilon}_{j_{mi}}=\mathbf{e}_{i}$.
As discussed earlier, the effective channel matrix corresponding to the strong-pilot UEs is represented as $\mathbf{\bar{G}}_{m}\mathbf{E}_{\mathcal{S}_{m}}$. Following the zero-forcing principle, the pseudo-inverse of this matrix can be expressed as $(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}})^{-1}$. To obtain the combining vector associated with a specific pilot, this pseudo-inverse term is multiplied by $\bm{\varepsilon}_{j_{mi}}$ and subsequently normalized. Thus, the local ZF combining vector for pilot $i\in\mathcal{S}_{m}$ at AP $m$ is given by:
$\mathbf{v}_{mi}^{\text{LZF}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\bm{\varepsilon}_{j_{mi}}}{\sqrt{\mathbb{E}\Big\{\Big\|\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\bm{\varepsilon}_{j_{mi}}\Big\|^{2}\Big\}}},$ (7)
Similarly, the MR combining vector for pilot $i\in\mathcal{W}_{m}$ at AP $m$ corresponds directly to the channel vector associated with that pilot and is expressed as:
$\mathbf{v}_{mi}^{\text{MR}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{e}_{i}}{\sqrt{\mathbb{E}\Big\{\Big\|\mathbf{\bar{G}}_{m}\mathbf{e}_{i}\Big\|^{2}\Big\}}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{e}_{i}}{\sqrt{A\theta_{mi}}},$ (8)
where $\theta_{mi}=\mathbb{E}\big\{|[\mathbf{\bar{G}}_{m}\mathbf{e}_{i}]_{l}|^{2}\big\}=\big(\sum_{k=1}^{T}p^{p}_{k}L_{p}\beta_{mk}\left|\bm{\psi}^{H}_{i}\bm{\psi}_{i_{k}}\right|^{2}+1\big)$. Following [19, 20, 21], the normalization term in (7), for $L_{\mathcal{S}_{m}}\times L_{\mathcal{S}_{m}}$ complex Wishart matrix, $\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}$, with $A$ degree of freedom satisfying $A-1\geq L_{\mathcal{S}_{m}}$, can be rewritten as
$\sqrt{\mathbb{E}\Big\{\bm{\varepsilon}^{H}_{j_{mi}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\bm{\varepsilon}_{j_{mi}}\Big\}}=\sqrt{\mathbb{E}\Big\{\big[(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\big]_{j_{mi},j_{mi}}\Big\}}=\frac{1}{\sqrt{(A-L_{\mathcal{S}_{m}})\theta_{mi}}},$ (9)
Also, the G-PFZF scheme suppresses the interference among the UEs whose pilot lies in $\mathcal{S}_{m}$ by sacrificing the $L_{\mathcal{S}_{m}}$ degree of freedom from the total $A$ degrees to boost the desired signal. Therefore, for any UE $t$ such that $i_{t}\in\mathcal{S}_{m}$
$(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}=\begin{cases}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t},\end{cases}$ (10)
$\begin{split}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}})\gamma_{mk}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},&\text{if }i_{k}\notin\mathcal{S}_{m}\end{cases}\end{split}$ (11)
Similarly, for any UE $t$ such that $i_{t}\in\mathcal{W}_{m}$
$\mathbb{E}\{(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{g}_{mk}}\}=\begin{cases}\sqrt{A\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}.\end{cases}$ (12)
$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\mathbb{E}\Big\{\Big|(\frac{1}{\sqrt{A\theta_{mi_{t}}}c_{mt}}\hat{\mathbf{g}}^{H}_{mt}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A+1)\gamma_{mk}&\text{if }i_{k}=i_{t},\\ \gamma_{mk},&\text{if }i_{k}\neq i_{t},\end{cases}$ (13)
Note 1: From (11) and (13), it is evident that local ZF achieves superior interference suppression compared to MR, since each AP can eliminate interference among UEs with pilots grouped as strong. However, this interference mitigation comes at the cost of $L_{\mathcal{S}_{m}}$ degrees of freedom, which reduces the array gain available for desired signal enhancement.
By using the combining vectors in (7) and (8), the closed-form expression of achievable SE for G-PFZF combining can be computed using Theorem 1 with the corresponding SINR given by:
[…]
where $\delta_{mi_{t}}$ is defined as:
$\delta_{mi_{t}}=\begin{cases}1,&\text{if }\ i_{t}\in\mathcal{S}_{m},\\ 0,&\text{otherwise}.\end{cases}$ (15)
The derivation of SINR expression is given in Appendix -A.
Local Weight Design for G-PFZF Scheme: A fundamental limitation of the conventional LSFD architecture is its reliance on centralized optimization [13]. The optimal LSFD weights are calculated at the CPU using knowledge of global channel statistics, which creates a significant fronthaul signaling overhead and computational burden, thus undermining the scalability of the network. In this work, we break this dependency by introducing a novel decentralized decoding architecture where each AP $m$ independently computes a local interference-suppressing weight $a_{mt}$ for each UE $t$, based solely on its local channel information. Since these weights are derived from the LSFCs, they evolve much slower than the small-scale fading components, ensuring long-term stability and minimal operational overhead.
The primary objectives of this local weight design are twofold: (1) to maximize the contribution of the local estimate to the final SINR at the CPU, and (2) to actively mitigate the dominant interference components, specifically, pilot contamination and inter-user interference, at their source, prior to transmission. This approach eliminates the need for the APs to send the channel estimates to the CPU and the CPU to perform any complex optimization, reducing its role to that of a simple aggregator that sums the pre-weighted estimates from all APs.
The main idea behind designing decoding weights is that it should amplify the data signal received from AP with strong local desired signal component and suppress those dominated by interference. In LSFD design, weights are computed through coordination among the APs, specifically the coherent interference terms. In contrast, the proposed decentralized framework eliminates inter-AP coordination, requiring each AP to determine its weight independently. Accordingly, an AP with a higher local SINR is given a larger weight than one with a lower SINR in order to enhance the global SINR. For simplicity, the local weight is chosen to be equal to the local SINR.
The closed-form expression of the local SINR for the UE $t$ at the AP $m$ can be obtained from (2) by expanding it in the same manner as in (4) and following the derivation procedure outlined in Appendix -A. Accordingly, the closed-form of the local SINR or local weight with respect to the AP $m$ of the UE $t$ is given by:
$a_{mt}=\text{SINR}^{\text{G-PFZF}}_{mt}=\frac{p^{{u}}_{t}(b_{tt}^{m})^{2}}{\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{{u}}_{k}(b_{kt}^{m})^{2}+W_{mt}}$ (16)
where
$b_{kt}^{m}=\sqrt{(A-\delta_{{m}i_{t}}L_{\mathcal{S}_{m}})\gamma_{{m}k}},$
${W}_{{m}t}=\sum\limits_{k=1}^{T}p^{{u}}_{k}(\beta_{{m}k}-\delta_{{m}i_{t}}\delta_{{m}i_{k}}\gamma_{{m}k})+1.$
The numerator represents the desired signal strength for the desired UE $t$ at the AP $m$, capturing the gain achieved through the chosen combining strategy. The denominator is designed to penalize the weight by two main sources of impairment: the coherent interference from pilot contamination and the combined power of non-coherent interference and noise.
By applying these weight to its local estimate, each AP independently produces a ”soft” decision that is already pre-optimized for local interference suppression. This design ensures that the signals forwarded to the CPU require no further large-scale fading decoding, thereby completely eliminating the associated fronthaul overhead for channel statistics and the computational cost of centralized weight calculation. The result is a radically simplified and highly scalable uplink processing architecture that retains the performance benefits of distributed APs design.
Adaptive Pilot Grouping for G-PFZF Combining: This subsection details the local optimization framework that each AP employs to independently partition its set of pilots into strong and weak groups. This classification dictates the appropriate combining strategy, G-PFZF for strong pilots or MR for weak ones, and crucially, generalizes to pure FZF or pure MR should the optimization deem them to be optimal.
The foundation of this optimization is the local SINR expression given in (16). As $L_{\mathcal{S}m}=\sum_{i=1}^{L_{p}}\delta_{mi}$ and substituting $L_{\mathcal{S}m}$ into (16) makes the dependency explicit:
[…]
This expression reveals a fundamental trade-off governed by the size of the strong pilot set, $L_{\mathcal{S}m}$. Allocating a pilot to the strong group (i.e., setting $\delta_{mi}=1$) employs ZF to suppress the non-coherent interference from that pilot’s UEs as well as other strong pilots UEs. However, this comes at the cost of a reduced array gain, $(A-L_{\mathcal{S}_{m}})$, for all UEs served by the AP, which diminishes the desired signal power. Conversely, classifying a pilot as weak preserves the full array gain but forgoes any interference suppression for its UEs. This inherent trade-off between maximizing signal strength and minimizing interference makes the grouping problem non-trivial and renders a fixed, network-wide threshold highly suboptimal.
To navigate the above mentioned trade-off, the AP $m$ independently solves a local optimization problem. The objective is to select the binary assignment variables $\delta_{mi}\in\{0,1\}$ for all pilots that maximize the sum of the local SEs for all UEs, as defined by
$\underset{\bm{\delta}_{m}}{\max~}$ $\ \sum\limits_{t=1}^{T}\log_{2}\left(1+\frac{S_{mt}}{I_{mt}}\right),$ (18a)
subject to: $\ \delta_{mi}\in[0,1],\ \forall i,$ (18b)
$A-1\geq\sum_{i=1}^{L_{p}}\delta_{mi},$ (18c)
where $\bm{\delta}_{m}=[\delta_{m1},\delta_{m2},\cdots,\delta_{mL_{p}}]^{\intercal}$. The constraint (18c) represents the necessary condition of Wishart matrix for interference suppression. Also, intuitively, this would otherwise lead to invalid negative values for the desired signal power and coherent interference in the SINR expression.
The optimization problem in (18) is a Mixed-Integer Non-Linear Program (MINLP), a class of problems known for its exponential computational cost. To develop a more tractable formulation, we begin by relaxing the binary constraint (18b) into a continuous one:
$0\leq\delta_{mi}\leq 1\ \forall i.$ (19)
This relaxation transforms the original MINLP into a Non-Linear Program (NLP). To ensure that the solutions of this new NLP satisfy the original binary requirement ($\delta_{mi}\in{0,1}$), we introduce the following complementary constraint for each $i$:
$\delta_{mi}-\delta^{2}_{mi}\leq 0.$ (20)
The combined constraints (19) and (20) are equivalent to the original binary constraint (18b), as the only values between 0 and 1 that satisfy $\delta_{mi}(1-\delta_{mi})\leq 0$ are precisely 0 and 1. Thus, the problem can be written as
$\underset{\bm{\delta}_{m}}{\max~}$ $\ \sum\limits_{t=1}^{T}\log(1+\text{SINR}^{\text{G-PFZF}}_{mt}),$ (21a)
[…]
This problem can be solved using non-linear solvers or the successive convex approximation optimization method. Both of these are computationally expensive for large network deployment. Therefore, to solve this problem with low computation cost, we employ gradient based proximal gradient ascent (PGA) method [22, 23]. We first reformulate the problem to be unconstrained except for the box constraint (19). This is achieved by moving the other constraints into the objective function via a penalty method. The reformulated problem is:
$\underset{\bm{\delta}_{m}}{\max~}$ $f(\bm{\delta}_{m}){=}\sum\limits_{t=1}^{T}\log_{2}(1+\text{SINR}^{\text{G-PFZF}}_{mt})-\chi\Bigg(\!\lambda_{1}\!\sum_{i=1}^{L_{p}}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)^{2}+\lambda_{2}\max\!\Big(0,\sum_{i=1}^{L_{p}}\delta_{mi}-A+1\Big)^{2}\!\Bigg),$ (22a)
[…]
where $\chi$ is a penalty parameter, and $\lambda_{1}$ and $\lambda_{2}$ are penalty weights.
Lemma 1.
: For a sufficiently large penalty parameter $\chi^{*}$, the solution to Problem (22) is equivalent to the solution of Problem (21).
Proof: As $\chi\rightarrow+\infty$, the penalty terms $\chi\lambda_{1}$ and $\chi\lambda_{2}$ enforce that any constraint violation is driven to zero for the solution to remain finite. Given that the original problem’s feasible set is bounded (as argued in [24, Proposition 1]), there exists a finite $\chi^{*}$ such that the solutions coincide.
To solve Problem (22) using the PGA method at each AP independently, we follow the steps outlined in Algorithm 1. Since the objective function (22a) is smooth and differentiable, the proximal step reduces to a projection. We initialize $\bm{\delta}_{m}$ and update it along with the gradient ascent direction of $f(\bm{{\delta}}^{(j)}_{m})$ with step size $\alpha$, followed by projection onto $[0,1]$ to satisfy the box constraint (19):
$\bm{{\delta}}^{(j+1)}_{m}=\text{proj}_{[0,1]}\!\left(\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)=\min\!\left(1,\max\!\left(0,\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)\right).$ (23)
The components of gradient $\nabla f(\bm{\delta}_{m})=\Big[\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{m1}},\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{m2}},\dots,\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mL_{p}}}\Big]^{\intercal}$ can be written as
$\begin{split}\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}=&\sum_{t=1}^{T}\frac{I_{mt}}{I_{mt}+S_{mt}}\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}-\chi\Bigg(\!2\lambda_{1}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)\Big(1-2\delta_{mi}\Big)\\ &+2\lambda_{2}\max\!\Big(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\Big)\!\Bigg),\end{split}$ (24)
where
$\frac{\partial S_{mt}}{\partial\delta_{mi}}=\begin{cases}-p^{u}_{t}\Big(1+\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\delta_{mj}\Big)\gamma_{mt},&\text{if }\ i_{t}=i,\\ -p^{u}_{t}\delta_{mi_{t}}\gamma_{mt},&\text{if }\ i_{t}\neq i,\end{cases}$ (25)
$\frac{\partial I_{mt}}{\partial\delta_{mi}}=\begin{cases}-\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{u}_{k}\Big(1+\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\delta_{mj}\Big)\gamma_{m}-\sum\limits_{k\in\mathcal{P}_{i_{t}}}p^{u}_{k}\gamma_{mk},&\text{if }\ i_{t}=i,\\ -\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{u}_{k}\delta_{mi_{t}}\gamma_{mk}-\sum\limits_{k\in\mathcal{P}_{i}}p^{u}_{k}\delta_{mi_{t}}\gamma_{mk},&\text{if }\ i_{t}\neq i,\end{cases}$ (26)
1:for all $m\in\mathcal{M}$ independently do
2:  Initialize: $j=1$, $k=1$, $\bm{\delta}_{m}^{(1)}$, $s^{(k)}=F(\bm{\delta}_{m}^{(1)})$, $\chi=1$, $\Delta>1$, $\epsilon=5e^{-3}$
3:  repeat
4:   repeat
5:    Compute $\bm{{\delta}}^{(j+1)}_{m}$ using (23)
6:    Update $j=j+1$
7:   until $\Big|\frac{f(\bm{{\delta}}^{(j)}_{m})-f(\bm{{\delta}}^{(j-1)}_{m})}{f(\bm{{\delta}}^{(j-1)}_{m})}\leq\epsilon\Big|$
8:   Update $\chi=\chi\times\Delta$
9:   Update $k=k+1$
10:   Update $s^{(k)}=f(\bm{{\delta}}^{(j)}_{m})$
11:  until $\big|\frac{s^{(k)}-s^{(k-1)}}{s^{(k-1)}}\leq\epsilon\big|$
12:end for
Algorithm 1 Proximal Gradient Ascent (PGA) Method
Convergence and Computation Cost Analysis: Since the feasible set defined by the box constraint (19) is bounded and the gradient function $\nabla f(\bm{\delta}_{m})$ is Lipschitz continuous (proof is given in Appendix -B) with constant $L>0$, one should choose $\alpha\in(0,1/L]$ to ensure that the gradient step remains within the stability region of the PGA method [22, 23]. In numerical simulations, small step size parameter $\alpha$ has been observed to achieve convergence for Algorithm 1.
At each AP, the computation cost of Algorithm 1 in each iteration depends upon the gradient computation step. The computation cost of calculating $f(\bm{\delta}_{m})$ is $\mathcal{O}(T^{2})$. Thus, the cost of calculating $\nabla f(\bm{{\delta}}_{m})$ is also $\mathcal{O}(T^{2})$. Since we have only $L_{p}$ variables, it converges in a few iterations only.
Note 2: To reduce the computational cost, we use pilot-based grouping instead of UE-based grouping, as discussed in [19, 13]. In UE-based grouping, each AP would have $T$ binary variables (one per UE), leading to $TM$ total variables across the network. In the proposed pilot-based grouping, each AP has $L_{p}$ binary variables, leading to $ML_{p}$ total variables. Since $L_{p}\ll T$ in dense deployments, this significantly reduces the computational cost.
III-B Generalized Protected Weak Partial Full-Pilot Zero Forcing Combining Scheme
In the G-PFZF scheme, the combiner for the UEs assigned to a strong pilot is designed only to suppress interference only from other UEs that are also on strong pilots. This intra-group suppression leaves UEs on weak pilots vulnerable to dominant interference from the strong UEs.
The proposed G-PWPFZF scheme introduces a more comprehensive, protective strategy. In this framework, the combiners for all UEs, whether assigned to a strong or weak pilot, are designed to actively suppress interference from all UEs that use a strong pilot. This is achieved by projecting the combining vectors for weak UEs onto the orthogonal complement of the subspace spanned by the strong UEs’ effective channels. Consequently, the G-PWPFZF scheme provides universal protection against the most significant sources of interference, dramatically improving performance for far users on weak pilots, at cost of higher computation cost of designing the vector for weak pilot UEs.
In the G-PWPFZF scheme, the strong pilots are assigned with the vectors in the same way as in (7). The combining vector allocated to the pilot $i_{t}\in\mathcal{W}_{m}$ corresponding to the AP $m$ is simply the normalized MR combining vector, as in (8), with a projection matrix and is given by:
${\mathbf{v}}^{\text{PMR}}_{mi_{t}}=\frac{1}{\sqrt{(A-L_{\mathcal{S}_{m}})\theta_{mi_{t}}}}\mathbf{B}_{m}\mathbf{\bar{G}}_{m}e_{i_{t}},$ (27)
where $\mathbf{B}_{m}$ is a projection matrix that projects the received signal onto the orthogonal complement of the subspace spanned by the effective channels of all strong pilots. It is defined as [19]:
$\mathbf{B}_{m}=\mathbf{I}_{A}-\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}})^{-1}\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m},$
This projection ensures that the combiner for a weak pilot is orthogonal to the channel estimates of all strong-pilot UEs, thereby nullifying the dominant interference from that group.
The expected value of the effective channel gain for the UE $t$ such that $i_{t}\in\mathcal{W}_{m}$
$\mathbb{E}\{(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}_{mk}}\}=\begin{cases}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}.\end{cases}$ (28)
Furthermore, the second moment of the effective channel gain is given by:
$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}}+1)\gamma_{mk}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},&\text{if }i_{k}\neq i_{t},\end{cases}$ (29)
These equations confirm that the projected MR (PMR) combiner successfully nullifies interference from all UEs sharing strong pilots $i_{k}\in\mathcal{S}_{m}$. This design effectively mitigates the most significant intra-network interference, enhancing the performance of the UEs with weaker channel conditions.
By using the combining vector (7) for strong pilot and (27), the closed-form expression for achievable SE for G-PWPFZF combining, as in Theorem 1, can be obtained with the SINR given by:
[…]
The derivation of SINR expression is given in Appendix -C.
Local Weights Design for G-PWPFZF Scheme: The design of local weights for the G-PWPFZF scheme follows a similar design to that of the G-PFZF scheme, aiming to maximize the local contribution to the final SINR while suppressing interference in a fully distributed manner. Here too the local weights are same as the local SINR.
The closed-form expression of the local SINR for UE $t$ at AP $m$ can be obtained from (2) by expanding it in the same manner as in (4) and following the derivation procedure outlined in Appendix -C. Accordingly, the closed-form of the local SINR or local weight is given by:
$a_{mt}=\text{SINR}^{\text{G-PWPFZF}}_{mt}=\frac{p^{{u}}_{t}(b_{tt}^{m})^{2}}{\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{{u}}_{k}(b_{kt}^{m})^{2}+W_{mt}}$ (31)
where
$b_{kt}^{m}=\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{{m}k}},$
${W}_{{m}t}=\sum\limits_{k=1}^{T}p^{{u}}_{k}(\beta_{{m}k}-\delta_{{m}i_{k}}\gamma_{{m}k})+1.$
By employing these locally computed weights, each AP independently scales its data estimates to prioritize signals with high desired gain and low interference. This process effectively embeds interference suppression functionality at the source, thereby completely eliminating the need for centralized LSFD at the CPU and the associated fronthaul overhead for channel statistics.
Adaptive Pilot Grouping for G-PWPFZF Combining: This subsection discusses the local optimization framework for pilot grouping in the G-PWPFZF scheme. The objective and constraint structure remain identical to those of the G-PFZF scheme. The sole distinction lies in the expression for the local SINR, which forms the foundation of the utility function $f(\bm{\delta}_{m})$ for pilot-grouping. The local SINR for the UE $t$ at the AP $m$ for G-PWPFZF scheme, given in (31) after substituting $L_{\mathcal{S}m}=\sum_{i=1}^{L_{p}}\delta_{mi}$, can be rewritten as:
[…]
Consequently, the optimization algorithm (Algorithm 1) and the overall structure of the gradient $\nabla f(\bm{\delta}_{m})$ remain unchanged. The impact of the different SINR is confined solely to the calculation of the partial derivatives $\frac{\partial S_{mt}}{\partial\delta_{mi}}$ and $\frac{\partial I_{mt}}{\partial\delta_{mi}}$ within the gradient. For the G-PWPFZF SINR in (32), these derivatives are:
$\frac{\partial S_{mt}}{\partial\delta_{mi}}=-p^{u}_{t}\gamma_{mt},$ (33)
$\frac{\partial I_{mt}}{\partial\delta_{mi}}=-\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{u}_{k}\gamma_{mk}-\sum\limits_{k\in\mathcal{P}_{i}}p^{u}_{k}\gamma_{mk}.$ (34)
These expressions replace their G-PFZF counterparts in the gradient calculation step of Algorithm 1. All other steps of the algorithm proceed identically.
IV Numerical Simulations
TABLE I: Baseline Simulation Parameters
Parameter Value
Number of APs ($M$) 100
Antennas per AP ($A$) 8
Number of UEs ($T$) 100
Channel bandwidth ($B$) 20 MHz
Uplink pilot symbols ($L_{p}$) 7
Coherence block length ($L_{c}$) 200 symbols
Maximum UE transmit power ($p_{\text{max}}$) 100 mW
Shadow fading standard deviation 8 dB
This section demonstrates the influence of the proposed improvements on the SE performance across multiple distributed combining strategies and local weights. We consider a network with $M$ APs and $T$ UEs, uniformly distributed in a square area of size $1\times 1$ km2. The parameters are mentioned in Table I. Large-scale fading coefficients (LSFCs) are generated according to the model in [3], incorporating shadow fading. Pilots are assigned using the scalable pilot assignment strategy from [3], ensuring efficient reuse under limited pilot resources. All APs are assumed to operate at full transmit power during both pilot training and data transmission. For the baseline PFZF and PWPFZF schemes, UE grouping is performed using a $90\%$ of total LSFCs received at the AP, which we identified using numerical simulation as providing the best performance. To ensure statistical reliability, all reported results are averaged over 500 independent simulation realizations.
IV-A Computation cost and Fronthaul cost
In this subsection, we analyze the computational cost and fronthaul signaling cost. The costs are computed following the approach outlined in [16, 19, 15, 10], where only multiplication and division operations are counted per coherence block, as these dominate the computational load.
TABLE II: Decoding Computation and Fronthaul Overhead per UE per Coherence Block
Decoding Type Fronthaul Cost (Complex Scalars) Combining Weights Computation Cost
o-LSFD [16, 15] $L_{u}M+\frac{3MT+M}{2}$ $\left(\frac{M^{2}+M}{2}\right)T+\frac{M^{3}-M}{3}+M^{2}$
Proposed Decentralized Decoding $L_{u}M$ $\left(\frac{M^{2}+M}{2}\right)T+M+M^{2}$
Table II compares the fronthaul and computational costs for optimal o-LSFD and the proposed decentralized decoding architecture. In conventional o-LSFD, each AP must transmit both channel estimates and local data estimates to the CPU, resulting in substantial fronthaul overhead of $L_{u}M+\frac{3MT+M}{2}$ complex scalars per UE per coherence block. Furthermore, o-LSFD requires computationally expensive matrix operations such as matrix inversion at the CPU that scale as $\mathcal{O}(TM^{2}+M^{3})$ per UE, making it impractical for large-scale deployments.
The proposed decentralized decoding architecture eliminates these bottlenecks by avoiding centralized LSFD entirely. Each AP independently computes and applies interference-suppressing local weights, reducing the fronthaul load to only $L_{u}M$ complex scalars. The computational cost is similarly streamlined to $\mathcal{O}(TM^{2})$ by eliminating the cubic $M^{3}$ term associated with large matrix inversions. This combination of reduced fronthaul overhead and scalable computation enables practical deployment in dense network scenarios.
Figure 2: Computational cost of weights calculation versus number of UEs, showing the significant computation cost reduction of the proposed decentralized decoding approach compared to o-LSFD.
Figure 3: Fronthaul cost versus number of UEs, demonstrating the reduction achieved by the proposed decentralized decoding architecture by eliminating channel estimate transmission.
As demonstrated in Fig. 2, the proposed decentralized decoding weights computation provides substantial computation cost reduction compared to o-LSFD, particularly as the number of UEs increases. Similarly, Fig. 3 shows that eliminating channel estimate transmission reduces fronthaul overhead.
TABLE III: Computational cost per AP per Coherence Block
Scheme Combining Vector Computation
PFZF[19] $\frac{3L_{\mathcal{S}_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}AT$
G-PFZF $\frac{3L_{\mathcal{S}_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}AL_{p}$
PWPFZF[19] $\frac{3L_{S_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}2(L_{p}{-}L_{\mathcal{S}_{l}})L_{\mathcal{S}_{l}}A{+}AT$
G-PWPFZF $\frac{3L_{S_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}2(L_{p}{-}L_{\mathcal{S}_{l}})L_{\mathcal{S}_{l}}A{+}AL_{p}$
Figure 4: Computational cost of combining vectors versus number of UEs, illustrating the scalability advantage of generalized schemes (G-PFZF, G-PWPFZF) over traditional approaches.
The computational cost per AP per coherence block for calculating the combining vectors is detailed in Table III. The primary distinction between the traditional and proposed generalized schemes lies in their scaling behavior relative to the number of users. The traditional PFZF and PWPFZF schemes exhibit a computational cost that scales linearly with the total number of user equipments, $T$, as in by the $AT$ term for PFZF and PWPFZF. This dependence creates a significant computational bottleneck in densed networks with massive connectivity. In contrast, the proposed G-PFZF and G-PWPFZF schemes achieve a fundamental scalability advantage by scaling with the number of pilots, $L_{p}$, rather than the number of UEs. This is reflected in the $AL_{p}$ term for G-PFZF and G-PWPFZF, which is a fixed system parameter independent of the user population. Fig. 4 further validates the scalability benefits of the generalized schemes, where computational cost saturates once the number of UEs exceeds the available pilot sequences, unlike traditional schemes that continue to scale linearly with user density.
IV-B Performance Evaluation
Figure 5: Distribution of Strong Pilot Decisions: How Each AP Adapts?
Fig. 5 shows the distribution of strong pilot decisions ($L_{\mathcal{S}_{m}}$) across all APs. The baseline PFZF and PWPFZF schemes exhibit a pronounced peak at $L_{\mathcal{S}_{m}}=6$-77, revealing their rigid threshold-based strategy that sacrifices excessive degrees of freedom for interference suppression regardless of local conditions. In striking contrast, the proposed G-PFZF and G-PWPFZF schemes show fundamentally different behavior: G-PFZF peaks at 2-4 strong pilots while G-PWPFZF shows an even more conservative distribution peaking at 1-3. This efficient allocation preserves spatial resources for signal enhancement. Most notably, the significant non-zero count at $L_{\mathcal{S}_{m}}=0$ demonstrates our algorithm’s ability to adaptively default to simple MR combining when it is the optimal strategy for a given AP’s local conditions. This demonstrates a seamless, adaptive switching between local zero-forcing and local MR processing that is impossible for the threshold-bound baseline.
Figure 6: The uplink sum SE comparison for various combining schemes.
Fig. 6 depicts the sum spectral efficiency (SE) versus the number of UEs ($T$) for various combining schemes. The proposed distributed optimization framework delivers significant and scalable performance gains, with G-PFZF outperforming its baseline PFZF by 6.5% to 9.5% and G-PWPFZF surpassing PWPFZF by 5.5% to 10%, with the margin widening as the number of UEs increases. This superior performance highlights the critical limitation of the baseline PFZF and PWPFZF schemes, which rely on a hard, network-wide threshold using LSFCs to rigidly classify users as strong or weak at every AP. In contrast, the proposed framework introduces a distributed optimization algorithm, executed independently by each AP, to determine a better pilot-based grouping strategy. This allows an AP to adaptively select none, a subset, or all of the pilots to designate as strong, based on its local channel conditions and specific geometric configuration. This per-AP flexibility enables a more efficient and dynamic trade-off between utilizing degrees of freedom for interference suppression and desired signal enhancement. Consequently, the proposed G-PWPFZF scheme achieves superior performance over the G-PFZF scheme by suppressing interference from strong UEs for the weak UEs without a significant sacrifice in degrees of freedom.
Figure 7: The uplink sum SE comparison for various combining schemes.
Fig. 7 illustrates the sum SE versus the number of orthogonal pilots $L_{p}$, demonstrating the superior performance of the proposed G-PFZF and G-PWPFZF schemes over their respective baseline counterparts. The proposed schemes achieve significant performance improvements of 6-14% for G-PFZF and 6-13% for G-PWPFZF compared to their respective baselines across different pilot lengths, with the performance gap widening as the number of pilots increases. This expanding margin highlights the advantage of the proposed adaptive user grouping strategy over the rigid network-wide threshold approach used in baseline schemes. The baseline schemes, constrained by a rigid network-wide threshold, must rigidly limit the number of strong pilots to avoid violating the fundamental condition for zero-forcing ($L_{\mathcal{S}_{m}}>A$), which leads to inefficient interference management. In contrast, our proposed distributed framework empowers each AP to independently and optimally select its set of strong pilots, inherently ensuring the selection is both locally optimal and always feasible ($L_{\mathcal{S}_{m}}<A$). This inherent adaptability allows our methods to leverage larger pilot sets more effectively, maximizing array gain and suppressing interference, which explains the observed widening of the performance gap. The G-PWPFZF scheme provides a further advantage by proactively protecting weak UEs from the strong UEs’ interference.
Figure 8: The uplink sum SE comparison for various combining schemes.
Fig. 8 illustrates the sum spectral efficiency (SE) versus the number of antennas per AP ($A$) for various combining schemes, revealing a characteristic performance pattern that highlights the adaptability of our proposed framework. The G-PFZF and G-PWPFZF schemes demonstrate consistent superiority across all antenna configurations, with the performance gap exhibiting a distinctive trajectory that reflects adaptive resource utilization. At the minimal antenna configuration (2 antennas), where spatial degrees of freedom $A-L_{\mathcal{S}_{m}}$ are severely constrained, both proposed and baseline schemes face fundamental challenges in interference suppression. Nevertheless, our adaptive grouping strategy achieves measurable gains of 2.5% for G-PFZF and 6% for G-PWPFZF through an adaptive selection of combining schemes, dynamically switching between the G-PFZF/G-PWPFZF, and the MR at each AP based on local conditions. In the mid-range antenna configuration (4-8 antennas), where careful trade-off between interference suppression and signal enhancement becomes crucial, our proposed method demonstrates performance gain of 10-14% for G-PFZF and 10-18% for G-PWPFZF, showcasing their superior grouping capabilities. As antenna numbers increase further (10-16 antennas), providing abundant spatial resources that benefit all schemes, the performance gap narrows to 4.5-6.5% for G-PFZF and 3.5-5.5% for G-PWPFZF, though our approach maintains consistent superiority. These results confirm that our adaptive optimization framework achieves maximum relative advantage in precisely those scenarios where adaptive resource allocation is most valuable, particularly in resource-constrained environments, while still delivering meaningful gains in antenna-rich deployments.
Figure 9: The uplink sum SE comparison for various combining schemes.
Fig. 9 illustrates the sum spectral efficiency (SE) versus the number of UEs ($T$), comparing local and LSFD-based combining schemes. For our G-PFZF and G-PWPFZF schemes, the performance gap between local weights and optimal LSFD remains minimal at just 3% across all UE densities. This marginal performance difference is dramatically outweighed by the substantial system benefits: our local weight approach eliminates the need for complex network-wide coordination, reducing computational cost at the CPU by avoiding large-scale matrix inversions and cutting fronthaul overhead by eliminating the need to share channel state information across the network. Despite this streamlined architecture, our locally-weighted G-PFZF and G-PWPFZF schemes still outperform the baseline PFZF and PWPFZF schemes with optimal LSFD by margins of 4.5-9% and 3.5-7.5% respectively. This demonstrates that the gains from our adaptive per-AP grouping strategy are so substantial that they outperform the benefits of optimal coordination applied to suboptimal threshold-based grouping. These results confirm that our fully distributed approach achieves an optimal balance between performance and practicality.
Figure 10: The uplink sum SE comparison for various combining schemes.
Fig. 10 illustrates the sum spectral efficiency (SE) versus the number of pilots ($L_{p}$), comparing local and LSFD-based combining schemes. The performance gap between local and LSFD-based weights is less than 5.5% for sparse pilot configurations, where high pilot contamination makes centralized LSFD better equipped to handle interference. However, as the number of pilots increases to moderate and high values, the gap reduces to less than 1%, demonstrating near-identical performance. This trend occurs because with sufficient orthogonal pilots, both approaches have similar capability to mitigate pilot contamination. Notably, even in the worst-case scenario with limited pilots, the performance penalty remains modest, validating that the substantial fronthaul and complexity reductions of our decentralized decoding approach come with acceptable performance trade-offs across all pilot sequences. Furthermore, the proposed G-PFZF and G-PWPFZF schemes with local weights consistently outperform traditional PFZF and PWPFZF with LSFD across moderate to high pilot lengths, confirming the advantage of adaptive per-AP combining strategies.
Figure 11: The uplink SE per user comparison for various combining schemes.
Fig. 11 presents the 90%-likely per-user spectral efficiency across different combining schemes, demonstrating remarkable improvements in user fairness through our distributed optimization framework. The proposed G-PFZF scheme achieves a substantial 45% improvement in 90%-likely SE compared to conventional PFZF, while G-PWPFZF shows a 34% gain over its baseline PWPFZF. These dramatic improvements confirm that our adaptive per-AP grouping strategy not only enhances overall system capacity but fundamentally transforms user experience across the network. The gains in 90%-likely SE particularly highlight our framework’s effectiveness in improving the performance of weak users who traditionally suffer from poor service quality. These results demonstrate that our distributed optimization framework successfully addresses both system-level efficiency and user-level fairness requirements in D-mMIMO deployments, achieving gains in quality-of-service uniformity.
Figure 12: The uplink SE per user comparison for various combining schemes.
Fig. 12 presents the 90%-likely per-user spectral efficiency comparing local versus LSFD processing, revealing crucial insights about the performance-complexity trade-off in our distributed framework. While our G-PFZF and G-PWPFZF schemes with local weights experience a modest performance reduction of approximately 10-10.7% compared to their LSFD-based counterparts, the 10% performance difference between local and LSFD processing represents a reasonable trade-off for achieving full distributability, substantially reduced fronthaul overhead, and lower computational cost. Remarkably, our locally-weighted G-PFZF scheme outperforms the baseline PFZF with optimal LSFD by 32%, and G-PWPFZF with local weights surpasses the LSFD-enhanced PWPFZF baseline by 22.5%. These results demonstrate that the performance gains from our adaptive per-AP grouping strategy are so significant that they outweigh the benefits of optimal LSFD coordination applied to suboptimal threshold-based grouping. This analysis confirms that our distributed optimization framework provides a balance between performance and practical implementation constraints, delivering better SE per user while maintaining architectural advantages crucial for deployments.
V Conclusion
This paper has addressed the critical bottlenecks of sub-optimal performance, high fronthaul load, and computational cost in distributed massive MIMO networks by introducing a novel, decentralized decoding uplink architecture that fundamentally departs from conventional LSFD-based designs.
Through the proposed local ZF framework, we have demonstrated that enabling each AP to independently determine its combining strategy via local optimization, classifying pilots as strong or weak without fixed thresholds and dynamically switching among PFZF/PWPFZF, FZF, or MR, yields significant performance improvements. The resulting generalized schemes, G-PFZF and G-PWPFZF, consistently outperform their fixed-threshold counterparts across all evaluated scenarios, achieving substantial gains in both sum and per user spectral efficiency. Moreover, the introduction of pilot-dependent combining vectors and interference-suppressing local weights applied distributively at each AP eliminates the need for centralized LSFD, thereby drastically reducing fronthaul overhead and computational cost. Remarkably, this distributed approach incurs only a minimal performance penalty compared to idealized centralized coordination, while even outperforming conventional threshold-based schemes using optimal LSFD.
These findings collectively establish that adaptive processing at the network, not complex centralized decoding, is the key to scalability and efficiency in future D-mMIMO systems. The proposed framework offers a practical and high-performance pathway toward realizing scalable distributed MIMO networks without compromising on quality-of-service or imposing prohibitive infrastructure costs.
-A Proof of closed-form SINR expression for G-PFZF combining:
To derive the closed-form expression, we need few expectation properties,
$\begin{split}\mathbb{E}\Big\{\Big|\sum_{m}a_{m}\mathbf{v}_{m}^{H}\mathbf{g}_{m}\Big|^{2}\Big\}=\sum_{m}|a_{m}|^{2}\mathbb{E}\Big\{\Big|\mathbf{v}_{m}^{H}\mathbf{g}_{m}\Big|^{2}\Big\}+\Big|\sum_{m}a_{m}\mathbb{E}\Big\{\mathbf{v}_{m}^{H}\mathbf{g}_{m}\Big\}\Big|^{2}-\sum_{m}\Big|a_{m}\mathbb{E}\Big\{\mathbf{v}_{m}^{H}\mathbf{g}_{m}\Big\}\Big|^{2},\end{split}$ (35)
Also, if $\mathbf{v}_{m}$ and $\mathbf{g}_{m}$ are indpendent random vectors of N length with zero mean and elements following i.i.d., then
$\mathbb{E}\Big\{\Big|\mathbf{v}_{m}^{H}\mathbf{g}_{m}\Big|^{2}\Big\}=\frac{\mathbb{E}\Big\{\Big|\mathbf{v}_{m}\Big|^{2}\Big\}\mathbb{E}\Big\{\Big|\mathbf{g}_{m}\Big|^{2}\Big\}}{N}.$ (36)
We derive the closed-form expression of desired signal term in (6) using the combining vectors in (7) and (8).
$|\text{DS}_{t}|^{2}=p^{u}_{t}\Big|\mathbb{E}\Big\{\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mt}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mt}\Big\}\Big|^{2},$ (37)
where $\mathcal{Z}_{i_{t}}$ is the subset of APs for which pilot $i_{t}$ is strong and $\mathcal{Y}_{i_{t}}$ is the subset of APs for which pilot $i_{t}$ is weak. Using (10) and (12), the (37) can be written as:
$|\text{DS}_{t}|^{2}=p^{u}_{t}\Bigg|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}\sqrt{A\gamma_{mk}}\Bigg|^{2}=p^{u}_{t}\Bigg|\sum_{m=1}^{M}a_{mt}\sqrt{(A-\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2},$ (38)
where $\delta_{mi_{t}}$ is defined in (15). The first term of interference in (6) can be expanded as:
$\mathbb{E}\Big\{|\text{BU}_{t}|^{2}\Big\}=p^{u}_{t}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mt}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mt}\Big|^{2}\Bigg\}-|\text{DS}_{t}|^{2},$ (39)
We focus on first term of $\mathbb{E}\Big\{|\text{BU}_{t}|^{2}\Big\}$, which can be written as:
$\begin{split}\!\!\!\!\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}\Big|^{2}\!\Bigg\}+\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mt}\Big|^{2}\Bigg\}+2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nt}\!\Bigg\},\end{split}$ (40)
The first term of (40) can be written as:
$\begin{split}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!(\hat{\mathbf{g}}_{mt}+\tilde{\mathbf{g}}_{mt})\Big|^{2}\!\Bigg\}&=\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mt}\Big|^{2}\!\Bigg\}+\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mt}\Big|^{2}\!\Bigg\}\\ &\overset{(a)}{=}\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\mathbb{E}\Big\{\!\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mt}\Big|^{2}\!\Big\}+\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mt}\!\Big\}\Big|^{2}\\ &-\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mt}\!\Big\}\Big|^{2}+\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\mathbb{E}\Big\{\!\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mt}\Big|^{2}\!\Big\}\\ &+\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mt}\!\Big\}\Big|^{2}-\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mt}\!\Big\}\Big|^{2}\\ &\overset{(b)}{=}\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(A-L_{\mathcal{S}_{m}})\gamma_{mt}+\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}\\ &-\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(\beta_{mt}-\gamma_{mt})\\ &\overset{(c)}{=}\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(\beta_{mt}-\gamma_{mt})\end{split}$ (41)
we get (a) using same way as in (35) and (b) using the (10), (11) and (36). Similarly, the second term of (40) is first expanded same as in (a) in (41) using (35), then using the (12), (13) and (36), can be written as
$\begin{split}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!(\hat{\mathbf{g}}_{mt}+\tilde{\mathbf{g}}_{mt})\Big|^{2}\!\Bigg\}&{=}\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(A+1)\gamma_{mt}+\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mt}}\Big|^{2}\\ &-\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\sqrt{A\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(\beta_{mt}-\gamma_{mt})\\ &=\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\beta_{mt}.\end{split}$ (42)
The expectation in closed-form in third term of (40) can be written as:
$\begin{split}2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nt}\!\Bigg\}=2\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Bigg)\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!a_{mt}\sqrt{A\gamma_{mt}}\Bigg).\end{split}$ (43)
Using (41), (42) and (43) in (40) and using this in (39), we get
$\mathbb{E}\Big\{|\text{BU}_{t}|^{2}\Big\}=|\text{DS}_{t}|^{2}+p^{u}_{t}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mt}-\delta_{mi_{t}}\gamma_{mt})-|\text{DS}_{t}|^{2}=p^{u}_{t}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mt}-\delta_{mi_{t}}\gamma_{mt}).$ (44)
The second term i.e., the pilot contamination term for the UE $t$ with the pilot sharing UE $k$ in (6), can be evaluated in closed-form as
$\begin{split}&\mathbb{E}\Big\{|\text{PC}_{tk}|^{2}\Big\}=p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mk}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mk}\Big|^{2}\Bigg\},\end{split}$ (45)
The RHS of (45) can be written as
$\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\Bigg\}+\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}+2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}$
Following the same steps as in (41), (42) and (43)
$\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}=\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}+\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(\beta_{mk}-\gamma_{mk}),$ (46)
$\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}{=}\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mk}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\beta_{mk},$ (47)
$2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}=2\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg)\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!a_{mt}\sqrt{A\gamma_{mk}}\Bigg),$ (48)
Therefore
$\begin{split}&\mathbb{E}\Big\{|\text{PC}_{tk}|^{2}\Big\}=p^{u}_{k}\Big|\sum_{m=1}^{M}a_{mt}\sqrt{(A-\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}+p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (49)
The third interference term, for any non-pilot sharing UEs pair $t,k$, in (6) can be written as:
$\begin{split}\mathbb{E}\Big\{|\text{UI}_{tk}|^{2}\Big\}&=p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mk}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\Bigg\}+p^{u}_{k}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}\\ &+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}(1-\delta_{mi_{k}})\gamma_{mk}+p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}(\beta_{mk}-\gamma_{mk})\\ &+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\gamma_{mk}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}(\beta_{mk}-\gamma_{mk})\\ &=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (50)
The last term of interference in (6) can be written as:
$\begin{split}\mathbb{E}\Big\{|\text{GN}_{t}|^{2}\Big\}&=p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}+p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (51)
Using the (38), (44), (49), (50) and (51) in (6) gives the SINR closed-form expression in (14).
-B Proof of Lipschitz Continuity of $\nabla f(\bm{\delta}_{m})$:
To establish the Lipschitz continuity of $\nabla f(\bm{\delta}_{m})$, we analyze its component functions and their boundedness over the compact domain $\bm{\delta}_{m}\in[0,1]^{L_{p}}$.
Consider the gradient component:
$g(\delta_{mi})=\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}=$ $\sum_{t=1}^{T}\frac{I_{mt}}{I_{mt}+S_{mt}}\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}$
$-\chi\left[2\lambda_{1}\max\left(0,\delta_{mi}-\delta^{2}_{mi}\right)\left(1-2\delta_{mi}\right)+2\lambda_{2}\max\left(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\right)\right]$
Boundedness Analysis:
Signal and Interference Terms: For $\delta_{mi}\in[0,1]$:
  • $S_{mt}>0$ and $I_{mt}>0$ (positive definite)
  • $S_{mt},I_{mt}\leq C_{1}$ (bounded above by network parameters)
  • $\frac{\partial S_{mt}}{\partial\delta_{mi}},\frac{\partial I_{mt}}{\partial\delta_{mi}}\leq C_{2}$ (bounded derivatives from (25) and (26))
Penalty Terms:
  • $\max(0,\delta_{mi}-\delta^{2}_{mi})\in[0,\frac{1}{4}]$ (bounded on $[0,1]$)
  • $\max\left(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\right)\leq L_{p}$ (finite sum constraint)
Composite Function Bounds:
  • $\frac{I_{mt}}{I_{mt}+S_{mt}}\in(0,1)$
  • $\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}\leq C_{3}$
Lipschitz Constant Derivation:
The partial derivatives of $g(\delta_{mi})$ exist and are bounded on the compact set $0,1]^{L_{p}}$:
$\begin{split}\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}&=\sum_{t=1}^{T}\Bigg[\frac{S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}-I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}}{\big(I_{mt}+S_{mt}\big)^{2}}\cdot\frac{I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}\\ &\quad+\frac{I_{mt}}{I_{mt}+S_{mt}}\left(\frac{I_{mt}\,\frac{\partial^{2}S_{mt}}{\partial\delta_{mi}^{2}}-S_{mt}\,\frac{\partial^{2}I_{mt}}{\partial\delta_{mi}^{2}}}{I_{mt}^{2}}-\frac{2\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big(I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big)}{I_{mt}^{3}}\right)\Bigg]\\ &\quad-\chi\Bigg[2\lambda_{1}\,\mathbf{1}_{\{\delta_{mi}-\delta_{mi}^{2}>0\}}\Big((1-2\delta_{mi})^{2}-2(\delta_{mi}-\delta_{mi}^{2})\Big)+2\lambda_{2}\,\mathbf{1}_{\{2-A>0\}}\Bigg].\end{split}$
For all $\delta_{mi}\in[0,1]$, $\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}$ is bounded:
$\left|\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}\right|\leq C_{4}$
Since all component functions are bounded and continuously differentiable on the compact domain, there exists a global Lipschitz constant $L$ such that:
$\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\|\leq L\|\mathbf{x}-\mathbf{y}\|,\quad\forall\mathbf{x},\mathbf{y}\in[0,1]$
This establishes the Lipschitz continuity of $\nabla f(\bm{\delta}_{m})$, ensuring convergence of the proximal gradient algorithm with appropriate step size selection.
-C Proof of closed-form SINR expression for G-PWPFZF combining:
The derivation of G-PWPFZF SINR closed-form expression follows the same step as the derivation of G-PFZF SINR closed-form expression, the only difference is that instead of ${\mathbf{v}}^{\text{MR}}_{mi_{t}}$, we have ${\mathbf{v}}^{\text{PMR}}_{mi_{t}}$ combining vector.
The desired signal expression in (6) using the combining vectors in (7) and (27).
$|\text{DS}_{t}|^{2}=p^{u}_{t}\Big|\mathbb{E}\Big\{\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mt}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\mathbf{{g}}_{mt}\Big\}\Big|^{2},$ (52)
Using (10) and (28), the (52) can be written as:
$|\text{DS}_{t}|^{2}=p^{u}_{t}\Bigg|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2}=p^{u}_{t}\Bigg|\sum_{m=1}^{M}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2},$ (53)
The first term $\mathbb{E}\Big\{|\text{BU}_{t}|^{2}\Big\}$ of interference in (6) can be expanded as:
$\begin{split}\!\!\!\!\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}\Big|^{2}\!\Bigg\}+\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{mt}\Big|^{2}\Bigg\}+2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}(\mathbf{v}_{ni_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{nt}\!\Bigg\},\end{split}$ (54)
The first term of (54) is same as (41), the second term of (54) can be written as
$\begin{split}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\!(\hat{\mathbf{g}}_{mt}+\tilde{\mathbf{g}}_{mt})\Big|^{2}\!\Bigg\}&{=}\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(A-L_{\mathcal{S}_{m}}+1)\gamma_{mt}+\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}\\ &-\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(\beta_{mt}-\gamma_{mt})\\ &=\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\beta_{mt}.\end{split}$ (55)
The expectation in closed-form in third term of (54) can be written as:
$\begin{split}2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}(\mathbf{v}_{ni_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{nt}\!\Bigg\}=2\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Bigg)\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Bigg).\end{split}$ (56)
Using (41), (55) and (56) in (54), we get
$\mathbb{E}\Big\{|\text{BU}_{t}|^{2}\Big\}=p^{u}_{t}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mt}-\delta_{mi_{t}}\gamma_{mt}).$ (57)
The second term i.e., the pilot contamination term $\mathbb{E}\Big\{|\text{PC}_{tk}|^{2}\Big\}$ for UE $t$ with pilot sharing UE $k$ in (6), can be evaluated as
$\begin{split}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\Bigg\}+\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}+2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}\end{split}$ (58)
The first term of (58) is same as (46). The second and third term of (58) can be computed as:
$\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\!{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}{=}\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\beta_{mk},$ (59)
$2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}=2\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg)\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg),$ (60)
Using (46), (59) and (60)
$\begin{split}&\mathbb{E}\Big\{|\text{PC}_{tk}|^{2}\Big\}=p^{u}_{k}\Big|\sum_{m=1}^{M}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}+p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (61)
The third interference term, for any non-pilot sharing UEs pair $t,k$, in (6) can be written as:
$\begin{split}\mathbb{E}\Big\{|\text{UI}_{tk}|^{2}\Big\}&={=}p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}\\ &+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}\\ &{=}p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}(1-\delta_{mi_{k}})\gamma_{mk}+p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}(\beta_{mk}-\gamma_{mk})\\ &+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}(1-\delta_{mi_{k}})\gamma_{mk}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}(\beta_{mk}-\gamma_{mk})\\ &=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (62)
The last term of interference in (6) can be written as:
$\begin{split}\mathbb{E}\Big\{|\text{GN}_{t}|^{2}\Big\}&=p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}+p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (63)
Using the (53), (57), (61), (62) and (63) in (6) gives the SINR closed-form expression in (30).

### v2

@@
 A major bottleneck in uplink distributed massive multiple-input multiple-output networks is the sub-optimal performance of local combining schemes, coupled with high fronthaul load and computational cost inherent in centralized large scale fading decoding (LSFD) architectures.
-This paper introduces a decentralized decoding architecture that fundamentally breaks from the conventional LSFD, by allowing each AP calculates interference-suppressing local weights independently and applies them to its data estimates before transmission.
-Furthermore, two generalized local zero-forcing (ZF) framework, generalized partial full-pilot ZF (G-PFZF) and generalized protected weak PFZF (G-PWPFZF), are introduced, where each access point (AP) adaptively and independently determines its combining strategy through a local sum spectral efficiency optimization that classifies user equipments (UEs) as strong or weak using only local information, eliminating the fixed thresholds used in PFZF and PWPFZF.
-To further enhance scalability, pilot-dependent combining vectors instead of user-dependent ones are introduced and are shared among users with the same pilot.
-The corresponding closed-form spectral efficiency expressions are derived.
-Numerical results show that the proposed generalized schemes consistently outperform fixed-threshold counterparts, while the introduction of local weights yields lower overhead and computation costs with minimal performance penalty compared to them.
+This paper introduces a decentralized decoding architecture that fundamentally breaks from the LSFD, by allowing each access point (AP) to calculate interference-suppressing local weights independently and apply them to its data estimates before transmission.
+Furthermore, two generalized local zero-forcing (ZF) frameworks, generalized partial full-pilot ZF (G-PFZF) and generalized protected weak PFZF (G-PWPFZF), are introduced, where each AP adaptively and independently determines its combining strategy through a local sum spectral efficiency (SE) optimization that classifies user equipments (UEs) as strong or weak, eliminating the fixed thresholds used in the PFZF and PWPFZF schemes.
+To enhance scalability, pilot-dependent combining vectors instead of user-dependent ones are introduced and are shared among users with the same pilot.
+The closed-form SE expressions corresponding to the proposed schemes are derived.
+Numerical results show that the proposed schemes consistently outperform fixed-threshold counterparts, while the introduction of local weights yields lower overheads and computation costs with lower performance penalty compared to them.
 I Introduction
-Distributed massive MIMO (D-mMIMO) networks are a promising framework for next-generation wireless systems, as they can enhance weak users’ signal and guarantee consistent service quality across wide coverage areas by enabling numerous distributed access points (APs) to coherently serve user equipments (UEs) [1, 2].
-Unlike conventional cellular architectures, such distributed systems enhance coverage and reliability through cooperative processing and coherent transmission/reception [3].
-To fully use the advantages of D-mMIMO, while ensuring scalability, distributed signal processing at the APs is essential [4].
+Distributed massive MIMO (D-mMIMO) networks are a promising framework for next-generation wireless systems, as they can enhance weak users’ signal and guarantee uniform service quality across wide coverage areas by enabling numerous distributed access points (APs) to coherently serve user equipments (UEs) [1, 2].
+To fully use the advantages of D-mMIMO, while ensuring scalability, distributed signal processing at the APs is essential [3, 4].
 A major challenge in this regard is the development of effective uplink combining schemes that can alleviate inter-user interference within practical limitations.
@@
 These partial centralized ZF schemes reduce the fronthaul load compared to fully centralized ZF, but the overhead still scales with the number of UEs, as for each UE, a subset of APs need to send the channel estimates to the CPU, which limits their suitability for ultra-dense network scenarios.
-These centralized and partial centralized combining schemes follow the the fully centralized architecture as shown in Fig.
-1, which has very high computational cost and fronthaul overhead, as for all UEs, channel need to be estimated at the CPU with the cooperation of all or a subset of APs.
+These centralized and partial centralized combining schemes follow the fully centralized architecture as shown in Fig.
+1, which has very high computational cost and fronthaul overhead, as for all UEs, the channel needs to be estimated at the CPU with the cooperation of all or a subset of APs.
 Several distributed combining approaches have been proposed in the literature.
 Maximum ratio (MR) combining, introduced in [1], is computationally simple and fronthaul-efficient, but its performance is fundamentally limited by its inability to effectively mitigate inter-user interference.
-To improve upon MR, a local partial MMSE scheme is proposed in [3, 10], which achieves superior interference suppression and performance.
+To improve upon the MR, a local partial MMSE scheme is proposed in [3, 10], which achieves superior interference suppression and performance.
 However, a key drawback of this approach is that closed-form expressions for the spectral efficiency (SE) cannot be derived.
-As a result, resource allocation and system optimization must rely on computationally expensive Monte Carlo simulations, and the lack of tractable expressions makes theoretical performance analysis and gaining insights into the system behavior difficult to obtain.
-In [1, 3], local combining schemes follow an architecture of simple decoding as shown in Fig.
-1, thus reducing the overhead apart from local data estimates, yet it suffers from sub-optimality.
-(a) FullyCentralized(b) Local Combining and LSFD(c) Local Combining and Simple decoding (d) Proposed Decentralized decoding APs CPU Received SignalsChannel Estimates Global Data Estimates $\hat{x}_{t}$ APs Local Combining $\mathbf{v}_{mt}^{H}\mathbf{y}_{m}\rightarrow\hat{x}_{mt}$ CPU Local Data Estimates $\hat{x}_{mt}$Local Channel Estimates Calculates LSFD Weights $\mathbf{a}_{t}$ Final Estimates $\hat{x}_{t}=\mathbf{a}_{t}^{H}\hat{\mathbf{x}}_{t}$ APs Local Combining $\mathbf{v}_{mt}^{H}\mathbf{y}_{m}\rightarrow\hat{x}_{mt}$ CPU Local Data Estimates $\hat{x}_{mt}$ Final Estimates $\hat{x}_{t}=\sum_{m}\hat{x}_{mt}$ APs Local Combining $\mathbf{v}_{mt}^{H}\mathbf{y}_{m}\rightarrow\hat{x}_{mt}$ Local Weighting $a_{mt}\cdot\hat{x}_{mt}$ CPU Pre-weighted DataEstimates $a_{mt}\hat{x}_{mt}$ Final Data Estimates $\hat{x}_{t}=\sum_{m}a_{mt}\hat{x}_{mt}$
-Figure 1: Comparison of uplink processing architectures in D-mMIMO: (a) Fully centralized processing is computationally prohibitive.
-(b) Local combining and LSFD requires sharing soft estimates and global CSI to compute optimal weights at the CPU.
-(c) Local combining and Non-coherent decoding is simple but sub-optimal.
-(d) The proposed decentralized decoding architecture: each AP performs local combining and weighting independently, reducing the CPU to a simple aggregator and eliminating fronthaul overhead for coordination.
-The local combining schemes such as full-pilot ZF (FZF), partial FZF (PFZF), and protected weak PFZF (PWPFZF) are proposed in [13, 14], offering tractable closed-form spectral efficiency (SE) expressions.
-These techniques construct combining vectors locally at each AP using channel statistics to suppress interference within designated user groups.
-In traditional FZF, PFZF, and PWPFZF, each AP generates a single combining vector per orthogonal pilot sequence and assigns scaled versions of this vector to individual UEs based on their estimated channel variances [13].
-The fundamental limitations of PFZF and PWPFZF stem from two factors.
+As a result, resource allocation and system optimization must rely on computationally expensive Monte Carlo simulations, and the lack of tractable expressions makes theoretical performance analysis and gaining insights into the system behavior difficult.
+In [1, 3], the authors follow a simple decoding architecture as in Fig.
+1, thus reducing the overhead apart from local data estimates, yet it suffers from sub-optimality with respect to centralized decoding.
+The local combining schemes such as full-pilot ZF (FZF), partial FZF (PFZF), and protected weak PFZF (PWPFZF) are proposed in [13, 14], offering tractable closed-form SE expressions.
+In these schemes, each AP classifies the UEs into strong and weak groups, generates one combining vector per orthogonal pilot: the ZF for strong UEs and the MR for weak UEs, and assigns scaled versions to individual UEs based on their estimated channel variances [13].
+The fundamental limitations of the PFZF and PWPFZF schemes stem from two factors.
 First, the rigid, threshold-based user grouping may misclassify UEs, leading to suboptimal interference suppression and inefficient use of spatial resources.
@@
 The large-scale fading decoding (LSFD) as shown in Fig.
-1, is employed at the central processing unit (CPU) to enhance uplink performance in D-mMIMO networks [8, 15, 13, 14].
-In the optimal LSFD (o-LSFD) approach [13], each AP must transmit both channel estimates and data estimates for all served UEs to the CPU, effectively doubling the fronthaul load while imposing computational costs that scale with the number of users [16].
-To alleviate this burden, partial LSFD (p-LSFD) is proposed in [15], which reduces computational cost and scalability requirements by considering only partial interference in the combining process.
-However, p-LSFD still requires matrix inversions of the same dimension as o-LSFD and the transmission of channel estimates to the CPU.
-As network size and UE density increase, these operations become increasingly demanding with respect to computation and overhead, significantly limiting the scalability and practicality of LSFD-based architectures in ultra-dense D-mMIMO deployments.
+1, is employed at the CPU to enhance uplink performance in D-mMIMO networks [8, 15, 13, 14].
+In the optimal LSFD approach [13], each AP must transmit both channel and data estimates for all served UEs to the CPU, effectively doubling the fronthaul load while imposing computational costs that scale with the number of users [16].
+To alleviate this burden, partial LSFD is proposed in [15], which reduces computational cost and scalability requirements by considering only partial interference in the combining process.
+However, partial LSFD still requires matrix inversions of the same dimension as optimal LSFD and transmission of the channel estimates to the CPU.
+As network size and UE density increase, these operations become increasingly demanding with respect to computation and overhead, significantly limiting the scalability and practicality of the LSFD-based architectures in ultra-dense D-mMIMO deployments.
+(a) FullyCentralized(b) LSFD(c) SimpleDecoding(d) Proposed DecentralizedDecoding APs CPU All Data Global Combining Global Estimates APs Local Combining CPU Local estimates $\hat{x}_{mt}$Channel estimates LSFD Weights Final Estimates APs Local Combining CPU $\hat{x}_{mt}$ $\hat{x}_{t}=\sum\limits_{m}\hat{x}_{mt}$ APs Local Combining Local Weights $a_{mt}$ CPU $a_{mt}\hat{x}_{mt}$ $\hat{x}_{t}=\sum\limits_{m}a_{mt}\hat{x}_{mt}$
+Figure 1: Uplink processing architectures: (a) Fully centralized, (b) LSFD with coordination, (c) Simple decoding, (d) Proposed decentralized decoding.
 I-A Contributions
-To overcome the above mentioned issues, this work introduces a novel decentralized decoding framework for D-mMIMO that fundamentally enhances scalability, adaptability, and analytical tractability.
-The principal contributions are summarized as follows:
+This work addresses some of the above mentioned issues. Our principal contributions are as follows:
 Decentralized Decoding Architecture with Local Weighting: We propose a novel uplink processing architecture as shown in Fig.
 1, that completely eliminates the need for centralized LSFD and its associated fronthaul overhead.
-Central to this architecture is the design of interference-suppressing local weights computed independently by each AP using only its local channel information.
+Central to this architecture is the design of interference-suppressing local weights, computed independently by each AP using only its local channel information.
 These weights are applied to data estimates before transmission, enabling the central processor to operate as a simple aggregator rather than a coordinator.
-This eliminates the need for any inter-AP coordination or channel information sharing with the CPU, reducing it to a low-complexity aggregator.
-This transition from a coordinated to a truly distributed architecture is the key enabler for the practical, large-scale deployment of cell-free networks.
+This eliminates the need for sharing any channel information with the CPU.
+This transition from a coordinated to a truly decentralized architecture is the key enabler for the practical, large-scale deployment.
 Adaptive Combining Strategy: We introduce generalized PFZF (G-PFZF) and generalized PWPFZF (G-PWPFZF).
-Unlike conventional PFZF and PWPFZF, where all APs must employ the same scheme, the generalized framework allows each AP to independently adapt its combining strategy.
-Specifically, under G-PFZF an AP may locally switch among PFZF, FZF, or MR; under G-PWPFZF, it may switch among PWPFZF, FZF, or MR.
-Thereby eliminating rigid threshold-based grouping and unlocking per-AP adaptability for improved scalability and performance.
-  • Novel Distributed Local Optimization for Pilot Partitioning: We also propose novel distributed optimization framework where each AP independently partitions pilots (and consequently, their associated UEs) into strong or weak groups based on a local sum spectral efficiency metric.
-This pilot-level optimization is solved using projected gradient ascent optimization and therefore maintaining low computational cost while maintaining functional equivalence with user-level partitioning.
-  • Closed-Form Spectral Efficiency Expressions: We derive closed-form expressions for the achievable spectral efficiency of the proposed system.
+Unlike the conventional PFZF and PWPFZF schemes, where all APs must employ the same scheme, the generalized framework allows each AP to independently adapt its combining strategy.
+Specifically, under G-PFZF an AP may locally switch among PFZF, FZF, or MR; under G-PWPFZF, it may switch among PWPFZF, FZF, or MR, thereby eliminating rigid threshold-based grouping and unlocking per-AP adaptability for improved scalability and performance.
+Novel Decentralized Local Optimization for Pilot Partitioning: We also propose novel decentralized optimization framework where each AP independently partitions pilots (and consequently, their associated UEs) into strong or weak groups based on a local sum SE metric.
+This pilot-level optimization is solved using projected gradient ascent optimization, maintaining low computational cost while maintaining functional equivalence with user-level partitioning.
+Closed-Form Spectral Efficiency Expressions: We derive closed-form expressions for the achievable SE of the proposed system.
 This analysis is essential due to the new statistical properties introduced by pilot-dependent processing and local weighting, and it provides critical insights for system design without the need for Monte Carlo simulation as in MMSE based schemes.
-  • Performance Analysis: Extensive numerical results demonstrate that our decentralized framework and adaptive grouping strategy achieve good performance with reduced computation cost with state-of-art baselines.
-The adaptive UE partitioning outperforms traditional fixed-threshold schemes, while the overall design significantly reduces fronthaul and computational costs in comparison to LSFD framewrok and fixed threshold partitioning, validating the approach as a practical solution for large-scale deployment.
+Performance Analysis: Extensive numerical results demonstrate that our decentralized framework and adaptive grouping strategy achieve good SE with reduced computation cost with the state-of-art baselines.
+The adaptive UE partitioning outperforms traditional fixed-threshold schemes, while the overall design significantly reduces fronthaul overhead and computational costs in comparison to the LSFD framework and fixed threshold partitioning, validating the approach as a practical solution for large-scale deployment.
 Organization: The remainder of the paper is organized as follows.
@@
 Section III presents the proposed combining schemes and derives their corresponding closed-form SE expressions.
-Section IV provides numerical simulations to evaluate performance.
+Section IV provides a rigorous performance evaluation.
 Finally, Section V concludes the paper and outlines potential directions for future work.
@@
 The sets of real and complex numbers are denoted by $\mathbb{R}$ and $\mathbb{C}$.
-The cardinality of a set $\mathcal{S}$ is denoted by $|\mathcal{S}|$.
 The expectation operator is denoted by $\mathbb{E}[\cdot]$.
@@
 II System Model
-The system consist of $T$ UEs with a single antenna and $M$ APs, where each AP has $A$ antennas, ensuring that $T\ll MA$.
+The system consist of $T$ UEs with a single antenna and $M$ APs, where each AP has $A$ antennas.
 Within the coverage area of interest, there is a uniform distribution of both APs and UEs, and APs work together to service the UEs using the same frequency and time resources.
-Every AP is linked to the CPU by fronthaul link, so that the network can operate together and process signals.
-Using uplink pilot transmissions, the system estimates the AP-UE channels under the time-division duplexing (TDD) regime.
-The wireless channel is represented using a block fading model, characterized by a coherence block of length $L_{c}$ symbols of which $L_{p}$ pilot symbols are designated for uplink pilot training.
+Every AP is linked to the CPU by a fronthaul link, so that the network can operate together and process signals.
+The wireless channel is represented using a block fading model, characterized by a coherence block of length $L_{c}$ symbols of which $L_{p}$ pilot symbols are designated for pilot training.
 The small-scale fading is represented by a Rayleigh fading vector $\textbf{h}_{mt}\in\mathbb{C}^{A\times 1}$, and the large-scale fading coefficient (LSFC), encompassing both path-loss and shadowing effects, is denoted by $\beta_{mt}$.
-The comprehensive channel vector $\textbf{g}_{mt}\in\mathbb{C}^{A\times 1}$ between the AP $m$ and the UE $t$, encompassing both small- and large-scale fading, is defined as $\mathbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}$.
-We suppose that $\textbf{h}_{mt}\sim\mathcal{N}_{\mathbb{C}}(0,\textbf{I}_{A})$, for all $m$ and $t$, is an independent and identically distributed complex Gaussian random vector, thus $\mathbf{g}_{mt}\sim\mathcal{N}_{\mathbb{C}}(0,\beta_{mt}\textbf{I}_{A})$.
+The channel vector between AP $m$ and UE $t$, encompassing both small- and large-scale fading, is defined as $\mathbf{g}_{mt}=\beta_{mt}^{1/2}\textbf{h}_{mt}\in\mathbb{C}^{A\times 1}$.
+We suppose that $\textbf{h}_{mt}\sim\mathcal{CN}(0,\textbf{I}_{A})$, for all $m$ and $t$, is an independent and identically distributed complex Gaussian random vector, thus $\mathbf{g}_{mt}\sim\mathcal{CN}(0,\beta_{mt}\textbf{I}_{A})$.
 II-A Channel Estimation
-A pilot sequence $\sqrt{L_{p}}\bm{\psi}_{i_{t}}\in\mathbb{C}^{L_{p}\times 1}$ is transmitted by each UE $t$ during the uplink training stage, with $|\bm{\psi}_{i_{t}}|^{2}=1$, where $i_{t}$ is the pilot index of the UE $t$.
-At the $m$-th AP, the received signal $\textbf{y}^{\text{pilot}}_{m}\in\mathbb{C}^{A\times L_{p}}$ is provided by:
-$\textbf{y}^{pilot}_{m}=\sum_{t=1}^{T}\sqrt{p^{p}_{t}L_{p}}\textbf{g}_{mt}\bm{\psi}_{i_{t}}^{{H}}+\textbf{N}_{m},$
+A pilot sequence $\sqrt{L_{p}}\bm{\psi}_{i_{t}}\in\mathbb{C}^{L_{p}\times 1}$ is transmitted by each UE $t$ during the uplink training stage, with $|\bm{\psi}_{i_{t}}|^{2}=1$, where $i_{t}$ is the pilot index of UE $t$.
+At the $m$-th AP, the received signal is:
+$\textbf{y}^{pilot}_{m}=\sum\nolimits_{t=1}^{T}\sqrt{p^{p}_{t}L_{p}}\textbf{g}_{mt}\bm{\psi}_{i_{t}}^{{H}}+\textbf{N}_{m},$
 where $\textbf{N}_{m}\in\mathbb{C}^{A\times L_{p}}$ represents the additive white Gaussian noise (AWGN) matrix characterized by independent and identically distributed complex Gaussian entries.
-Additionally, $p^{p}_{t}$ represents the normalized signal-to-noise power ratio for the UE $t$ during pilot transmission.
-The minimum mean square error (MMSE) estimate $\hat{\mathbf{g}}_{mt}\in\mathbb{C}^{A\times 1}$ of the true channel vector $\mathbf{g}_{mt}$ is provided by [1]:
-$\hat{\textbf{g}}_{mt}=c_{mt}\textbf{y}^{pilot}_{m}\bm{\psi}_{i_{t}},$
-where
-${c}_{mt}=\frac{\sqrt{p^{p}_{t}L_{p}}\beta_{mt}}{\sum_{k=1}^{T}p^{p}_{k}L_{p}\beta_{mk}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+1}.$
-The estimate $\hat{\mathbf{g}}_{mt}$ and estimated error $\tilde{\mathbf{g}}_{mt}={\mathbf{g}}_{mt}-\hat{\mathbf{g}}_{mt}$ are independent Gaussian with distributions $\hat{\mathbf{g}}_{mt}\sim\mathcal{N}_{\mathbb{C}}(0,\gamma_{mt}\textbf{I}_{A})$ and $\tilde{\mathbf{g}}_{mt}\sim\mathcal{N}_{\mathbb{C}}(0,(\beta_{mt}-\gamma_{mt})\textbf{I}_{A})$, where
-$\gamma_{mt}=\mathbb{E}\big\{|[\hat{\mathbf{g}}_{mt}]_{l}|^{2}\big\}=\left(\frac{p^{p}_{t}L_{p}\beta_{mt}^{2}}{\sum_{k=1}^{T}p^{p}_{k}L_{p}\beta_{mk}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+1}\right),$ (1)
-The matrix of estimated channels at the AP $m$ is denoted by $\mathbf{\hat{G}}_{m}=[\hat{\mathbf{g}}_{m1},\hat{\mathbf{g}}_{m2},\ldots,\hat{\mathbf{g}}_{mT}]\in\mathbb{C}^{A\times T}$.
+Additionally, $p^{p}_{t}$ represents the normalized pilot power for UE $t$ during pilot transmission.
+The MMSE estimate of the true channel vector $\mathbf{g}_{mt}$ is provided by [1]:
+$\hat{\textbf{g}}_{mt}=\sqrt{p^{p}_{t}L_{p}}\beta_{mt}\theta^{-1}_{mi_{t}}\textbf{y}^{pilot}_{m}\bm{\psi}_{i_{t}},$
+where $\theta_{mi_{t}}{=}\big(\sum_{k=1}^{T}p^{p}_{k}L_{p}\beta_{mk}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+1\big)$.
+The estimate $\hat{\mathbf{g}}_{mt}$ and estimated error $\tilde{\mathbf{g}}_{mt}{=}{\mathbf{g}}_{mt}{-}\hat{\mathbf{g}}_{mt}$ are independent Gaussian with distributions $\hat{\mathbf{g}}_{mt}{\sim}\mathcal{CN}(0,\gamma_{mt}\textbf{I}_{A})$ and $\tilde{\mathbf{g}}_{mt}{\sim}\mathcal{CN}(0,(\beta_{mt}{-}\gamma_{mt})\textbf{I}_{A})$, where
+$\gamma_{mt}=\mathbb{E}\big\{|[\hat{\mathbf{g}}_{mt}]_{l}|^{2}\big\}=p^{p}_{t}L_{p}\beta_{mt}^{2}\theta^{-1}_{mi_{t}},$ (1)
+Remark 1: The channel estimates of UEs sharing the same pilot sequence are linearly dependent. As a result, an AP cannot fully distinguish among these UEs, leading to residual pilot contamination.
+The matrix of estimated channels at AP $m$ is denoted by $\mathbf{\hat{G}}_{m}{=}[\hat{\mathbf{g}}_{m1},\hat{\mathbf{g}}_{m2},\ldots,\hat{\mathbf{g}}_{mT}]{\in}\mathbb{C}^{A\times T}$.
 Owing to pilot reuse among the UEs, the estimated channel vectors $\hat{\mathbf{g}}_{mt}$ and $\hat{\mathbf{g}}_{mk}$ corresponding to two UEs $t$ and $k$ that share the same pilot sequence become linearly dependent.
 As a result, the columns of $\mathbf{\hat{G}}_{m}$ exhibit linear dependence, rendering the matrix $\mathbf{\hat{G}}_{m}$ rank-deficient.
-The full rank matrix is designed by removing the linear dependent columns and by including only those columns that contain linearly independent channel estimates, representing one vector for each pilot sequence.
-Thus, the full rank matrix $\mathbf{\bar{G}}_{m}\in\mathbb{C}^{A\times L_{p}}$ is designed as:
-$\mathbf{\bar{G}}_{m}=\textbf{y}^{pilot}_{m}\bm{\Psi},$
-where $\bm{\Psi}=[\bm{\psi}_{1},\bm{\psi}_{2},\ldots,\bm{\psi}_{L_{p}}]\in\mathbb{C}^{L_{p}\times L_{p}}$. The channel estimate $\hat{\mathbf{g}}_{mt}$ is rewritten as:
+The full rank matrix is designed by removing the linearly dependent columns and by including only those columns that contain linearly independent channel estimates, representing one vector for each pilot sequence.
+Thus, the matrix $\mathbf{\bar{G}}_{m}$ is designed as:
+$\mathbf{\bar{G}}_{m}=\textbf{y}^{pilot}_{m}\bm{\Psi},\ \ \in\mathbb{C}^{A\times L_{p}}$
+where $\bm{\Psi}=[\bm{\psi}_{1},\bm{\psi}_{2},\ldots,\bm{\psi}_{L_{p}}]$. The channel estimate $\hat{\mathbf{g}}_{mt}$ is rewritten as:
 $\hat{\mathbf{g}}_{mt}=c_{mt}\mathbf{\bar{G}}_{m}\mathbf{e}_{i_{t}},$
@@
 II-B Uplink Transmission and Spectral Efficiency
-Let $x_{t}$ denote the unit power uplink data signal transmitted by the UE $t$, satisfying $\mathbb{E}{|x_{t}|^{2}}=1$.
-The uplink normalised signal-to-noise power ratio corresponding to the UE $t$ and additive white Gaussian noise are denoted by $p^{u}_{t}$ and $\textbf{n}_{m}\sim\mathcal{N}_{\mathbb{C}}(0,\textbf{I}_{A})$, respectively.
-The uplink signal $\textbf{y}^{u}_{m}\in\mathbb{C}^{A\times 1}$ received at the $m$-th AP is expressed as:
-$\textbf{y}^{u}_{m}=\sum_{t\in\mathcal{T}}\textbf{g}_{mt}\sqrt{p^{u}_{t}}x_{t}+\textbf{n}_{m}.$
+Let $x_{t}$ denote the unit power uplink data signal transmitted by UE $t$.
+The uplink normalised data power corresponding to UE $t$ and additive white Gaussian noise are denoted by $p^{u}_{t}$ and $\textbf{n}_{m}\sim\mathcal{N}_{\mathbb{C}}(0,\textbf{I}_{A})$, respectively.
+The uplink signal received at the $m$-th AP is:
+$\textbf{y}^{u}_{m}=\sum\nolimits_{t=1}^{T}\textbf{g}_{mt}\sqrt{p^{u}_{t}}x_{t}+\textbf{n}_{m}.$
 In decentralized data detection, unlike centralized approach where each AP sends received signal and channel estimates to CPU, each AP performs data detection for its served UEs using a local linear received combining vector.
@@
 These weights are applied to the locally detected data symbols before being sent to the CPU.
-This differs from LSFD, where local data estimates and channel estimates are transmitted to the CPU for centralized processing.
-The locally weighted data symbol $\hat{x}_{mt}$ for the UE $t$ at the AP $m$ is given by
+This differs from LSFD, where local data and channel estimates are transmitted to the CPU for centralized processing.
+The locally weighted data symbol $\hat{x}_{mt}$ for UE $t$ at AP $m$ is given by
 $\hat{x}_{mt}=a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{y}^{u}_{m},$ (2)
-where $\mathbf{v}_{m{i_{t}}}\in\mathbb{C}^{A\times 1}$ denotes the local combining vector for the UE $t$ at the AP $m$, and $a_{mt}$ is the corresponding local weight.
-These weighted local estimates are forwarded to the CPU, which aggregates them to obtain the final estimate of $x_{t}$:
-$\hat{x}_{t}=\sum_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{y}^{u}_{m}.$ (3)
+where $\mathbf{v}_{m{i_{t}}}\in\mathbb{C}^{A\times 1}$ denotes the local combining vector for UE $t$ at AP $m$, and $a_{mt}$ is the corresponding local weight.
+These weighted estimates are forwarded to the CPU, which aggregates them to obtain the final estimate of $x_{t}$:
+$\hat{x}_{t}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{y}^{u}_{m}.$ (3)
 After expanding $\textbf{y}^{u}_{m}$, the (3) can also be rewritten as
-$\hat{x}_{t}=\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}+\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\tilde{\textbf{g}}_{mt}+\sum_{k=1,k\neq t}^{T}\sum_{m=1}^{M}\sqrt{p^{u}_{k}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk}x_{k}+\sum_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m}.$ (4)
-Here, the first term is the desired signal for the UE $t$, the second term is interference term due to imperfect channel estimation, the third term is interference term due to all other UEs and last term the noise.
-Based on the decomposition in (4), the achievable uplink spectral efficiency is obtained using the bounding technique and the Shannon capacity lower bound [17, 18, 16] is given by Theorem 1.
+$\begin{split}\hat{x}_{t}&=\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}+\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\tilde{\textbf{g}}_{mt}\\ &+\sum_{\begin{subarray}{c}k=1\\ k\neq t\end{subarray}}^{T}\sum_{m=1}^{M}\sqrt{p^{u}_{k}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk}x_{k}+\sum_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m}.\end{split}$ (4)
+Here, the first term is the desired signal for UE $t$, the second term is interference term due to imperfect channel estimation, the third term is interference term due to all other UEs and last term is the noise.
+Based on the decomposition in (4), the achievable uplink spectral efficiency is obtained using the bounding technique and the corresponding Shannon capacity lower bound [17, 18, 16] is given by Theorem 1.
 Theorem 1.
-[16, 13] : A lower bound on the uplink ergodic capacity for UE $t$ is given by
+[16, 13] : A lower bound on the uplink ergodic SE for UE $t$ is given by
 $\textsc{SE}^{u}_{t}=L_{u}\log_{2}\big(1+\textsc{SINR}_{t}\big),$ (5)
-where $L_{u}$ is $\Big(\frac{1-\frac{L_{p}}{L_{c}}}{2}\Big)$ and the SINR for the UE $t$ is defined by:
-[…]
+where $L_{u}=\Big(\frac{1-\frac{L_{p}}{L_{c}}}{2}\Big)$ and $\textsc{SINR}_{t}$ is
+$\textsc{SINR}_{t}=\frac{p^{u}_{t}|\textsc{DS}_{t}|^{2}}{\!\!\sum\limits_{k=1}^{T}\!\!p^{u}_{k}\mathbb{E}\left\{|\textsc{UI}_{tk}|^{2}\right\}-p^{u}_{t}|\textsc{DS}_{t}|^{2}+\mathbb{E}\left\{|\textsc{GN}_{t}|^{2}\right\}},$ (6)
 where,
-$\begin{split}&\textsc{DS}_{t}=\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbb{E}\left\{\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}\right\}\\ &\textsc{BU}_{t}=\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}(\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mt}-\mathbb{E}\left\{\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mt}\right\})\\ &\textsc{UI}_{tk}=\textsc{PC}_{tk}=\sum_{m=1}^{M}\sqrt{p^{u}_{k}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk}\\ &\textsc{GN}_{t}=\sum_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m},\end{split}$
-where $\mathcal{P}_{i_{t}}$ represents the subset of UEs sharing the pilot $i_{t}$. Also, the SE lower bound, in (5), is valid regardless of the combining scheme used.
+$\begin{split}&\textsc{DS}_{t}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbb{E}\left\{\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}\right\},\\ &\textsc{UI}_{tk}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk},\\ &\textsc{GN}_{t}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m}.\end{split}$
+The SE lower bound, in (5), is valid regardless of the combining scheme used.
 III Adaptive Combining and Local Weights Analysis
-In this section, we analyze the proposed generalized combining schemes, G-PFZF and G-PWPFZF, from both performance and computational cost perspectives.
-We first derive closed-form expressions for the uplink SE and local combining weights for both schemes.
-These expressions provide analytical insights into how the proposed framework improves performance while maintain scalability.
-Furthermore, we evaluate the computational and fronthauling costs associated with each scheme, highlighting the significant reductions achieved through decentralized decoding system design.
-In [13], the combining vectors for any UE $t$ is dependent on $c_{mt}$, which is unique to each UE.
-As a result, generating a distinct ZF vector for every UE entails a number of multiplications that scales linearly with the total number of UEs.
-To reduce computational cost, we seek an alternative formulation in which the number of multiplications scales with the number of orthogonal pilot sequences, rather than with the number of UEs.
+In this section, we first derive closed-form expressions for the uplink SE and local combining weights for both G-PFZF and G-PWPFZF schemes.
+We also formulate an optimization problem to partition the pilots into strong and weak.
 III-A Generalized Partial Full-Pilot Zero Forcing Combining Scheme
-In the proposed G-PFZF combining scheme, each AP independently designs a combining vector for each of its pilots and pilot-sharing UEs use the same pilot, as APs are not able to distinguish among individual UEs sharing the same pilot.
+In [13], the combining vectors for UE $t$ is dependent on $\beta_{mt}$, which is unique to each UE.
+As a result, generating a distinct combining vector for every UE entails a number of multiplications that scales linearly with the total number of UEs.
+To reduce computational cost, we propose G-PFZF combining scheme, where each AP independently designs a combining vector for each of its pilots and pilot-sharing UEs use the same pilot, as APs are not able to distinguish among individual UEs sharing the same pilot.
 The design of combining vectors depends on whether a pilot is strong or weak.
 Thus, each AP categorizes the pilots into two groups, strong and weak based on the local sum SE optimization.
-Unlike traditional PFZF, G-PFZF does not forces the APs to select atleast one pilot as strong and allows the APs to choose the combining schemes among PFZF or MR or FZF.
-For the AP $m$, a set of strong pilots is denoted by $\mathcal{S}_{m}$ and a complementary set of weak UEs as $\mathcal{W}_{m}$.
-Let $L_{\mathcal{S}_{m}}$ denote the number of distinct strong pilots used in the set $\mathcal{S}_{m}$, and define the corresponding set of pilot indices as $\mathcal{R}_{\mathcal{S}_{m}}=\{r_{m,1},r_{m,2},\dots,r_{m,L_{\mathcal{S}_{m}}}\}$.
+Unlike traditional PFZF, G-PFZF does not force the APs to select atleast one pilot as strong and allows the APs to choose the combining schemes among PFZF or MR or FZF.
+Remark 2: In our proposed combining schemes, a strong pilot indicates that a ZF combining vector is assigned to that pilot (and its corresponding UEs), while a weak pilot is assigned an MR combining vector.
+For AP $m$, a set of strong pilots is denoted by $\mathcal{S}_{m}$ and a complementary set of weak UEs as $\mathcal{W}_{m}$.
+Let $L_{\mathcal{S}_{m}}$ denote the number of distinct strong pilots used in set $\mathcal{S}_{m}$, and define the corresponding set of pilot indices as $\mathcal{R}_{\mathcal{S}_{m}}{=}\{r_{m,1},r_{m,2},{\dots},r_{m,L_{\mathcal{S}_{m}}}\}$.
 The matrix $\mathbf{\bar{G}}_{m}{\in}\mathbb{C}^{A\times L_{p}}$ contains the estimated channel vectors corresponding to all $L_{p}$ orthogonal pilot sequences, whereas the strong pilots utilize only $L_{\mathcal{S}_{m}}$ of them.
 To isolate the relevant columns of $\mathbf{\bar{G}}_{m}$, we define the selection matrix $\mathbf{E}_{\mathcal{S}_{m}}{=}[\mathbf{e}_{r_{m,1}},{\dots},\mathbf{e}_{r_{m,L_{\mathcal{S}_{m}}}}]{\in}\mathbb{C}^{L_{p}\times L_{\mathcal{S}_{m}}}$, where $\mathbf{e}_{r}$ denotes the $r$-th column of the identity matrix $\mathbf{I}_{L_{p}}$.
-The resulting product $\mathbf{\bar{G}}_{m}\mathbf{E}_{\mathcal{S}_{m}}$ yields a full-rank matrix containing only the $L_{\mathcal{S}_{m}}$ linearly independent columns corresponding to the strong pilots.
-For a given pilot $i\in\mathcal{S}_{m}$, let $j_{mi}\in\{1,2,\dots,L_{\mathcal{S}_{m}}\}$ denote the index of the pilot $i$ within $\mathcal{R}_{\mathcal{S}_{m}}$, and define the vector $\bm{\varepsilon}_{j_{mi}}\in\mathbb{C}^{L_{\mathcal{S}_{m}}}$ as the $j_{mi}$-th column of the identity matrix $\mathbf{I}_{L_{\mathcal{S}_{m}}}$, such that $\mathbf{E}_{\mathcal{S}_{m}}\bm{\varepsilon}_{j_{mi}}=\mathbf{e}_{i}$.
-As discussed earlier, the effective channel matrix corresponding to the strong-pilot UEs is represented as $\mathbf{\bar{G}}_{m}\mathbf{E}_{\mathcal{S}_{m}}$.
+The resulting product $\mathbf{\bar{G}}_{m}\mathbf{E}_{\mathcal{S}_{m}}$ yields a full-rank matrix containing only $L_{\mathcal{S}_{m}}$ linearly independent columns corresponding to the strong pilots.
+For a given pilot $i{\in}\mathcal{S}_{m}$, let $j_{mi}{\in}\{1,2,\dots,L_{\mathcal{S}_{m}}\}$ denote the index of pilot $i$ within $\mathcal{R}_{\mathcal{S}_{m}}$, and define the vector $\bm{\varepsilon}_{j_{mi}}{\in}\mathbb{C}^{L_{\mathcal{S}_{m}}}$ as the $j_{mi}$-th column of the identity matrix $\mathbf{I}_{L_{\mathcal{S}_{m}}}$, such that $\mathbf{E}_{\mathcal{S}_{m}}\bm{\varepsilon}_{j_{mi}}{=}\mathbf{e}_{i}$.
 Following the zero-forcing principle, the pseudo-inverse of this matrix can be expressed as $(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}})^{-1}$.
@@
 Similarly, the MR combining vector for pilot $i\in\mathcal{W}_{m}$ at AP $m$ corresponds directly to the channel vector associated with that pilot and is expressed as:
-$\mathbf{v}_{mi}^{\text{MR}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{e}_{i}}{\sqrt{\mathbb{E}\Big\{\Big\|\mathbf{\bar{G}}_{m}\mathbf{e}_{i}\Big\|^{2}\Big\}}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{e}_{i}}{\sqrt{A\theta_{mi}}},$ (8)
-where $\theta_{mi}=\mathbb{E}\big\{|[\mathbf{\bar{G}}_{m}\mathbf{e}_{i}]_{l}|^{2}\big\}=\big(\sum_{k=1}^{T}p^{p}_{k}L_{p}\beta_{mk}\left|\bm{\psi}^{H}_{i}\bm{\psi}_{i_{k}}\right|^{2}+1\big)$.
-Following [19, 20, 21], the normalization term in (7), for $L_{\mathcal{S}_{m}}\times L_{\mathcal{S}_{m}}$ complex Wishart matrix, $\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}$, with $A$ degree of freedom satisfying $A-1\geq L_{\mathcal{S}_{m}}$, can be rewritten as
-$\sqrt{\mathbb{E}\Big\{\bm{\varepsilon}^{H}_{j_{mi}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\bm{\varepsilon}_{j_{mi}}\Big\}}=\sqrt{\mathbb{E}\Big\{\big[(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\big]_{j_{mi},j_{mi}}\Big\}}=\frac{1}{\sqrt{(A-L_{\mathcal{S}_{m}})\theta_{mi}}},$ (9)
-Also, the G-PFZF scheme suppresses the interference among the UEs whose pilot lies in $\mathcal{S}_{m}$ by sacrificing the $L_{\mathcal{S}_{m}}$ degree of freedom from the total $A$ degrees to boost the desired signal.
+$\mathbf{v}_{mi}^{\text{MR}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{e}_{i}}{\sqrt{\mathbb{E}\big\{\big\|\mathbf{\bar{G}}_{m}\mathbf{e}_{i}\big\|^{2}\big\}}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{e}_{i}}{\sqrt{A\theta_{mi}}}.$ (8)
+Following [19, 20, 21], the normalization term in (7), for $L_{\mathcal{S}_{m}}\times L_{\mathcal{S}_{m}}$ complex Wishart matrix ($\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}$) with degree of freedom satisfying $A\geq L_{\mathcal{S}_{m}}+1$, can be rewritten as
+${\mathbb{E}\Big\{\bm{\varepsilon}^{H}_{j_{mi}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\bm{\varepsilon}_{j_{mi}}\Big\}}=\frac{1}{{(A-L_{\mathcal{S}_{m}})\theta_{mi}}},$ (9)
+Also, the G-PFZF scheme suppresses the interference among the UEs whose pilot lies in $\mathcal{S}_{m}$ by sacrificing $L_{\mathcal{S}_{m}}$ degrees of freedom from total $A$ degrees to boost the desired signal.
 Therefore, for any UE $t$ such that $i_{t}\in\mathcal{S}_{m}$
 $(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}=\begin{cases}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t},\end{cases}$ (10)
-$\begin{split}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}})\gamma_{mk}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},&\text{if }i_{k}\notin\mathcal{S}_{m}\end{cases}\end{split}$ (11)
+$\begin{split}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}})\gamma_{mk}\ \ \text{if }i_{k}=i_{t},\\ 0,\ \ \text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},\ \ \text{if }i_{k}\notin\mathcal{S}_{m}\end{cases}\end{split}$ (11)
 Similarly, for any UE $t$ such that $i_{t}\in\mathcal{W}_{m}$
-$\mathbb{E}\{(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{g}_{mk}}\}=\begin{cases}\sqrt{A\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}.\end{cases}$ (12)
-$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\mathbb{E}\Big\{\Big|(\frac{1}{\sqrt{A\theta_{mi_{t}}}c_{mt}}\hat{\mathbf{g}}^{H}_{mt}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A+1)\gamma_{mk}&\text{if }i_{k}=i_{t},\\ \gamma_{mk},&\text{if }i_{k}\neq i_{t},\end{cases}$ (13)
+$\mathbb{E}\{(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}{\mathbf{\hat{g}}_{mk}}\}=\begin{cases}\sqrt{A\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}.\end{cases}$ (12)
+$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A+1)\gamma_{mk}&\text{if }i_{k}=i_{t},\\ \gamma_{mk},&\text{if }i_{k}\neq i_{t},\end{cases}$ (13)
 Note 1: From (11) and (13), it is evident that local ZF achieves superior interference suppression compared to MR, since each AP can eliminate interference among UEs with pilots grouped as strong.
 However, this interference mitigation comes at the cost of $L_{\mathcal{S}_{m}}$ degrees of freedom, which reduces the array gain available for desired signal enhancement.
-By using the combining vectors in (7) and (8), the closed-form expression of achievable SE for G-PFZF combining can be computed using Theorem 1 with the corresponding SINR given by:
+By using the combining vectors in (7) and (8), the closed-form expression of achievable SE for G-PFZF combining can be computed using Theorem 1 with the corresponding SINR given by
 […]
-where $\delta_{mi_{t}}$ is defined as:
+where $\mathcal{P}_{i_{t}}$ represents the subset of UEs sharing pilot $i_{t}$ and $\delta_{mi_{t}}$ is defined as:
 $\delta_{mi_{t}}=\begin{cases}1,&\text{if }\ i_{t}\in\mathcal{S}_{m},\\ 0,&\text{otherwise}.\end{cases}$ (15)
-The derivation of SINR expression is given in Appendix -A.
+The derivation of the SINR expression is given in Appendix -A.
 Local Weight Design for G-PFZF Scheme: A fundamental limitation of the conventional LSFD architecture is its reliance on centralized optimization [13].
@@
 In this work, we break this dependency by introducing a novel decentralized decoding architecture where each AP $m$ independently computes a local interference-suppressing weight $a_{mt}$ for each UE $t$, based solely on its local channel information.
-Since these weights are derived from the LSFCs, they evolve much slower than the small-scale fading components, ensuring long-term stability and minimal operational overhead.
+Since these weights are derived from the LSFCs, they evolve much slowly than the small-scale fading components, ensuring long-term stability and minimal operational overhead.
 The primary objectives of this local weight design are twofold: (1) to maximize the contribution of the local estimate to the final SINR at the CPU, and (2) to actively mitigate the dominant interference components, specifically, pilot contamination and inter-user interference, at their source, prior to transmission.
 This approach eliminates the need for the APs to send the channel estimates to the CPU and the CPU to perform any complex optimization, reducing its role to that of a simple aggregator that sums the pre-weighted estimates from all APs.
-The main idea behind designing decoding weights is that it should amplify the data signal received from AP with strong local desired signal component and suppress those dominated by interference.
+The main idea behind designing decoding weights is that it should amplify the data signal received from the APs with strong local desired signal component and suppress those dominated by interference.
 In LSFD design, weights are computed through coordination among the APs, specifically the coherent interference terms.
 In contrast, the proposed decentralized framework eliminates inter-AP coordination, requiring each AP to determine its weight independently.
-Accordingly, an AP with a higher local SINR is given a larger weight than one with a lower SINR in order to enhance the global SINR.
+Accordingly, an AP with a higher local SINR is given a larger weight than the one with a lower SINR in order to enhance the global SINR.
 For simplicity, the local weight is chosen to be equal to the local SINR.
-The closed-form expression of the local SINR for the UE $t$ at the AP $m$ can be obtained from (2) by expanding it in the same manner as in (4) and following the derivation procedure outlined in Appendix -A.
-Accordingly, the closed-form of the local SINR or local weight with respect to the AP $m$ of the UE $t$ is given by:
+The closed-form expression of the local SINR for UE $t$ at AP $m$ can be obtained from (2) by expanding it in the same manner as in (4) and following the derivation procedure outlined in Appendix -A.
+Accordingly, the closed-form of the local SINR or local weight with respect to AP $m$ of UE $t$ is given by:
 $a_{mt}=\text{SINR}^{\text{G-PFZF}}_{mt}=\frac{p^{{u}}_{t}(b_{tt}^{m})^{2}}{\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{{u}}_{k}(b_{kt}^{m})^{2}+W_{mt}}$ (16)
@@
 ${W}_{{m}t}=\sum\limits_{k=1}^{T}p^{{u}}_{k}(\beta_{{m}k}-\delta_{{m}i_{t}}\delta_{{m}i_{k}}\gamma_{{m}k})+1.$
-The numerator represents the desired signal strength for the desired UE $t$ at the AP $m$, capturing the gain achieved through the chosen combining strategy.
-The denominator is designed to penalize the weight by two main sources of impairment: the coherent interference from pilot contamination and the combined power of non-coherent interference and noise.
-By applying these weight to its local estimate, each AP independently produces a ”soft” decision that is already pre-optimized for local interference suppression.
+The numerator represents the desired signal strength for the desired UE $t$ at AP $m$, capturing the gain achieved through the chosen combining strategy.
+The denominator is designed to penalize the weight by two main sources of impairment: the coherent interference from pilot contamination and the combined power of non-coherent interference due to channel estimation error and noise.
+By applying these weight to its local estimate, each AP independently produces a soft decision that is already pre-optimized for local interference suppression.
 This design ensures that the signals forwarded to the CPU require no further large-scale fading decoding, thereby completely eliminating the associated fronthaul overhead for channel statistics and the computational cost of centralized weight calculation.
@@
 This classification dictates the appropriate combining strategy, G-PFZF for strong pilots or MR for weak ones, and crucially, generalizes to pure FZF or pure MR should the optimization deem them to be optimal.
-The foundation of this optimization is the local SINR expression given in (16).
-As $L_{\mathcal{S}m}=\sum_{i=1}^{L_{p}}\delta_{mi}$ and substituting $L_{\mathcal{S}m}$ into (16) makes the dependency explicit:
+The basis of this optimization is the local SINR expression given in (16).
+As $L_{\mathcal{S}m}{=}\sum_{i=1}^{L_{p}}\delta_{mi}$ and substituting $L_{\mathcal{S}m}$ into (16) makes the dependency explicit as given in (17).
 […]
 This expression reveals a fundamental trade-off governed by the size of the strong pilot set, $L_{\mathcal{S}m}$.
-Allocating a pilot to the strong group (i.e., setting $\delta_{mi}=1$) employs ZF to suppress the non-coherent interference from that pilot’s UEs as well as other strong pilots UEs.
+Allocating a pilot to the strong group ( $\delta_{mi}{=}1$) employs ZF to suppress the non-coherent interference from that pilot’s UEs as well as other strong pilots UEs.
 However, this comes at the cost of a reduced array gain, $(A-L_{\mathcal{S}_{m}})$, for all UEs served by the AP, which diminishes the desired signal power.
@@
 This inherent trade-off between maximizing signal strength and minimizing interference makes the grouping problem non-trivial and renders a fixed, network-wide threshold highly suboptimal.
-To navigate the above mentioned trade-off, the AP $m$ independently solves a local optimization problem.
-The objective is to select the binary assignment variables $\delta_{mi}\in\{0,1\}$ for all pilots that maximize the sum of the local SEs for all UEs, as defined by
+To navigate the above mentioned trade-off, AP $m$ independently solves a local optimization problem, which is formulated as
 $\underset{\bm{\delta}_{m}}{\max~}$ $\ \sum\limits_{t=1}^{T}\log_{2}\left(1+\frac{S_{mt}}{I_{mt}}\right),$ (18a)
@@
 To develop a more tractable formulation, we begin by relaxing the binary constraint (18b) into a continuous one:
-$0\leq\delta_{mi}\leq 1\ \forall i.$ (19)
-This relaxation transforms the original MINLP into a Non-Linear Program (NLP).
-To ensure that the solutions of this new NLP satisfy the original binary requirement ($\delta_{mi}\in{0,1}$), we introduce the following complementary constraint for each $i$:
-$\delta_{mi}-\delta^{2}_{mi}\leq 0.$ (20)
+$0\leq\delta_{mi}\leq 1,\ \forall i.$ (19)
+To ensure that the solutions of this new NLP satisfy the original binary requirement ($\delta_{mi}\in{0,1}$), we introduce the following complementary constraint:
+$\delta_{mi}-\delta^{2}_{mi}\leq 0,\ \forall i.$ (20)
 The combined constraints (19) and (20) are equivalent to the original binary constraint (18b), as the only values between 0 and 1 that satisfy $\delta_{mi}(1-\delta_{mi})\leq 0$ are precisely 0 and 1.
 Thus, the problem can be written as
-$\underset{\bm{\delta}_{m}}{\max~}$ $\ \sum\limits_{t=1}^{T}\log(1+\text{SINR}^{\text{G-PFZF}}_{mt}),$ (21a)
+$\underset{\bm{\delta}_{m}}{\max~}$ $\ \sum\limits_{t=1}^{T}\log_{2}\left(1+\frac{S_{mt}}{I_{mt}}\right),$ (21a)
 […]
-This problem can be solved using non-linear solvers or the successive convex approximation optimization method.
-Both of these are computationally expensive for large network deployment.
+This problem can be solved using non-convex solvers or the successive convex approximation optimization method.
+Both of these are computationally expensive for large network deployment as non-convex solvers.
 Therefore, to solve this problem with low computation cost, we employ gradient based proximal gradient ascent (PGA) method [22, 23].
 We first reformulate the problem to be unconstrained except for the box constraint (19).
-This is achieved by moving the other constraints into the objective function via a penalty method.
 The reformulated problem is:
-$\underset{\bm{\delta}_{m}}{\max~}$ $f(\bm{\delta}_{m}){=}\sum\limits_{t=1}^{T}\log_{2}(1+\text{SINR}^{\text{G-PFZF}}_{mt})-\chi\Bigg(\!\lambda_{1}\!\sum_{i=1}^{L_{p}}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)^{2}+\lambda_{2}\max\!\Big(0,\sum_{i=1}^{L_{p}}\delta_{mi}-A+1\Big)^{2}\!\Bigg),$ (22a)
+$\underset{\bm{\delta}_{m}}{\max~}$ $f(\bm{\delta}_{m}),$ (22a)
 […]
+where
+$\begin{split}f(\bm{\delta}_{m})&=\!\!\sum\limits_{t=1}^{T}\log_{2}(1+\frac{S_{mt}}{I_{mt}})-\chi\Bigg(\!\lambda_{1}\!\sum_{i=1}^{L_{p}}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)^{2}+\lambda_{2}\max\!\Big(0,\sum_{i=1}^{L_{p}}\delta_{mi}-A+1\Big)^{2}\Bigg),\end{split}$
 where $\chi$ is a penalty parameter, and $\lambda_{1}$ and $\lambda_{2}$ are penalty weights.
@@
 We initialize $\bm{\delta}_{m}$ and update it along with the gradient ascent direction of $f(\bm{{\delta}}^{(j)}_{m})$ with step size $\alpha$, followed by projection onto $[0,1]$ to satisfy the box constraint (19):
-$\bm{{\delta}}^{(j+1)}_{m}=\text{proj}_{[0,1]}\!\left(\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)=\min\!\left(1,\max\!\left(0,\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)\right).$ (23)
-The components of gradient $\nabla f(\bm{\delta}_{m})=\Big[\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{m1}},\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{m2}},\dots,\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mL_{p}}}\Big]^{\intercal}$ can be written as
+$\begin{split}\bm{{\delta}}^{(j+1)}_{m}&=\text{proj}_{[0,1]}\!\left(\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)=\min\!\left(1,\max\!\left(0,\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)\right).\end{split}$ (23)
+The components of gradient $\nabla f(\bm{\delta}_{m})$ can be written as
 $\begin{split}\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}=&\sum_{t=1}^{T}\frac{I_{mt}}{I_{mt}+S_{mt}}\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}-\chi\Bigg(\!2\lambda_{1}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)\Big(1-2\delta_{mi}\Big)\\ &+2\lambda_{2}\max\!\Big(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\Big)\!\Bigg),\end{split}$ (24)
@@
 $\frac{\partial S_{mt}}{\partial\delta_{mi}}=\begin{cases}-p^{u}_{t}\Big(1+\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\delta_{mj}\Big)\gamma_{mt},&\text{if }\ i_{t}=i,\\ -p^{u}_{t}\delta_{mi_{t}}\gamma_{mt},&\text{if }\ i_{t}\neq i,\end{cases}$ (25)
-$\frac{\partial I_{mt}}{\partial\delta_{mi}}=\begin{cases}-\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{u}_{k}\Big(1+\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\delta_{mj}\Big)\gamma_{m}-\sum\limits_{k\in\mathcal{P}_{i_{t}}}p^{u}_{k}\gamma_{mk},&\text{if }\ i_{t}=i,\\ -\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{u}_{k}\delta_{mi_{t}}\gamma_{mk}-\sum\limits_{k\in\mathcal{P}_{i}}p^{u}_{k}\delta_{mi_{t}}\gamma_{mk},&\text{if }\ i_{t}\neq i,\end{cases}$ (26)
+$\frac{\partial I_{mt}}{\partial\delta_{mi}}=\begin{cases}-\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!p^{u}_{k}\ \Big(1+\!\!\!\!\!\!\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\!\!\!\!\!\!\!\delta_{mj}\Big)\gamma_{m}-\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}}\!\!\!\!p^{u}_{k}\gamma_{mk},\ \ \text{if }i_{t}=i,\\ \vskip 5.69054pt-\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!p^{u}_{k}\delta_{mi_{t}}\gamma_{mk}-\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}}\!\!\!\!p^{u}_{k}\delta_{mi_{t}}\gamma_{mk},\ \ \text{if }i_{t}\neq i,\end{cases}$ (26)
 1:for all $m\in\mathcal{M}$ independently do
@@
 4:   repeat
-5:    Compute $\bm{{\delta}}^{(j+1)}_{m}$ using (23)
+5:    Compute:
+$\bm{{\delta}}^{(j+1)}_{m}{=}\min\!\big(1,\max\!\big(0,\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\big)\big)$
 6:    Update $j=j+1$
@@
 Algorithm 1 Proximal Gradient Ascent (PGA) Method
-Convergence and Computation Cost Analysis: Since the feasible set defined by the box constraint (19) is bounded and the gradient function $\nabla f(\bm{\delta}_{m})$ is Lipschitz continuous (proof is given in Appendix -B) with constant $L>0$, one should choose $\alpha\in(0,1/L]$ to ensure that the gradient step remains within the stability region of the PGA method [22, 23].
+Convergence and Computation Cost Analysis: Since the feasible set defined by the box constraint (19) is bounded and the gradient function $\nabla f(\bm{\delta}_{m})$ is Lipschitz continuous (proved in Appendix -B) with constant $L>0$, one should choose $\alpha\in(0,1/L]$ to ensure the convergence of the PGA method [22, 23].
 In numerical simulations, small step size parameter $\alpha$ has been observed to achieve convergence for Algorithm 1.
@@
 Since we have only $L_{p}$ variables, it converges in a few iterations only.
-Note 2: To reduce the computational cost, we use pilot-based grouping instead of UE-based grouping, as discussed in [19, 13].
+Note 2: To reduce the computational cost, we use pilot-based grouping instead of UE-based grouping as in [19, 13].
 In UE-based grouping, each AP would have $T$ binary variables (one per UE), leading to $TM$ total variables across the network.
@@
 III-B Generalized Protected Weak Partial Full-Pilot Zero Forcing Combining Scheme
-In the G-PFZF scheme, the combiner for the UEs assigned to a strong pilot is designed only to suppress interference only from other UEs that are also on strong pilots.
+In the G-PFZF scheme, the combiner for the UEs assigned to a strong pilot is designed only to suppress interference from other UEs that are also on strong pilots.
 This intra-group suppression leaves UEs on weak pilots vulnerable to dominant interference from the strong UEs.
 The proposed G-PWPFZF scheme introduces a more comprehensive, protective strategy.
-In this framework, the combiners for all UEs, whether assigned to a strong or weak pilot, are designed to actively suppress interference from all UEs that use a strong pilot.
+In this framework, the combiners for all UEs, whether assigned to a strong or weak pilot, are designed to actively suppress interference from all UEs that have a strong pilot.
 This is achieved by projecting the combining vectors for weak UEs onto the orthogonal complement of the subspace spanned by the strong UEs’ effective channels.
-Consequently, the G-PWPFZF scheme provides universal protection against the most significant sources of interference, dramatically improving performance for far users on weak pilots, at cost of higher computation cost of designing the vector for weak pilot UEs.
-In the G-PWPFZF scheme, the strong pilots are assigned with the vectors in the same way as in (7).
-The combining vector allocated to the pilot $i_{t}\in\mathcal{W}_{m}$ corresponding to the AP $m$ is simply the normalized MR combining vector, as in (8), with a projection matrix and is given by:
-${\mathbf{v}}^{\text{PMR}}_{mi_{t}}=\frac{1}{\sqrt{(A-L_{\mathcal{S}_{m}})\theta_{mi_{t}}}}\mathbf{B}_{m}\mathbf{\bar{G}}_{m}e_{i_{t}},$ (27)
-where $\mathbf{B}_{m}$ is a projection matrix that projects the received signal onto the orthogonal complement of the subspace spanned by the effective channels of all strong pilots.
-It is defined as [19]:
+Consequently, the G-PWPFZF scheme provides universal protection against the most significant sources of interference, improving performance for users on weak pilots, at the cost of higher computation in comparison to G-PFZF scheme due to projection matrix computation at the AP.
+In the G-PWPFZF scheme, the strong pilot $i{\in}\mathcal{S}_{m}$ are assigned with the combining vectors ${\mathbf{v}}^{\text{LFZ}}_{mi}$ as in (7).
+The combining vector allocated to pilot $i{\in}\mathcal{W}_{m}$ at AP $m$ is the projected MR (PMR) combining vector, and is given by:
+${\mathbf{v}}^{\text{PMR}}_{mi}=\frac{1}{\sqrt{(A-L_{\mathcal{S}_{m}})\theta_{mi}}}\mathbf{B}_{m}\mathbf{\bar{G}}_{m}e_{i},$ (27)
+where $\mathbf{B}_{m}$ is a matrix that projects the received signal onto the orthogonal complement of the subspace spanned by the channels of all strong pilots. It is defined as [19]:
 $\mathbf{B}_{m}=\mathbf{I}_{A}-\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}})^{-1}\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m},$
 This projection ensures that the combiner for a weak pilot is orthogonal to the channel estimates of all strong-pilot UEs, thereby nullifying the dominant interference from that group.
-The expected value of the effective channel gain for the UE $t$ such that $i_{t}\in\mathcal{W}_{m}$
-$\mathbb{E}\{(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}_{mk}}\}=\begin{cases}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}.\end{cases}$ (28)
-Furthermore, the second moment of the effective channel gain is given by:
-$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}}+1)\gamma_{mk}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},&\text{if }i_{k}\neq i_{t},\end{cases}$ (29)
+The expected value of the effective channel gain for UE $t$ such that $i_{t}\in\mathcal{W}_{m}$
+$\mathbb{E}\{(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}{\mathbf{\hat{g}}_{mk}}\}=\begin{cases}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}.\end{cases}$ (28)
+Furthermore
+$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}}+1)\gamma_{mk}\ \ \text{if }i_{k}=i_{t},\\ 0,\ \ \text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},\ \ \text{if }i_{k}\neq i_{t},\end{cases}$ (29)
 These equations confirm that the projected MR (PMR) combiner successfully nullifies interference from all UEs sharing strong pilots $i_{k}\in\mathcal{S}_{m}$.
 This design effectively mitigates the most significant intra-network interference, enhancing the performance of the UEs with weaker channel conditions.
-By using the combining vector (7) for strong pilot and (27), the closed-form expression for achievable SE for G-PWPFZF combining, as in Theorem 1, can be obtained with the SINR given by:
+By using the combining vector (7) for strong pilot and (27), the closed-form expression for achievable SE for G-PWPFZF combining, as in Theorem 1, can be obtained with the SINR given by
 […]
-The derivation of SINR expression is given in Appendix -C.
-Local Weights Design for G-PWPFZF Scheme: The design of local weights for the G-PWPFZF scheme follows a similar design to that of the G-PFZF scheme, aiming to maximize the local contribution to the final SINR while suppressing interference in a fully distributed manner.
-Here too the local weights are same as the local SINR.
+The derivation of the SINR expression is given in Appendix -C.
+Local Weights Design for G-PWPFZF Scheme: The design of local weights for the G-PWPFZF scheme follows a similar design to that of the G-PFZF scheme, aiming to maximize the local contribution to the final SINR while suppressing interference in a fully decentralized manner.
+Here too the local weights are the same as the local SINR.
 The closed-form expression of the local SINR for UE $t$ at AP $m$ can be obtained from (2) by expanding it in the same manner as in (4) and following the derivation procedure outlined in Appendix -C.
@@
 The sole distinction lies in the expression for the local SINR, which forms the foundation of the utility function $f(\bm{\delta}_{m})$ for pilot-grouping.
-The local SINR for the UE $t$ at the AP $m$ for G-PWPFZF scheme, given in (31) after substituting $L_{\mathcal{S}m}=\sum_{i=1}^{L_{p}}\delta_{mi}$, can be rewritten as:
-[…]
+The local SINR for UE $t$ at AP $m$ for G-PWPFZF scheme is given in (31).
+After substituting $L_{\mathcal{S}m}=\sum_{i=1}^{L_{p}}\delta_{mi}$ in (31):
+$S_{mt}=p^{u}_{t}\Big(A-\sum\limits_{i=1}^{L_{p}}\delta_{mi}\Big)\gamma_{mt}$
+$I_{mt}=\!\!\!\!\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!p^{u}_{k}\Big(\!A-\!\!\sum\limits_{i=1}^{L_{p}}\delta_{mi}\!\Big)\gamma_{mk}+\!\sum\limits_{k=1}^{K}\!p^{u}_{k}(\beta_{mk}{-}\delta_{mi_{k}}\gamma_{mk}){+}1.$
 Consequently, the optimization algorithm (Algorithm 1) and the overall structure of the gradient $\nabla f(\bm{\delta}_{m})$ remain unchanged.
 The impact of the different SINR is confined solely to the calculation of the partial derivatives $\frac{\partial S_{mt}}{\partial\delta_{mi}}$ and $\frac{\partial I_{mt}}{\partial\delta_{mi}}$ within the gradient.
-For the G-PWPFZF SINR in (32), these derivatives are:
+For the G-PWPFZF scheme, these derivatives are:
 $\frac{\partial S_{mt}}{\partial\delta_{mi}}=-p^{u}_{t}\gamma_{mt},$ (32)
@@
 These expressions replace their G-PFZF counterparts in the gradient calculation step of Algorithm 1. All other steps of the algorithm proceed identically.
-IV Numerical Simulations
+IV Performance Evaluation
 TABLE I: Baseline Simulation Parameters
@@
 Shadow fading standard deviation 8 dB
-This section demonstrates the influence of the proposed improvements on the SE performance across multiple distributed combining strategies and local weights.
+This section demonstrates the gains of the our proposed decentralized decoding and combining schemes in terms of the sum SE, per user SE and computation cost.
 We consider a network with $M$ APs and $T$ UEs, uniformly distributed in a square area of size $1\times 1$ km2.
-The parameters are mentioned in Table I.
-Large-scale fading coefficients (LSFCs) are generated according to the model in [3], incorporating shadow fading.
+The relevant system parameters are mentioned in Table I.
+The LSFCs are generated according to the model in [3], incorporating shadow fading.
 Pilots are assigned using the scalable pilot assignment strategy from [3], ensuring efficient reuse under limited pilot resources.
 All APs are assumed to operate at full transmit power during both pilot training and data transmission.
-For the baseline PFZF and PWPFZF schemes, UE grouping is performed using a $90\%$ of total LSFCs received at the AP, which we identified using numerical simulation as providing the best performance.
-To ensure statistical reliability, all reported results are averaged over 500 independent simulation realizations.
+For the baseline PFZF and PWPFZF schemes, UE grouping is performed using a $90\%$ of total LSFCs received at the AP, are identified using numerical simulation, to provide the best performance.
+To ensure statistical reliability, all reported results are averaged over 500 independent network realizations.
 IV-A Computation cost and Fronthaul cost
-In this subsection, we analyze the computational cost and fronthaul signaling cost.
+In this subsection, we analyze the computational and fronthaul signaling costs of our proposed decentralized decoding and combining schemes.
 The costs are computed following the approach outlined in [16, 19, 15, 10], where only multiplication and division operations are counted per coherence block, as these dominate the computational load.
 TABLE II: Decoding Computation and Fronthaul Overhead per UE per Coherence Block
-Decoding Type Fronthaul Cost (Complex Scalars) Combining Weights Computation Cost
-o-LSFD [16, 15] $L_{u}M+\frac{3MT+M}{2}$ $\left(\frac{M^{2}+M}{2}\right)T+\frac{M^{3}-M}{3}+M^{2}$
-Proposed Decentralized Decoding $L_{u}M$ $\left(\frac{M^{2}+M}{2}\right)T+M+M^{2}$
-Table II compares the fronthaul and computational costs for optimal o-LSFD and the proposed decentralized decoding architecture.
-In conventional o-LSFD, each AP must transmit both channel estimates and local data estimates to the CPU, resulting in substantial fronthaul overhead of $L_{u}M+\frac{3MT+M}{2}$ complex scalars per UE per coherence block.
-Furthermore, o-LSFD requires computationally expensive matrix operations such as matrix inversion at the CPU that scale as $\mathcal{O}(TM^{2}+M^{3})$ per UE, making it impractical for large-scale deployments.
+Decoding Type Fronthaul Cost Computation Cost
+LSFD [16, 15] $L_{u}M+\frac{3MT+M}{2}$ $\left(\frac{M^{2}+M}{2}\right)T+\frac{M^{3}-M}{3}+M^{2}$
+Proposed Decoding $L_{u}M$ $\left(\frac{M^{2}+M}{2}\right)T+M+M^{2}$
+Table II compares the fronthaul and computational costs for optimal LSFD and the proposed decentralized decoding architecture.
+In conventional LSFD, each AP must transmit both channel estimates and local data estimates to the CPU, resulting in substantial fronthaul overhead of $L_{u}M+\frac{3MT+M}{2}$ complex scalars per UE per coherence block.
+Furthermore, LSFD requires computationally expensive matrix operations such as matrix inversion at the CPU that scale as $\mathcal{O}(TM^{2}+M^{3})$ per UE, making it impractical for large-scale network deployments.
 The proposed decentralized decoding architecture eliminates these bottlenecks by avoiding centralized LSFD entirely.
 Each AP independently computes and applies interference-suppressing local weights, reducing the fronthaul load to only $L_{u}M$ complex scalars.
-The computational cost is similarly streamlined to $\mathcal{O}(TM^{2})$ by eliminating the cubic $M^{3}$ term associated with large matrix inversions.
+The computational cost is similarly streamlined to $\mathcal{O}(TM^{2})$ by eliminating the cubic $M^{3}$ term associated with large matrix inversions required in the LSFD architecture due to centralized AP coordination.
 This combination of reduced fronthaul overhead and scalable computation enables practical deployment in dense network scenarios.
-Figure 2: Computational cost of weights calculation versus number of UEs, showing the significant computation cost reduction of the proposed decentralized decoding approach compared to o-LSFD.
-Figure 3: Fronthaul cost versus number of UEs, demonstrating the reduction achieved by the proposed decentralized decoding architecture by eliminating channel estimate transmission.
+Figure 2: Comparison of computational and fronthaul costs for different decoding and combining schemes.
 As demonstrated in Fig.
-2, the proposed decentralized decoding weights computation provides substantial computation cost reduction compared to o-LSFD, particularly as the number of UEs increases.
+2a, the proposed decentralized decoding weights computation provides substantial computation cost reduction compared to LSFD, particularly as the number of UEs increases.
 Similarly, Fig.
-3 shows that eliminating channel estimate transmission reduces fronthaul overhead.
+2b shows that eliminating channel estimate transmission reduces fronthaul overhead.
 TABLE III: Computational cost per AP per Coherence Block
@@
 G-PWPFZF $\frac{3L_{S_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}2(L_{p}{-}L_{\mathcal{S}_{l}})L_{\mathcal{S}_{l}}A{+}AL_{p}$
-Figure 4: Computational cost of combining vectors versus number of UEs, illustrating the scalability advantage of generalized schemes (G-PFZF, G-PWPFZF) over traditional approaches.
 The computational cost per AP per coherence block for calculating the combining vectors is detailed in Table III.
 The primary distinction between the traditional and proposed generalized schemes lies in their scaling behavior relative to the number of users.
-The traditional PFZF and PWPFZF schemes exhibit a computational cost that scales linearly with the total number of user equipments, $T$, as in by the $AT$ term for PFZF and PWPFZF.
+The traditional PFZF and PWPFZF schemes exhibit computational costs that scale linearly with the total number of user equipments, $T$, as in the $AT$ term for PFZF and PWPFZF.
 This dependence creates a significant computational bottleneck in densed networks with massive connectivity.
@@
 Fig.
-4 further validates the scalability benefits of the generalized schemes, where computational cost saturates once the number of UEs exceeds the available pilot sequences, unlike traditional schemes that continue to scale linearly with user density.
+2c further validates the scalability benefits of the generalized schemes, where computational cost saturates once the number of UEs exceeds the available pilot sequences, unlike traditional schemes that continue to scale linearly with user density.
 IV-B Performance Evaluation
-Figure 5: Distribution of Strong Pilot Decisions: How Each AP Adapts?
+Figure 3: Distribution of strong pilot decisions.
 Fig.
@@
 Most notably, the significant non-zero count at $L_{\mathcal{S}_{m}}{=}0$ demonstrates our algorithm’s ability to adaptively default to simple MR combining when it is the optimal strategy for a given AP’s local conditions.
-This demonstrates a seamless, adaptive switching between local zero-forcing and local MR processing that is impossible for the threshold-bound baseline.
+This demonstrates a seamless, adaptive switching between local zero-forcing and local MR processing that is impossible for the threshold-bound baselines.
 Figure 4: The uplink sum SE comparison for various combining schemes.
 Fig.
-6 depicts the sum spectral efficiency (SE) versus the number of UEs ($T$) for various combining schemes.
-The proposed distributed optimization framework delivers significant and scalable performance gains, with G-PFZF outperforming its baseline PFZF by 6.5% to 9.5% and G-PWPFZF surpassing PWPFZF by 5.5% to 10%, with the margin widening as the number of UEs increases.
-This superior performance highlights the critical limitation of the baseline PFZF and PWPFZF schemes, which rely on a hard, network-wide threshold using LSFCs to rigidly classify users as strong or weak at every AP.
-In contrast, the proposed framework introduces a distributed optimization algorithm, executed independently by each AP, to determine a better pilot-based grouping strategy.
-This allows an AP to adaptively select none, a subset, or all of the pilots to designate as strong, based on its local channel conditions and specific geometric configuration.
+4 depicts the sum SE versus the number of UEs ($T$) for various combining schemes.
+The proposed decentralized optimization framework delivers significant and scalable performance gains, with G-PFZF outperforming its baseline PFZF by 6.5% to 9.5% and G-PWPFZF surpassing PWPFZF by 5.5% to 10%, with the gaps widening as the number of UEs increases.
+Better performance of the proposed schemes highlights the critical limitation of the baseline PFZF and PWPFZF schemes, which rely on a hard, network-wide threshold using LSFCs to rigidly classify users as strong or weak at every AP.
+In contrast, the proposed framework introduces a decentralized optimization algorithm, executed independently by each AP, to determine a better pilot-based grouping strategy.
+This allows an AP to adaptively select none, a subset, or all of the pilots as strong, based on its local channel conditions and specific geometric configuration.
 This per-AP flexibility enables a more efficient and dynamic trade-off between utilizing degrees of freedom for interference suppression and desired signal enhancement.
-Consequently, the proposed G-PWPFZF scheme achieves superior performance over the G-PFZF scheme by suppressing interference from strong UEs for the weak UEs without a significant sacrifice in degrees of freedom.
+Consequently, the proposed G-PWPFZF scheme achieves better performance over the G-PFZF scheme by suppressing interference of the weak UEs from the strong UEs at the APs.
 Figure 5: The uplink sum SE comparison for various combining schemes.
 Fig.
-7 illustrates the sum SE versus the number of orthogonal pilots $L_{p}$, demonstrating the superior performance of the proposed G-PFZF and G-PWPFZF schemes over their respective baseline counterparts.
+5 illustrates the sum SE versus the number of orthogonal pilots $L_{p}$, demonstrating the improved performance of the proposed G-PFZF and G-PWPFZF schemes over their respective baseline counterparts.
 The proposed schemes achieve significant performance improvements of 6-14% for G-PFZF and 6-13% for G-PWPFZF compared to their respective baselines across different pilot lengths, with the performance gap widening as the number of pilots increases.
-This expanding margin highlights the advantage of the proposed adaptive user grouping strategy over the rigid network-wide threshold approach used in baseline schemes.
-The baseline schemes, constrained by a rigid network-wide threshold, must rigidly limit the number of strong pilots to avoid violating the fundamental condition for zero-forcing ($L_{\mathcal{S}_{m}}>A$), which leads to inefficient interference management.
-In contrast, our proposed distributed framework empowers each AP to independently and optimally select its set of strong pilots, inherently ensuring the selection is both locally optimal and always feasible ($L_{\mathcal{S}_{m}}<A$).
-This inherent adaptability allows our methods to leverage larger pilot sets more effectively, maximizing array gain and suppressing interference, which explains the observed widening of the performance gap.
-The G-PWPFZF scheme provides a further advantage by proactively protecting weak UEs from the strong UEs’ interference.
+This expanding gap highlights the advantage of the proposed adaptive user grouping strategy over the rigid network-wide threshold approach used in baseline schemes.
+The baseline schemes, must rigidly limit the number of strong pilots to avoid violating the fundamental condition for zero-forcing ($L_{\mathcal{S}_{m}}{>}A$), which leads to inefficient interference management.
+In contrast, our proposed decentralized framework empowers each AP to independently and optimally select its set of strong pilots, inherently ensuring that the selection is both locally optimal and always feasible ($L_{\mathcal{S}_{m}}{<}A$).
+This inherent adaptability allows our methods to leverage larger pilot sets more effectively, maximizing array gain and suppressing interference, which explains the observed widening of the performance gap with the number of pilots.
 Figure 6: The uplink sum SE comparison for various combining schemes.
 Fig.
-8 illustrates the sum spectral efficiency (SE) versus the number of antennas per AP ($A$) for various combining schemes, revealing a characteristic performance pattern that highlights the adaptability of our proposed framework.
+6 illustrates the sum SE versus the number of antennas per AP ($A$) for various combining schemes, revealing a characteristic performance pattern that highlights the adaptability of our proposed framework.
 The G-PFZF and G-PWPFZF schemes demonstrate consistent superiority across all antenna configurations, with the performance gap exhibiting a distinctive trajectory that reflects adaptive resource utilization.
@@
 As antenna numbers increase further (10-16 antennas), providing abundant spatial resources that benefit all schemes, the performance gap narrows to 4.5-6.5% for G-PFZF and 3.5-5.5% for G-PWPFZF, though our approach maintains consistent superiority.
-These results confirm that our adaptive optimization framework achieves maximum relative advantage in precisely those scenarios where adaptive resource allocation is most valuable, particularly in resource-constrained environments, while still delivering meaningful gains in antenna-rich deployments.
 Figure 7: The uplink sum SE comparison for various combining schemes.
 Fig.
-9 illustrates the sum spectral efficiency (SE) versus the number of UEs ($T$), comparing local and LSFD-based combining schemes.
+7 illustrates the sum SE versus the number of UEs ($T$), comparing local and LSFD-based combining schemes.
 For our G-PFZF and G-PWPFZF schemes, the performance gap between local weights and optimal LSFD remains minimal at just 3% across all UE densities.
-This marginal performance difference is dramatically outweighed by the substantial system benefits: our local weight approach eliminates the need for complex network-wide coordination, reducing computational cost at the CPU by avoiding large-scale matrix inversions and cutting fronthaul overhead by eliminating the need to share channel state information across the network.
-Despite this streamlined architecture, our locally-weighted G-PFZF and G-PWPFZF schemes still outperform the baseline PFZF and PWPFZF schemes with optimal LSFD by margins of 4.5-9% and 3.5-7.5% respectively.
+This marginal performance difference is dramatically outweighed by the substantial system benefits: our local weight approach eliminates the need for complex network-wide coordination, reducing computational cost at the CPU by avoiding large-scale matrix inversions and reducing fronthaul overhead by eliminating the need to share channel state information across the network.
+Despite this streamlined architecture, our locally-weighted G-PFZF and G-PWPFZF schemes still outperform the baseline PFZF and PWPFZF schemes with LSFD by margins of 4.5-9% and 3.5-7.5% respectively.
 This demonstrates that the gains from our adaptive per-AP grouping strategy are so substantial that they outperform the benefits of optimal coordination applied to suboptimal threshold-based grouping.
-These results confirm that our fully distributed approach achieves an optimal balance between performance and practicality.
+These results confirm that our fully decentralized approach achieves a better balance between performance and practicality.
 Figure 8: The uplink sum SE comparison for various combining schemes.
 Fig.
-10 illustrates the sum spectral efficiency (SE) versus the number of pilots ($L_{p}$), comparing local and LSFD-based combining schemes.
+8 illustrates the sum SE versus the number of pilots ($L_{p}$), comparing local and LSFD-based combining schemes.
 The performance gap between local and LSFD-based weights is less than 5.5% for sparse pilot configurations, where high pilot contamination makes centralized LSFD better equipped to handle interference.
@@
 Notably, even in the worst-case scenario with limited pilots, the performance penalty remains modest, validating that the substantial fronthaul and complexity reductions of our decentralized decoding approach come with acceptable performance trade-offs across all pilot sequences.
-Furthermore, the proposed G-PFZF and G-PWPFZF schemes with local weights consistently outperform traditional PFZF and PWPFZF with LSFD across moderate to high pilot lengths, confirming the advantage of adaptive per-AP combining strategies.
+Furthermore, the proposed G-PFZF and G-PWPFZF schemes with local weights consistently outperform PFZF and PWPFZF with LSFD across moderate to high pilot lengths, confirming the advantage of adaptive per-AP combining strategies.
 Figure 9: The uplink SE per user comparison for various combining schemes.
 Fig.
-11 presents the 90%-likely per-user spectral efficiency across different combining schemes, demonstrating remarkable improvements in user fairness through our distributed optimization framework.
-The proposed G-PFZF scheme achieves a substantial 45% improvement in 90%-likely SE compared to conventional PFZF, while G-PWPFZF shows a 34% gain over its baseline PWPFZF.
-These dramatic improvements confirm that our adaptive per-AP grouping strategy not only enhances overall system capacity but fundamentally transforms user experience across the network.
-The gains in 90%-likely SE particularly highlight our framework’s effectiveness in improving the performance of weak users who traditionally suffer from poor service quality.
-These results demonstrate that our distributed optimization framework successfully addresses both system-level efficiency and user-level fairness requirements in D-mMIMO deployments, achieving gains in quality-of-service uniformity.
-Figure 12: The uplink SE per user comparison for various combining schemes.
+9 presents the 90%-likely per-user SE across different combining schemes, demonstrating significant improvements in user fairness through our decentralized optimization framework.
+The proposed G-PFZF scheme achieves 45% improvement in 90%-likely SE compared to conventional PFZF, while G-PWPFZF shows a 34% gain over its baseline PWPFZF.
+These gains confirm that our adaptive per-AP grouping strategy not only enhances the overall system capacity but fundamentally transforms user experience across the network.
+The gains in 90%-likely SE particularly highlight our framework’s effectiveness in improving the performance of weak users.
+Figure 10: The uplink SE per user comparison for various decoding schemes.
 Fig.
-12 presents the 90%-likely per-user spectral efficiency comparing local versus LSFD processing, revealing crucial insights about the performance-complexity trade-off in our distributed framework.
-While our G-PFZF and G-PWPFZF schemes with local weights experience a modest performance reduction of approximately 10-10.7% compared to their LSFD-based counterparts, the 10% performance difference between local and LSFD processing represents a reasonable trade-off for achieving full distributability, substantially reduced fronthaul overhead, and lower computational cost.
-Remarkably, our locally-weighted G-PFZF scheme outperforms the baseline PFZF with optimal LSFD by 32%, and G-PWPFZF with local weights surpasses the LSFD-enhanced PWPFZF baseline by 22.5%.
+10 presents the 90%-likely per-user SE comparing local versus LSFD processing, revealing crucial insights about the performance-complexity trade-off in our decentralized framework.
+While our G-PFZF and G-PWPFZF schemes with local weights experience a modest performance reduction of approximately 10-10.7% compared to their LSFD-based counterparts, the 10% performance difference between local and LSFD processing represents a reasonable trade-off for achieving fully decentralized architecture, substantially reduced fronthaul overhead, and lower computational cost.
+Noticeably, our locally-weighted G-PFZF scheme outperforms the baseline PFZF with optimal LSFD by 32%, and G-PWPFZF with local weights surpasses the LSFD-enhanced PWPFZF baseline by 22.5%.
 These results demonstrate that the performance gains from our adaptive per-AP grouping strategy are so significant that they outweigh the benefits of optimal LSFD coordination applied to suboptimal threshold-based grouping.
-This analysis confirms that our distributed optimization framework provides a balance between performance and practical implementation constraints, delivering better SE per user while maintaining architectural advantages crucial for deployments.
-V Conclusion
-This paper has addressed the critical bottlenecks of sub-optimal performance, high fronthaul load, and computational cost in distributed massive MIMO networks by introducing a novel, decentralized decoding uplink architecture that fundamentally departs from conventional LSFD-based designs.
+V Conclusion and Future Work
+This paper has addressed the critical bottlenecks of sub-optimal performance, high fronthaul load, and computational cost in D-mMIMO networks by introducing a novel, decentralized decoding uplink architecture that fundamentally departs from conventional LSFD-based designs.
 Through the proposed local ZF framework, we have demonstrated that enabling each AP to independently determine its combining strategy via local optimization, classifying pilots as strong or weak without fixed thresholds and dynamically switching among PFZF/PWPFZF, FZF, or MR, yields significant performance improvements.
@@
 Moreover, the introduction of pilot-dependent combining vectors and interference-suppressing local weights applied distributively at each AP eliminates the need for centralized LSFD, thereby drastically reducing fronthaul overhead and computational cost.
-Remarkably, this distributed approach incurs only a minimal performance penalty compared to idealized centralized coordination, while even outperforming conventional threshold-based schemes using optimal LSFD.
-These findings collectively establish that adaptive processing at the network, not complex centralized decoding, is the key to scalability and efficiency in future D-mMIMO systems.
-The proposed framework offers a practical and high-performance pathway toward realizing scalable distributed MIMO networks without compromising on quality-of-service or imposing prohibitive infrastructure costs.
--A Proof of closed-form SINR expression for G-PFZF combining:
-To derive the closed-form expression, we need few expectation properties,
-$\begin{split}\mathbb{E}\Big\{\Big|\sum_{m}a_{m}\mathbf{v}_{m}^{H}\mathbf{g}_{m}\Big|^{2}\Big\}=\sum_{m}|a_{m}|^{2}\mathbb{E}\Big\{\Big|\mathbf{v}_{m}^{H}\mathbf{g}_{m}\Big|^{2}\Big\}+\Big|\sum_{m}a_{m}\mathbb{E}\Big\{\mathbf{v}_{m}^{H}\mathbf{g}_{m}\Big\}\Big|^{2}-\sum_{m}\Big|a_{m}\mathbb{E}\Big\{\mathbf{v}_{m}^{H}\mathbf{g}_{m}\Big\}\Big|^{2},\end{split}$ (35)
-Also, if $\mathbf{v}_{m}$ and $\mathbf{g}_{m}$ are indpendent random vectors of N length with zero mean and elements following i.i.d., then
-$\mathbb{E}\Big\{\Big|\mathbf{v}_{m}^{H}\mathbf{g}_{m}\Big|^{2}\Big\}=\frac{\mathbb{E}\Big\{\Big|\mathbf{v}_{m}\Big|^{2}\Big\}\mathbb{E}\Big\{\Big|\mathbf{g}_{m}\Big|^{2}\Big\}}{N}.$ (36)
-We derive the closed-form expression of desired signal term in (6) using the combining vectors in (7) and (8).
-$|\text{DS}_{t}|^{2}=p^{u}_{t}\Big|\mathbb{E}\Big\{\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mt}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mt}\Big\}\Big|^{2},$ (37)
-where $\mathcal{Z}_{i_{t}}$ is the subset of APs for which pilot $i_{t}$ is strong and $\mathcal{Y}_{i_{t}}$ is the subset of APs for which pilot $i_{t}$ is weak.
-Using (10) and (12), the (37) can be written as:
-$|\text{DS}_{t}|^{2}=p^{u}_{t}\Bigg|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}\sqrt{A\gamma_{mk}}\Bigg|^{2}=p^{u}_{t}\Bigg|\sum_{m=1}^{M}a_{mt}\sqrt{(A-\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2},$ (38)
-where $\delta_{mi_{t}}$ is defined in (15). The first term of interference in (6) can be expanded as:
-$\mathbb{E}\Big\{|\text{BU}_{t}|^{2}\Big\}=p^{u}_{t}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mt}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mt}\Big|^{2}\Bigg\}-|\text{DS}_{t}|^{2},$ (39)
-We focus on first term of $\mathbb{E}\Big\{|\text{BU}_{t}|^{2}\Big\}$, which can be written as:
-$\begin{split}\!\!\!\!\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}\Big|^{2}\!\Bigg\}+\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mt}\Big|^{2}\Bigg\}+2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nt}\!\Bigg\},\end{split}$ (40)
-The first term of (40) can be written as:
-$\begin{split}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!(\hat{\mathbf{g}}_{mt}+\tilde{\mathbf{g}}_{mt})\Big|^{2}\!\Bigg\}&=\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mt}\Big|^{2}\!\Bigg\}+\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mt}\Big|^{2}\!\Bigg\}\\ &\overset{(a)}{=}\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\mathbb{E}\Big\{\!\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mt}\Big|^{2}\!\Big\}+\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mt}\!\Big\}\Big|^{2}\\ &-\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mt}\!\Big\}\Big|^{2}+\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\mathbb{E}\Big\{\!\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mt}\Big|^{2}\!\Big\}\\ &+\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mt}\!\Big\}\Big|^{2}-\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mt}\!\Big\}\Big|^{2}\\ &\overset{(b)}{=}\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(A-L_{\mathcal{S}_{m}})\gamma_{mt}+\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}\\ &-\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(\beta_{mt}-\gamma_{mt})\\ &\overset{(c)}{=}\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(\beta_{mt}-\gamma_{mt})\end{split}$ (41)
-we get (a) using same way as in (35) and (b) using the (10), (11) and (36).
-Similarly, the second term of (40) is first expanded same as in (a) in (41) using (35), then using the (12), (13) and (36), can be written as
-$\begin{split}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!(\hat{\mathbf{g}}_{mt}+\tilde{\mathbf{g}}_{mt})\Big|^{2}\!\Bigg\}&{=}\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(A+1)\gamma_{mt}+\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mt}}\Big|^{2}\\ &-\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\sqrt{A\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(\beta_{mt}-\gamma_{mt})\\ &=\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\beta_{mt}.\end{split}$ (42)
-The expectation in closed-form in third term of (40) can be written as:
-$\begin{split}2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nt}\!\Bigg\}=2\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Bigg)\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!a_{mt}\sqrt{A\gamma_{mt}}\Bigg).\end{split}$ (43)
-Using (41), (42) and (43) in (40) and using this in (39), we get
-$\mathbb{E}\Big\{|\text{BU}_{t}|^{2}\Big\}=|\text{DS}_{t}|^{2}+p^{u}_{t}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mt}-\delta_{mi_{t}}\gamma_{mt})-|\text{DS}_{t}|^{2}=p^{u}_{t}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mt}-\delta_{mi_{t}}\gamma_{mt}).$ (44)
-The second term i.e., the pilot contamination term for the UE $t$ with the pilot sharing UE $k$ in (6), can be evaluated in closed-form as
-$\begin{split}&\mathbb{E}\Big\{|\text{PC}_{tk}|^{2}\Big\}=p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mk}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mk}\Big|^{2}\Bigg\},\end{split}$ (45)
-The RHS of (45) can be written as
-$\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\Bigg\}+\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}+2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}$
-Following the same steps as in (41), (42) and (43)
-$\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}=\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}+\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(\beta_{mk}-\gamma_{mk}),$ (46)
-$\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}{=}\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mk}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\beta_{mk},$ (47)
-$2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}=2\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg)\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!a_{mt}\sqrt{A\gamma_{mk}}\Bigg),$ (48)
-Therefore
-$\begin{split}&\mathbb{E}\Big\{|\text{PC}_{tk}|^{2}\Big\}=p^{u}_{k}\Big|\sum_{m=1}^{M}a_{mt}\sqrt{(A-\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}+p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (49)
-The third interference term, for any non-pilot sharing UEs pair $t,k$, in (6) can be written as:
-$\begin{split}\mathbb{E}\Big\{|\text{UI}_{tk}|^{2}\Big\}&=p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mk}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\Bigg\}+p^{u}_{k}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}\\ &+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}(1-\delta_{mi_{k}})\gamma_{mk}+p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}(\beta_{mk}-\gamma_{mk})\\ &+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\gamma_{mk}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}(\beta_{mk}-\gamma_{mk})\\ &=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (50)
-The last term of interference in (6) can be written as:
-$\begin{split}\mathbb{E}\Big\{|\text{GN}_{t}|^{2}\Big\}&=p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}+p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (51)
-Using the (38), (44), (49), (50) and (51) in (6) gives the SINR closed-form expression in (14).
--B Proof of Lipschitz Continuity of $\nabla f(\bm{\delta}_{m})$:
+Remarkably, this decentralized approach incurs only a small performance penalty compared to idealized centralized coordination, while even outperforming conventional threshold-based schemes using optimal LSFD.
+These findings collectively establish that proposed decentralized decoding in the network, not centralized decoding, is the key to scalability and efficiency in future D-mMIMO systems without compromising on quality-of-service.
+In future, we plan to investigate the robustness of the proposed decentralized architecture and generalized combining schemes under more practical and challenging conditions, such as hardware impairments.
+-A Derivation of closed-form SINR expression for G-PFZF combining
+In order to derive the closed-form expression, we need few expectations computation or expansions
+$\begin{split}\mathbb{E}\Big\{\Big|\!\!\sum_{m}\!\!\mathbf{v}_{mi}^{H}\mathbf{g}_{mi}\!\Big|^{2}\!\Big\}\!\!=&\!\sum_{m}\!\mathbb{E}\Big\{\!\Big|\mathbf{v}_{mi}^{H}\mathbf{g}_{mi}\!\Big|^{2}\!\Big\}+\!\Big|\!\sum_{m}\!\mathbb{E}\Big\{\mathbf{v}_{mi}^{H}\mathbf{g}_{m}\Big\}\Big|^{2}-\sum_{m}\Big|\mathbb{E}\Big\{\mathbf{v}_{mi}^{H}\mathbf{g}_{mi}\Big\}\Big|^{2},\end{split}$ (A1)
+$\mathbb{E}\Big\{\Big|\mathbf{v}_{mi_{t}}^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Big\}=$ $\beta_{mk}-\gamma_{mk}$ (A2)
+$\mathbb{E}\Big\{\Big|\mathbf{v}_{mi_{t}}^{H}{\mathbf{n}}_{m}\Big|^{2}\Big\}=$ $1.$ (A3)
+The desired signal term in (6) can be written as
+$|\text{DS}_{t}|^{2}=\Big|\mathbb{E}\Big\{\!\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{{g}}}_{mt}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{{g}}}_{mt}\Big\}\Big|^{2},$
+where $\mathcal{Z}_{i_{t}}$ and $\mathcal{Y}_{i_{t}}$ denote the subsets of APs where pilot $i_{t}$ is classified as strong and weak, respectively.
+Using (10) and (12), the $|\text{DS}_{t}|^{2}$ can be written as:
+$\begin{split}|\text{DS}_{t}|^{2}&=\!\!\Big|\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mk}}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mk}}\Big|^{2}\\ &=\Big|\sum_{m=1}^{M}\!a_{mt}\sqrt{(A-\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2},\end{split}$ (A4)
+The interference term $\mathbb{E}\big\{|\text{UI}_{tk}|^{2}\big\}$ in (6) can be expanded as:
+$\begin{split}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mk}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}&=\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\!\Bigg\}\!+\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\!\Bigg\}\\ &+2\mathbb{E}\Bigg\{\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}.\end{split}$ (A5)
+Using (A1), the first term ($A_{1}$) of (A5) can be rewritten as:
+$\begin{split}A_{1}&=\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!(\hat{\mathbf{g}}_{mk}+\tilde{\mathbf{g}}_{mk})\Big|^{2}\!\Bigg\}=\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}\!+\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}\\ &=\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\Big|\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\!\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}\\ &+\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|(a_{mt}\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}+\!\Big|\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2},\end{split}$
+Using (10), (11) and (A2), we can rewrite $A_{1}$ as
+$\begin{split}A_{1}&=\Big|\!\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (A6)
+Using (A1), the second term ($A_{2}$) of (A5) is expanded as:
+$\begin{split}A_{2}&=\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\Big|\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\!\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}\\ &+\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|(a_{mt}\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}+\!\Big|\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2},\end{split}$
+Using the (12), (13) and (A2), we can rewrite $A_{2}$ as
+$\begin{split}A_{2}&=\Big|\!\!\!\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mk}}\Big|^{2}\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}\!\!+\!\!\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!|a_{mt}|^{2}\beta_{mk}.\end{split}$ (A7)
+Finally, the third term ($A_{3}$) of (A5) can be written as:
+$\begin{split}A_{3}=2\Big(\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mt}}\!\Big)\Big(\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!a_{mt}\sqrt{A\gamma_{mt}}\Big).\end{split}$ (A8)
+Using (A6), (A7) and (A8)
+$\begin{split}\mathbb{E}\Big\{\!|\text{UI}_{tk}|^{2}\!\Big\}&=\!\!\Big|\!\!\sum\limits_{m=1}^{M}\!\!\!a_{mt}\sqrt{(A{-}\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\!\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\!\right|^{2}+\sum\limits_{m=1}^{M}\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (A9)
+Using (A3), the term $\mathbb{E}\big\{|\text{GN}_{t}|^{2}\big\}$ of interference in (6) can be written as:
+$\begin{split}&\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!|a_{mt}|^{2}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Big\}+\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!\!a_{mt}|^{2}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Big\}=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (A10)
+Using the (A4), (A9) and (A10) in (6) gives the SINR closed-form expression in (14).
+-B Proof of Lipschitz Continuity of $\nabla f(\bm{\delta}_{m})$
 To establish the Lipschitz continuity of $\nabla f(\bm{\delta}_{m})$, we analyze its component functions and their boundedness over the compact domain $\bm{\delta}_{m}\in[0,1]^{L_{p}}$.
-Consider the gradient component:
-$g(\delta_{mi})=\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}=$ $\sum_{t=1}^{T}\frac{I_{mt}}{I_{mt}+S_{mt}}\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}$
-$-\chi\left[2\lambda_{1}\max\left(0,\delta_{mi}-\delta^{2}_{mi}\right)\left(1-2\delta_{mi}\right)+2\lambda_{2}\max\left(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\right)\right]$
-Boundedness Analysis:
-Signal and Interference Terms: For $\delta_{mi}\in[0,1]$:
-  • $S_{mt}>0$ and $I_{mt}>0$ (positive definite)
-  • $S_{mt},I_{mt}\leq C_{1}$ (bounded above by network parameters)
-  • $\frac{\partial S_{mt}}{\partial\delta_{mi}},\frac{\partial I_{mt}}{\partial\delta_{mi}}\leq C_{2}$ (bounded derivatives from (25) and (26))
-Penalty Terms:
+The gradient component in (24) is $g(\delta_{mi})=\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}$ bounded as
+  • $0<S_{mt},I_{mt}\leq C_{1}$ (always bounded)
+  • $\frac{\partial S_{mt}}{\partial\delta_{mi}},\frac{\partial I_{mt}}{\partial\delta_{mi}}\leq C_{2}$ (bounded from (25) and (26))
   • $\max(0,\delta_{mi}-\delta^{2}_{mi})\in[0,\frac{1}{4}]$ (bounded on $[0,1]$)
-  • $\max\left(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\right)\leq L_{p}$ (finite sum constraint)
-Composite Function Bounds:
-  • $\frac{I_{mt}}{I_{mt}+S_{mt}}\in(0,1)$
-  • $\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}\leq C_{3}$
-Lipschitz Constant Derivation:
-The partial derivatives of $g(\delta_{mi})$ exist and are bounded on the compact set $0,1]^{L_{p}}$:
-$\begin{split}\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}&=\sum_{t=1}^{T}\Bigg[\frac{S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}-I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}}{\big(I_{mt}+S_{mt}\big)^{2}}\cdot\frac{I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}\\ &\quad+\frac{I_{mt}}{I_{mt}+S_{mt}}\left(\frac{I_{mt}\,\frac{\partial^{2}S_{mt}}{\partial\delta_{mi}^{2}}-S_{mt}\,\frac{\partial^{2}I_{mt}}{\partial\delta_{mi}^{2}}}{I_{mt}^{2}}-\frac{2\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big(I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big)}{I_{mt}^{3}}\right)\Bigg]\\ &\quad-\chi\Bigg[2\lambda_{1}\,\mathbf{1}_{\{\delta_{mi}-\delta_{mi}^{2}>0\}}\Big((1-2\delta_{mi})^{2}-2(\delta_{mi}-\delta_{mi}^{2})\Big)+2\lambda_{2}\,\mathbf{1}_{\{2-A>0\}}\Bigg].\end{split}$
-For all $\delta_{mi}\in[0,1]$, $\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}$ is bounded:
-$\left|\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}\right|\leq C_{4}$
-Since all component functions are bounded and continuously differentiable on the compact domain, there exists a global Lipschitz constant $L$ such that:
-$\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\|\leq L\|\mathbf{x}-\mathbf{y}\|,\quad\forall\mathbf{x},\mathbf{y}\in[0,1]$
-This establishes the Lipschitz continuity of $\nabla f(\bm{\delta}_{m})$, ensuring convergence of the proximal gradient algorithm with appropriate step size selection.
--C Proof of closed-form SINR expression for G-PWPFZF combining:
+  • $\max\left(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\right)\leq L_{p}$
+  • $\frac{I_{mt}}{I_{mt}+S_{mt}}\in(0,1),\qquad C_{3}\leq\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}\leq C_{4}$,
+where $C_{1}$, $C_{3}$, $C_{3}$ and $C_{4}$ are finite constants.
+The partial derivatives of $g(\delta_{mi})$ is given by
+$\begin{split}\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}&=\!\!\sum_{t=1}^{T}\!\!\Bigg[\frac{S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}-I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}}{\big(I_{mt}+S_{mt}\big)^{2}}\frac{I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}+\frac{I_{mt}}{I_{mt}+S_{mt}}\Bigg(\frac{I_{mt}\,\frac{\partial^{2}S_{mt}}{\partial\delta_{mi}^{2}}-S_{mt}\,\frac{\partial^{2}I_{mt}}{\partial\delta_{mi}^{2}}}{I_{mt}^{2}}\\ &-\frac{2\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big(I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big)}{I_{mt}^{3}}\Bigg)\Bigg]-\chi\Bigg[2\lambda_{1}\,\mathbf{1}_{\{\delta_{mi}-\delta_{mi}^{2}>0\}}\Big((1-2\delta_{mi})^{2}\\ &-2(\delta_{mi}-\delta_{mi}^{2})\Big)+2\lambda_{2}\,\mathbf{1}_{\{2-A>0\}}\Bigg].\end{split}$
+For all $\delta_{mi}\in[0,1]$, since all component functions $\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}$ are bounded and continuously differentiable on the compact domain $0,1]^{L_{p}}$, the Hessian matrix $\nabla^{2}f(\bm{\delta}_{m})=\nabla g(\bm{\delta}_{m})$ exists and is continuous.
+Applying the mean value theorem for vector-valued functions, for any $\mathbf{x},\mathbf{y}\in[0,1]^{L_{p}}$, there exists $\mathbf{z}$ on the line segment between $\mathbf{x}$ and $\mathbf{y}$ such that:
+$\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\|\leq\|\nabla^{2}f(\mathbf{z})\|_{2}\|\mathbf{x}-\mathbf{y}\|\leq L\|\mathbf{x}-\mathbf{y}\|$
+This establishes the Lipschitz continuity of $\nabla f(\bm{\delta}_{m})$, ensuring convergence of the proximal gradient algorithm.
+-C Derivation of closed-form SINR expression for G-PWPFZF combining
 The derivation of G-PWPFZF SINR closed-form expression follows the same step as the derivation of G-PFZF SINR closed-form expression, the only difference is that instead of ${\mathbf{v}}^{\text{MR}}_{mi_{t}}$, we have ${\mathbf{v}}^{\text{PMR}}_{mi_{t}}$ combining vector.
-The desired signal expression in (6) using the combining vectors in (7) and (27).
-$|\text{DS}_{t}|^{2}=p^{u}_{t}\Big|\mathbb{E}\Big\{\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mt}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\mathbf{{g}}_{mt}\Big\}\Big|^{2},$ (52)
-Using (10) and (28), the (52) can be written as:
-$|\text{DS}_{t}|^{2}=p^{u}_{t}\Bigg|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2}=p^{u}_{t}\Bigg|\sum_{m=1}^{M}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2},$ (53)
-The first term $\mathbb{E}\Big\{|\text{BU}_{t}|^{2}\Big\}$ of interference in (6) can be expanded as:
-$\begin{split}\!\!\!\!\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}\Big|^{2}\!\Bigg\}+\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{mt}\Big|^{2}\Bigg\}+2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}(\mathbf{v}_{ni_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{nt}\!\Bigg\},\end{split}$ (54)
-The first term of (54) is same as (41), the second term of (54) can be written as
-$\begin{split}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\!(\hat{\mathbf{g}}_{mt}+\tilde{\mathbf{g}}_{mt})\Big|^{2}\!\Bigg\}&{=}\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(A-L_{\mathcal{S}_{m}}+1)\gamma_{mt}+\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}\\ &-\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}(\beta_{mt}-\gamma_{mt})\\ &=\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\beta_{mt}.\end{split}$ (55)
-The expectation in closed-form in third term of (54) can be written as:
-$\begin{split}2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mt}(\mathbf{v}_{ni_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{nt}\!\Bigg\}=2\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Bigg)\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}}\Bigg).\end{split}$ (56)
-Using (41), (55) and (56) in (54), we get
-$\mathbb{E}\Big\{|\text{BU}_{t}|^{2}\Big\}=p^{u}_{t}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mt}-\delta_{mi_{t}}\gamma_{mt}).$ (57)
-The second term i.e., the pilot contamination term $\mathbb{E}\Big\{|\text{PC}_{tk}|^{2}\Big\}$ for UE $t$ with pilot sharing UE $k$ in (6), can be evaluated as
-$\begin{split}\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\Bigg\}+\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}+2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}\end{split}$ (58)
-The first term of (58) is same as (46). The second and third term of (58) can be computed as:
-$\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\!{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}{=}\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!|a_{mt}|^{2}\beta_{mk},$ (59)
-$2\mathbb{E}\Bigg\{\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{PMR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}=2\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg)\Bigg(\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg),$ (60)
-Using (46), (59) and (60)
-$\begin{split}&\mathbb{E}\Big\{|\text{PC}_{tk}|^{2}\Big\}=p^{u}_{k}\Big|\sum_{m=1}^{M}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}+p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (61)
-The third interference term, for any non-pilot sharing UEs pair $t,k$, in (6) can be written as:
-$\begin{split}\mathbb{E}\Big\{|\text{UI}_{tk}|^{2}\Big\}&={=}p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}\\ &+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Bigg\}\\ &{=}p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}(1-\delta_{mi_{k}})\gamma_{mk}+p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}(\beta_{mk}-\gamma_{mk})\\ &+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}(1-\delta_{mi_{k}})\gamma_{mk}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}(\beta_{mk}-\gamma_{mk})\\ &=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (62)
-The last term of interference in (6) can be written as:
-$\begin{split}\mathbb{E}\Big\{|\text{GN}_{t}|^{2}\Big\}&=p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}+p^{u}_{k}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m\in\mathcal{Z}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}+p^{u}_{k}\sum_{m\in\mathcal{Y}_{i_{t}}}|a_{mt}|^{2}\mathbb{E}\Bigg\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Bigg\}\\ &=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (63)
-Using the (53), (57), (61), (62) and (63) in (6) gives the SINR closed-form expression in (30).
+The desired signal term in (6)
+$|\text{DS}_{t}|^{2}=\Big|\mathbb{E}\Big\{\!\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{{g}}}_{mt}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{{g}}}_{mt}\Big\}\Big|^{2}.$
+Using (10) and (28), then
+$|\text{DS}_{t}|^{2}=p^{u}_{t}\Bigg|\sum_{m=1}^{M}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2},$ (C1)
+Following the steps in Appendix -A, the first term in (6) can be written as
+$\begin{split}\mathbb{E}\Big\{\!|\text{UI}_{tk}|^{2}\!\Big\}&=\!\!\Big|\!\!\sum\limits_{m=1}^{M}\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\!\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+\sum\limits_{m=1}^{M}\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (C2)
+Similarly, the last term of interference in (6) can be written as:
+$\begin{split}\mathbb{E}\Big\{|\text{GN}_{t}|^{2}\Big\}=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (C3)
+Using the (C1), (C2) and (C3) in (6) gives the SINR closed-form expression in (30).

### v3

@@
 Abstract
-A major bottleneck in uplink distributed massive multiple-input multiple-output networks is the sub-optimal performance of local combining schemes, coupled with high fronthaul load and computational cost inherent in centralized large scale fading decoding (LSFD) architectures.
-This paper introduces a decentralized decoding architecture that fundamentally breaks from the LSFD, by allowing each access point (AP) to calculate interference-suppressing local weights independently and apply them to its data estimates before transmission.
-Furthermore, two generalized local zero-forcing (ZF) frameworks, generalized partial full-pilot ZF (G-PFZF) and generalized protected weak PFZF (G-PWPFZF), are introduced, where each AP adaptively and independently determines its combining strategy through a local sum spectral efficiency (SE) optimization that classifies user equipments (UEs) as strong or weak, eliminating the fixed thresholds used in the PFZF and PWPFZF schemes.
-To enhance scalability, pilot-dependent combining vectors instead of user-dependent ones are introduced and are shared among users with the same pilot.
-The closed-form SE expressions corresponding to the proposed schemes are derived.
-Numerical results show that the proposed schemes consistently outperform fixed-threshold counterparts, while the introduction of local weights yields lower overheads and computation costs with lower performance penalty compared to them.
+Efficient uplink processing in distributed massive multiple-input multiple-output (D-mMIMO) systems requires both effective local combining and scalable decoding to significantly mitigate inter-user interference.
+Recent zero-forcing (ZF)-based combining schemes, such as partial full-pilot ZF (PFZF) and protected weak PFZF (PWPFZF), rely on heuristic threshold-based user grouping that may lead to inefficient utilization of spatial degrees of freedom across access points (APs).
+To address this limitation, we propose adaptive pilot-aware local combining strategies, generalized PFZF (G-PFZF) and generalized PWPFZF (G-PWPFZF), that dynamically allocate spatial degrees of freedom based on local channel conditions and replace heuristic grouping with a decentralized pilot-level optimization framework.
+Thus providing substantial performance gains over conventional PFZF and PWPFZF.
+Further, centralized decoding has recently emerged as a promising technique for interference suppression in D-mMIMO systems.
+However, it incurs substantial fronthaul overhead and computational costs.
+We develop a decentralized large-scale fading decoding (d-LSFD) scheme in which each AP computes LSFD weights using only locally available channel statistics.
+We derive a lower bound on the signal-to-interference-plus-noise ratio that explicitly quantifies the performance gap between the proposed d-LSFD scheme and centralized LSFD (c-LSFD), and identifies conditions under which the proposed decentralized solution approaches the centralized optimum.
+Numerical results demonstrate that the proposed generalized combining and the d-LSFD scheme together achieve significantly higher sum spectral efficiency in comparison to any combination of existing local combining and decoding schemes, while also substantially reducing the computational cost and fronthaul overhead.
 I Introduction
-Distributed massive MIMO (D-mMIMO) networks are a promising framework for next-generation wireless systems, as they can enhance weak users’ signal and guarantee uniform service quality across wide coverage areas by enabling numerous distributed access points (APs) to coherently serve user equipments (UEs) [1, 2].
-To fully use the advantages of D-mMIMO, while ensuring scalability, distributed signal processing at the APs is essential [3, 4].
-A major challenge in this regard is the development of effective uplink combining schemes that can alleviate inter-user interference within practical limitations.
-Centralized zero-forcing (ZF) combining, as studied in [5, 6, 7], requires substantial fronthaul signaling and imposes prohibitively high computational costs.
-Similarly, centralized minimum mean square error (MMSE) combining [8] suffers from both high computation cost and fronthaul overhead, which makes it impractical for large-scale deployments.
-To alleviate this, partial MMSE combining is proposed in [9, 3, 10], where the interference from only a subset of UEs is considered when designing the combining vector for each UE.
-Along similar lines, partial centralized approaches have been explored in [11, 12], where joint ZF is applied across a subset of APs based on stronger channels.
-These partial centralized ZF schemes reduce the fronthaul load compared to fully centralized ZF, but the overhead still scales with the number of UEs, as for each UE, a subset of APs need to send the channel estimates to the CPU, which limits their suitability for ultra-dense network scenarios.
-These centralized and partial centralized combining schemes follow the fully centralized architecture as shown in Fig.
-1, which has very high computational cost and fronthaul overhead, as for all UEs, the channel needs to be estimated at the CPU with the cooperation of all or a subset of APs.
-Several distributed combining approaches have been proposed in the literature.
-Maximum ratio (MR) combining, introduced in [1], is computationally simple and fronthaul-efficient, but its performance is fundamentally limited by its inability to effectively mitigate inter-user interference.
-To improve upon the MR, a local partial MMSE scheme is proposed in [3, 10], which achieves superior interference suppression and performance.
-However, a key drawback of this approach is that closed-form expressions for the spectral efficiency (SE) cannot be derived.
-As a result, resource allocation and system optimization must rely on computationally expensive Monte Carlo simulations, and the lack of tractable expressions makes theoretical performance analysis and gaining insights into the system behavior difficult.
-In [1, 3], the authors follow a simple decoding architecture as in Fig.
-1, thus reducing the overhead apart from local data estimates, yet it suffers from sub-optimality with respect to centralized decoding.
-The local combining schemes such as full-pilot ZF (FZF), partial FZF (PFZF), and protected weak PFZF (PWPFZF) are proposed in [13, 14], offering tractable closed-form SE expressions.
-In these schemes, each AP classifies the UEs into strong and weak groups, generates one combining vector per orthogonal pilot: the ZF for strong UEs and the MR for weak UEs, and assigns scaled versions to individual UEs based on their estimated channel variances [13].
-The fundamental limitations of the PFZF and PWPFZF schemes stem from two factors.
-First, the rigid, threshold-based user grouping may misclassify UEs, leading to suboptimal interference suppression and inefficient use of spatial resources.
-Second, all APs are constrained to follow the same combining scheme, which prevents each AP from independently adapting its strategy to local channel conditions and interference patterns, further limiting overall system performance.
-The large-scale fading decoding (LSFD) as shown in Fig.
-1, is employed at the CPU to enhance uplink performance in D-mMIMO networks [8, 15, 13, 14].
-In the optimal LSFD approach [13], each AP must transmit both channel and data estimates for all served UEs to the CPU, effectively doubling the fronthaul load while imposing computational costs that scale with the number of users [16].
-To alleviate this burden, partial LSFD is proposed in [15], which reduces computational cost and scalability requirements by considering only partial interference in the combining process.
-However, partial LSFD still requires matrix inversions of the same dimension as optimal LSFD and transmission of the channel estimates to the CPU.
-As network size and UE density increase, these operations become increasingly demanding with respect to computation and overhead, significantly limiting the scalability and practicality of the LSFD-based architectures in ultra-dense D-mMIMO deployments.
-(a) FullyCentralized(b) LSFD(c) SimpleDecoding(d) Proposed DecentralizedDecoding APs CPU All Data Global Combining Global Estimates APs Local Combining CPU Local estimates $\hat{x}_{mt}$Channel estimates LSFD Weights Final Estimates APs Local Combining CPU $\hat{x}_{mt}$ $\hat{x}_{t}=\sum\limits_{m}\hat{x}_{mt}$ APs Local Combining Local Weights $a_{mt}$ CPU $a_{mt}\hat{x}_{mt}$ $\hat{x}_{t}=\sum\limits_{m}a_{mt}\hat{x}_{mt}$
-Figure 1: Uplink processing architectures: (a) Fully centralized, (b) LSFD with coordination, (c) Simple decoding, (d) Proposed decentralized decoding.
+Distributed massive MIMO (D-mMIMO) systems have emerged as a promising architecture for next-generation wireless systems, as they can provide uniform service quality across the network by enabling numerous distributed access points (APs) to coherently serve user equipments (UEs) [1, 2].
+However, achieving scalable signal processing in uplink D-mMIMO systems remains challenging because centralized processing architectures require significant fronthaul signaling and computational overhead [3, 4].
+Several centralized and partially centralized combining architectures have been proposed to mitigate inter-user interference in D-mMIMO systems by jointly processing signals or channel statistics across multiple APs.
+[5, 6, 7, 3, 8, 9, 10].
+While these approaches can achieve strong interference suppression, they require the central processing unit (CPU) to collect signals or channel statistics from multiple APs, resulting in substantial fronthaul overhead and computational costs, as shown in Fig.
+1(a).
+To reduce such costs, a two-layer decoding architecture has been proposed in [1, 3, 8], where each AP performs local combining while the CPU performs a second-stage decoding.
+Consequently, the overall performance of D-mMIMO systems critically depends on: the design of efficient local combining schemes at the APs and the design of scalable second-stage decoding techniques at the CPU.
+Local maximum ratio (MR) combining [1], offers low computational complexity, but suffers from limited interference suppression capability.
+The local partial minimum mean square error (LP-MMSE) combining scheme [3, 8], achieves better interference suppression and performance.
+However, it does not provide tractable closed-form spectral efficiency (SE) expressions, making system-level analysis and optimization challenging.
+To address these limitations of the MR and LP-MMSE combining, recent local combining schemes such as full-pilot ZF (FZF), partial FZF (PFZF), and protected weak PFZF (PWPFZF)  [11, 14], exploit the structure of pilot reuse to construct combining vectors that partially suppress interference while retaining analytical tractability.
+In these schemes, each AP classifies the UEs into strong and weak groups using large-scale fading thresholds, allowing some spatial degrees of freedom to be allocated for interference suppression while preserving array gain for weaker users.
+Although these schemes provide better interference suppression compared to the MR combining, their user grouping decisions rely on heuristic threshold-based rules.
+As a result, the number of spatial degrees of freedom allocated for interference mitigation is fixed by predetermined thresholds rather than being adapted to the local channel and interference conditions at each AP.
+To address such shortcomings of these zero-forcing schemes, we propose adaptive pilot-level combining strategies that dynamically allocate spatial degrees of freedom based on locally observed channel conditions.
+Specifically, we introduce two generalized combining schemes, referred to as generalized PFZF (G-PFZF) and generalized PWPFZF (G-PWPFZF).
+Since users sharing the same pilot cannot be fully distinguished by an AP, grouping decisions are naturally performed at the pilot-level rather than at the user level.
+Here, each AP independently partitions its set of pilots into strong and weak groups by solving a local sum SE optimization problem.
+Assigning a pilot to the strong group dedicates a spatial degree of freedom (DoF) to mitigate interference via ZF combining, whereas assigning a pilot to the weak group preserves that DoF for signal enhancement through MR combining.
+This pilot-level formulation allows each AP to adaptively operate across a range of combining strategies from pure MR to FZF.
+Performance evaluation of the proposed G-PFZF and G-PWPFZF combining schemes establishes that they significantly outperform PFZF and PWPFZF by enabling adaptive pilot-level allocation of spatial degrees of freedom.
+In addition to efficient local combining, scalable second-stage decoding is essential for practical uplink processing in D-mMIMO systems.
+In [1, 3], the authors have proposed a simple decoding technique, as shown in Fig.
+1(c) which suffers from suboptimal interference suppression performance.
+This motivates the use of centralized large-scale fading decoding (c-LSFD) [8, 11], which exploits large-scale channel statistics to mitigate inter-user interference across APs, as shown in Fig.
+1(b).
+While user-centric or scalable architectures reduce fronthaul signaling by limiting the set of serving APs for each UE, centralized LSFD (c-LSFD) designs still require the exchange of channel statistics among these APs.
+In the optimal c-LSFD scheme [8, 11], each AP forwards both channel statistics and locally detected signals to the CPU, which computes LSFD weights to suppress inter-user interference.
+However, this approach requires the exchange of large-scale channel statistics and matrix inversions whose dimensions grow with the number of UEs.
+Partial LSFD (p-LSFD) has been proposed to reduce the processing complexity by considering only dominant interference terms [8, 12].
+Nevertheless, p-LSFD still requires centralized decoding and matrix inversions at the CPU.
+As the network size and user density increase, centralized operations become increasingly expensive in terms of both fronthaul signaling and computational costs, limiting the scalability of the p-LSFD scheme.
+To overcome these scalability limitations, a decentralized decoding scheme, as shown in Fig.
+1(d), has been proposed in which each AP locally computes decoding weights [13].
+This decentralized heuristic LSFD (h-LSFD) scheme relies on heuristic weight selection and does not provide theoretical guarantees on their performance relative to c-LSFD and p-LSFD.
+Consequently, it remains unclear whether decentralized decoding with h-LSFD can achieve near c-LSFD performance using only locally computable weights without exchanging global channel statistics.
+To address this, we propose a principled d-LSFD scheme where each AP computes decoding weights using only locally available channel statistics through a diagonal approximation of the interference covariance matrix.
+We establish a theoretical connection between c-LSFD and d-LSFD by deriving an explicit SINR bound that quantifies the performance loss due to decentralization.
+Furthermore, we show that the effectiveness of the proposed d-LSFD scheme is intrinsically linked to the structure of the local combining scheme.
+It achieves near-optimal performance for combiners that yield a diagonally dominant large-scale interference covariance matrix (e.g., MR/ZF-based), and larger performance gap for combiners with strong global coupling, such as MMSE.
+Extensive numerical results demonstrate that the combination of the proposed G-PFZF and G-PWPFZF combining schemes and the proposed d-LSFD scheme, achieves significantly higher sum SE and per-user SE, and substantially reduced computational cost and fronthaul signaling compared to any combination of existing combining schemes and LSFD schemes.
+These results highlight the effectiveness of combining adaptive pilot-level interference management with d-LSFD for scalable uplink processing in D-mMIMO systems.
+(a) (b) (c) (d) APs All Data CPU Global Combining Global Estimates APs Local Combining Local data estimates ($\hat{x}_{mt}$) and Channel estimates CPU LSFD Weights & Final Estimates APs Local Combining $\hat{x}_{mt}$ CPU $\hat{x}_{t}=\sum\limits_{m}\hat{x}_{mt}$ APs Local Combining Local Weights ($a_{mt}$) $a_{mt}\hat{x}_{mt}$ CPU $\hat{x}_{t}{=}\sum\limits_{m}a_{mt}\hat{x}_{mt}$
+Figure 1: Uplink processing architectures: (a) Fully centralized, (b) LSFD with coordination, (c) Simple decoding, (d) Decentralized decoding.
 I-A Contributions
-This work addresses some of the above mentioned issues. Our principal contributions are as follows:
-Decentralized Decoding Architecture with Local Weighting: We propose a novel uplink processing architecture as shown in Fig.
-1, that completely eliminates the need for centralized LSFD and its associated fronthaul overhead.
-Central to this architecture is the design of interference-suppressing local weights, computed independently by each AP using only its local channel information.
-These weights are applied to data estimates before transmission, enabling the central processor to operate as a simple aggregator rather than a coordinator.
-This eliminates the need for sharing any channel information with the CPU.
-This transition from a coordinated to a truly decentralized architecture is the key enabler for the practical, large-scale deployment.
-Adaptive Combining Strategy: We introduce generalized PFZF (G-PFZF) and generalized PWPFZF (G-PWPFZF).
-Unlike the conventional PFZF and PWPFZF schemes, where all APs must employ the same scheme, the generalized framework allows each AP to independently adapt its combining strategy.
-Specifically, under G-PFZF an AP may locally switch among PFZF, FZF, or MR; under G-PWPFZF, it may switch among PWPFZF, FZF, or MR, thereby eliminating rigid threshold-based grouping and unlocking per-AP adaptability for improved scalability and performance.
-Novel Decentralized Local Optimization for Pilot Partitioning: We also propose novel decentralized optimization framework where each AP independently partitions pilots (and consequently, their associated UEs) into strong or weak groups based on a local sum SE metric.
-This pilot-level optimization is solved using projected gradient ascent optimization, maintaining low computational cost while maintaining functional equivalence with user-level partitioning.
-Closed-Form Spectral Efficiency Expressions: We derive closed-form expressions for the achievable SE of the proposed system.
-This analysis is essential due to the new statistical properties introduced by pilot-dependent processing and local weighting, and it provides critical insights for system design without the need for Monte Carlo simulation as in MMSE based schemes.
-Performance Analysis: Extensive numerical results demonstrate that our decentralized framework and adaptive grouping strategy achieve good SE with reduced computation cost with the state-of-art baselines.
-The adaptive UE partitioning outperforms traditional fixed-threshold schemes, while the overall design significantly reduces fronthaul overhead and computational costs in comparison to the LSFD framework and fixed threshold partitioning, validating the approach as a practical solution for large-scale deployment.
+We address the twin challenges of designing effective local combining and scalable decoding by developing adaptive pilot-level combining schemes and a decentralized LSFD scheme.
+In particular, our main contributions are as follows:
+Adaptive Combining Schemes: We propose generalized PFZF (G-PFZF) and generalized PWPFZF (G-PWPFZF) schemes that replace heuristic user-level grouping with a pilot-level spatial DoF allocation and provide an explicit mapping between pilot grouping decisions and the number of spatial degrees of freedom used for interference suppression.
+Decentralized LSFD Scheme: We propose a decentralized-LSFD (D-LSFD) scheme where each AP computes decoding weights using only locally available channel statistics.
+We derive locally computable d-LSFD weights and establish an explicit SINR bound that quantifies the performance gap between d-LSFD and c-LSFD.
+The analysis further reveals that d-LSFD is particularly effective for combining schemes that yield diagonally dominant interference covariance matrices, such as ZF- and MR-type combiners.
+Closed-form Expressions: Closed-form SINR and corresponding d-LSFD weights expressions are derived for the proposed combining schemes.
+Performance Analysis: Extensive numerical results demonstrate that the combination of the proposed G-PFZF/G-PWPFZF combining and the proposed d-LSFD outperforms any combination of traditional combining schemes with existing LSFD schemes by a significant margin, while significantly reducing the computation cost and fronthaul overhead.
 Organization: The remainder of the paper is organized as follows.
 Section II describes the system model, along with the channel estimation and data detection procedures.
-Section III presents the proposed combining schemes and derives their corresponding closed-form SE expressions.
+Section III presents the proposed combining schemes and derives their corresponding closed-form expressions.
 Section IV provides a rigorous performance evaluation.
@@
 Notation: Scalars are denoted in italics (e.g., $x$), vectors by bold lowercase letters (e.g., $\mathbf{x}$), and matrices by bold uppercase letters (e.g., $\mathbf{X}$).
-The transpose and Hermitian transpose are denoted by $(\cdot)^{T}$ and $(\cdot)^{H}$, respectively, while the complex conjugate of a scalar is written as $(\cdot)^{*}$.
+The transpose and Hermitian transpose are denoted by $(\cdot)^{T}$ and $(\cdot)^{H}$, respectively, while the complex conjugate is written as $(\cdot)^{*}$.
 The sets of real and complex numbers are denoted by $\mathbb{R}$ and $\mathbb{C}$.
 The expectation operator is denoted by $\mathbb{E}[\cdot]$.
-The identity matrix of size $N$ is denoted by $\mathbf{I}_{N}$.
+The size $N$ identity matrix is denoted by $\mathbf{I}_{N}$.
 A complex Gaussian random vector $\mathbf{x}$ with mean $\bm{\mu}$ and covariance matrix $\mathbf{K}$ is denoted as $\mathbf{x}\sim\mathcal{CN}(\bm{\mu},\mathbf{K})$.
@@
 II System Model
-The system consist of $T$ UEs with a single antenna and $M$ APs, where each AP has $A$ antennas.
-Within the coverage area of interest, there is a uniform distribution of both APs and UEs, and APs work together to service the UEs using the same frequency and time resources.
-Every AP is linked to the CPU by a fronthaul link, so that the network can operate together and process signals.
-The wireless channel is represented using a block fading model, characterized by a coherence block of length $L_{c}$ symbols of which $L_{p}$ pilot symbols are designated for pilot training.
+The system consists of $T$ UEs with a single antenna and $M$ APs, where each AP has $A$ antennas.
+Within the coverage area of interest, there is a uniform distribution of both APs and UEs, and APs serve the UEs using the same frequency and time resources.
+The wireless channel follows a block fading model with a coherence block of length $L_{c}$ symbols, of which $L_{p}$ symbols are allocated for pilot training.
 The small-scale fading is represented by a Rayleigh fading vector $\textbf{h}_{mt}{\in}\mathbb{C}^{A\times 1}$, and the large-scale fading coefficient (LSFC), encompassing both path-loss and shadowing effects, is denoted by $\beta_{mt}$.
 The channel vector between AP $m$ and UE $t$, encompassing both small- and large-scale fading, is defined as $\mathbf{g}_{mt}{=}\beta_{mt}^{1/2}\textbf{h}_{mt}{\in}\mathbb{C}^{A\times 1}$.
-We suppose that $\textbf{h}_{mt}\sim\mathcal{CN}(0,\textbf{I}_{A})$, for all $m$ and $t$, is an independent and identically distributed complex Gaussian random vector, thus $\mathbf{g}_{mt}\sim\mathcal{CN}(0,\beta_{mt}\textbf{I}_{A})$.
+We assume that $\textbf{h}_{mt}{\sim}\mathcal{CN}(0,\textbf{I}_{A})$, for all $m$ and $t$, is an independent and identically distributed (i.i.d) complex Gaussian random vector, thus $\mathbf{g}_{mt}{\sim}\mathcal{CN}(0,\beta_{mt}\textbf{I}_{A})$.
+To enhance scalability, we consider user-centric design [8], where $\mathcal{T}_{m}$ represents the set of UEs served by AP $m$ and $\mathcal{M}_{t}$ represents the set of APs that are serving UE $t$.
 II-A Channel Estimation
-A pilot sequence $\sqrt{L_{p}}\bm{\psi}_{i_{t}}\in\mathbb{C}^{L_{p}\times 1}$ is transmitted by each UE $t$ during the uplink training stage, with $|\bm{\psi}_{i_{t}}|^{2}=1$, where $i_{t}$ is the pilot index of UE $t$.
+A pilot $\sqrt{L_{p}}\bm{\psi}_{i_{t}}{\in}\mathbb{C}^{L_{p}\times 1}$ is transmitted by each UE $t$ during the uplink training stage, with $|\bm{\psi}_{i_{t}}|^{2}{=}1$, where $i_{t}$ is the pilot index of UE $t$.
 At the $m$-th AP, the received signal is:
 $\textbf{y}^{pilot}_{m}=\sum\nolimits_{t=1}^{T}\sqrt{p^{p}_{t}L_{p}}\textbf{g}_{mt}\bm{\psi}_{i_{t}}^{{H}}+\textbf{N}_{m},$
-where $\textbf{N}_{m}\in\mathbb{C}^{A\times L_{p}}$ represents the additive white Gaussian noise (AWGN) matrix characterized by independent and identically distributed complex Gaussian entries.
+where $\textbf{N}_{m}\in\mathbb{C}^{A\times L_{p}}$ represents the additive white Gaussian noise (AWGN) matrix characterized by i.i.d.
+complex Gaussian entries.
 Additionally, $p^{p}_{t}$ represents the normalized pilot power for UE $t$ during pilot transmission.
 The MMSE estimate of the true channel vector $\mathbf{g}_{mt}$ is provided by [1]:
-$\hat{\textbf{g}}_{mt}=\sqrt{p^{p}_{t}L_{p}}\beta_{mt}\theta^{-1}_{mi_{t}}\textbf{y}^{pilot}_{m}\bm{\psi}_{i_{t}},$
+$\hat{\textbf{g}}_{mt}=\sqrt{p^{p}_{t}L_{p}}\beta_{mt}\theta^{-1}_{mi_{t}}\textbf{y}^{pilot}_{m}\bm{\psi}_{i_{t}},$ (1)
 where $\theta_{mi_{t}}{=}\big(\sum_{k=1}^{T}p^{p}_{k}L_{p}\beta_{mk}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+1\big)$.
@@
 Remark 1: The channel estimates of UEs sharing the same pilot sequence are linearly dependent. As a result, an AP cannot fully distinguish among these UEs, leading to residual pilot contamination.
-The matrix of estimated channels at AP $m$ is denoted by $\mathbf{\hat{G}}_{m}{=}[\hat{\mathbf{g}}_{m1},\hat{\mathbf{g}}_{m2},\ldots,\hat{\mathbf{g}}_{mT}]{\in}\mathbb{C}^{A\times T}$.
-Owing to pilot reuse among the UEs, the estimated channel vectors $\hat{\mathbf{g}}_{mt}$ and $\hat{\mathbf{g}}_{mk}$ corresponding to two UEs $t$ and $k$ that share the same pilot sequence become linearly dependent.
-As a result, the columns of $\mathbf{\hat{G}}_{m}$ exhibit linear dependence, rendering the matrix $\mathbf{\hat{G}}_{m}$ rank-deficient.
+The channel estimates matrix at AP $m$ is denoted by $\mathbf{\hat{G}}_{m}{=}[\hat{\mathbf{g}}_{m1},\hat{\mathbf{g}}_{m2},{\ldots},\hat{\mathbf{g}}_{mT}]{\in}\mathbb{C}^{A\times T}$.
+Owing to pilot reuse, the columns of $\mathbf{\hat{G}}_{m}$ exhibit linear dependence, rendering the matrix $\mathbf{\hat{G}}_{m}$ rank-deficient.
 The full rank matrix is designed by removing the linearly dependent columns and by including only those columns that contain linearly independent channel estimates, representing one vector for each pilot sequence.
@@
 $\mathbf{\bar{G}}_{m}=\textbf{y}^{pilot}_{m}\bm{\Psi},\ \ \in\mathbb{C}^{A\times L_{p}}$
-where $\bm{\Psi}=[\bm{\psi}_{1},\bm{\psi}_{2},\ldots,\bm{\psi}_{L_{p}}]$. The channel estimate $\hat{\mathbf{g}}_{mt}$ is rewritten as:
-$\hat{\mathbf{g}}_{mt}=c_{mt}\mathbf{\bar{G}}_{m}\mathbf{e}_{i_{t}},$
-where $\mathbf{e}_{i_{t}}$ is the $i_{t}$-th column of identity matrix $\mathbf{I}_{L_{p}}$.
+where $\bm{\Psi}=[\bm{\psi}_{1},\bm{\psi}_{2},\ldots,\bm{\psi}_{L_{p}}]$.
+Then $\hat{\mathbf{g}}_{mt}=c_{mt}\mathbf{\bar{G}}_{m}\mathbf{e}_{i_{t}}$, where $\mathbf{e}_{i_{t}}$ is the $i_{t}$-th column of identity matrix $\mathbf{I}_{L_{p}}$.
 II-B Uplink Transmission and Spectral Efficiency
 Let $x_{t}$ denote the unit power uplink data signal transmitted by UE $t$.
-The uplink normalised data power corresponding to UE $t$ and additive white Gaussian noise are denoted by $p^{u}_{t}$ and $\textbf{n}_{m}\sim\mathcal{N}_{\mathbb{C}}(0,\textbf{I}_{A})$, respectively.
-The uplink signal received at the $m$-th AP is:
-$\textbf{y}^{u}_{m}=\sum\nolimits_{t=1}^{T}\textbf{g}_{mt}\sqrt{p^{u}_{t}}x_{t}+\textbf{n}_{m}.$
-In decentralized data detection, unlike centralized approach where each AP sends received signal and channel estimates to CPU, each AP performs data detection for its served UEs using a local linear received combining vector.
-Following the proposed decentralized decoding architecture, each AP computes its own interference-suppressing local weight based on the local signal-to-interference-plus-noise ratio (SINR).
+The uplink normalised data power corresponding to UE $t$ and AWGN vector are denoted by $p^{u}_{t}$ and $\textbf{n}_{m}\sim\mathcal{N}_{\mathbb{C}}(0,\textbf{I}_{A})$, respectively.
+The signal received at AP $m$
+$\textbf{y}_{m}=\sum\nolimits_{t=1}^{T}\textbf{g}_{mt}\sqrt{p^{u}_{t}}x_{t}+\textbf{n}_{m}.$
+In decentralized data detection, each AP performs data detection for its served UEs using a local linear received combining vector and also each AP computes its own interference-suppressing local weight based on the local signal-to-interference-plus-noise ratio (SINR).
 These weights are applied to the locally detected data symbols before being sent to the CPU.
-This differs from LSFD, where local data and channel estimates are transmitted to the CPU for centralized processing.
+This differs from c-LSFD, where local data and channel estimates are transmitted to the CPU for centralized processing.
 The locally weighted data symbol $\hat{x}_{mt}$ for UE $t$ at AP $m$ is given by
-$\hat{x}_{mt}=a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{y}^{u}_{m},$ (2)
-where $\mathbf{v}_{m{i_{t}}}\in\mathbb{C}^{A\times 1}$ denotes the local combining vector for UE $t$ at AP $m$, and $a_{mt}$ is the corresponding local weight.
+$\hat{x}_{mt}=a_{mt}\mathbf{v}_{m{t}}^{H}\textbf{y}_{m},$ (3)
+where $\mathbf{v}_{m{t}}\in\mathbb{C}^{A\times 1}$ denotes the local combining vector for UE $t$ at AP $m$, and $a_{mt}$ is the corresponding decoding weight.
 These weighted estimates are forwarded to the CPU, which aggregates them to obtain the final estimate of $x_{t}$:
-$\hat{x}_{t}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{y}^{u}_{m}.$ (3)
-After expanding $\textbf{y}^{u}_{m}$, the (3) can also be rewritten as
-$\begin{split}\hat{x}_{t}&=\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}+\sum_{m=1}^{M}\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\tilde{\textbf{g}}_{mt}\\ &+\sum_{\begin{subarray}{c}k=1\\ k\neq t\end{subarray}}^{T}\sum_{m=1}^{M}\sqrt{p^{u}_{k}}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk}x_{k}+\sum_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m}.\end{split}$ (4)
-Here, the first term is the desired signal for UE $t$, the second term is interference term due to imperfect channel estimation, the third term is interference term due to all other UEs and last term is the noise.
-Based on the decomposition in (4), the achievable uplink spectral efficiency is obtained using the bounding technique and the corresponding Shannon capacity lower bound [17, 18, 16] is given by Theorem 1.
+$\hat{x}_{t}=\sum\nolimits_{m\in\mathcal{M}_{t}}a_{mt}\mathbf{v}_{m{t}}^{H}\textbf{y}_{m}.$ (4)
+After expanding $\textbf{y}_{m}$, the (4) can also be rewritten as
+$\begin{split}\hat{x}_{t}&=\!\!\!\!\!\!\sum_{m\in\mathcal{M}_{t}}\!\!\!\!\!\!\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{t}}^{H}\hat{\textbf{g}}_{mt}x_{t}+\!\!\!\!\!\!\sum_{m\in\mathcal{M}_{t}}\!\!\!\!\!\!\sqrt{p^{u}_{t}}a_{mt}\mathbf{v}_{m{t}}^{H}\tilde{\textbf{g}}_{mt}x_{t}\\ &+\sum_{\begin{subarray}{c}k=1\\ k\neq t\end{subarray}}^{T}\sum_{m\in\mathcal{M}_{t}}\sqrt{p^{u}_{k}}a_{mt}\mathbf{v}_{m{t}}^{H}\textbf{g}_{mk}x_{k}+\!\!\!\!\!\sum_{m\in\mathcal{M}_{t}}\!\!\!a_{mt}\mathbf{v}_{m{t}}^{H}\textbf{n}_{m}.\end{split}$ (5)
+Based on (5), the achievable uplink SE is obtained using the bounding technique and the corresponding Shannon capacity lower bound [15, 16, 17] is is given by Theorem 1.
 Theorem 1.
@@
 $\textsc{SINR}_{t}=\frac{p^{u}_{t}|\textsc{DS}_{t}|^{2}}{\!\!\sum\limits_{k=1}^{T}\!\!p^{u}_{k}\mathbb{E}\left\{|\textsc{UI}_{tk}|^{2}\right\}-p^{u}_{t}|\textsc{DS}_{t}|^{2}+\mathbb{E}\left\{|\textsc{GN}_{t}|^{2}\right\}},$ (7)
-where,
-$\begin{split}&\textsc{DS}_{t}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbb{E}\left\{\mathbf{v}_{m{i_{t}}}^{H}\hat{\textbf{g}}_{mt}\right\},\\ &\textsc{UI}_{tk}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{g}_{mk},\\ &\textsc{GN}_{t}=\sum\nolimits_{m=1}^{M}a_{mt}\mathbf{v}_{m{i_{t}}}^{H}\textbf{n}_{m}.\end{split}$
+$\text{where,}\ \ \ \ \textsc{DS}_{t}=$ $\sum\nolimits_{m\in\mathcal{M}_{t}}a_{mt}\mathbb{E}\left\{\mathbf{v}_{m{t}}^{H}\hat{\textbf{g}}_{mt}\right\},$
+$\textsc{UI}_{tk}=$ $\sum\nolimits_{m\in\mathcal{M}_{t}}a_{mt}\mathbf{v}_{m{t}}^{H}\textbf{g}_{mk},$
+$\textsc{GN}_{t}=$ $\sum\nolimits_{m\in\mathcal{M}_{t}}a_{mt}\mathbf{v}_{m{t}}^{H}\textbf{n}_{m}.$
 The SE lower bound, in (6), is valid regardless of the combining scheme used.
-III Adaptive Combining and Local Weights Analysis
-In this section, we first derive closed-form expressions for the uplink SE and local combining weights for both G-PFZF and G-PWPFZF schemes.
-We also formulate an optimization problem to partition the pilots into strong and weak.
-III-A Generalized Partial Full-Pilot Zero Forcing Combining Scheme
-In [13], the combining vectors for UE $t$ is dependent on $\beta_{mt}$, which is unique to each UE.
-As a result, generating a distinct combining vector for every UE entails a number of multiplications that scales linearly with the total number of UEs.
-To reduce computational cost, we propose G-PFZF combining scheme, where each AP independently designs a combining vector for each of its pilots and pilot-sharing UEs use the same pilot, as APs are not able to distinguish among individual UEs sharing the same pilot.
-The design of combining vectors depends on whether a pilot is strong or weak.
-Thus, each AP categorizes the pilots into two groups, strong and weak based on the local sum SE optimization.
-Unlike traditional PFZF, G-PFZF does not force the APs to select atleast one pilot as strong and allows the APs to choose the combining schemes among PFZF or MR or FZF.
-Remark 2: In our proposed combining schemes, a strong pilot indicates that a ZF combining vector is assigned to that pilot (and its corresponding UEs), while a weak pilot is assigned an MR combining vector.
-For AP $m$, a set of strong pilots is denoted by $\mathcal{S}_{m}$ and a complementary set of weak UEs as $\mathcal{W}_{m}$.
-Let $L_{\mathcal{S}_{m}}$ denote the number of distinct strong pilots used in set $\mathcal{S}_{m}$, and define the corresponding set of pilot indices as $\mathcal{R}_{\mathcal{S}_{m}}{=}\{r_{m,1},r_{m,2},{\dots},r_{m,L_{\mathcal{S}_{m}}}\}$.
-The matrix $\mathbf{\bar{G}}_{m}{\in}\mathbb{C}^{A\times L_{p}}$ contains the estimated channel vectors corresponding to all $L_{p}$ orthogonal pilot sequences, whereas the strong pilots utilize only $L_{\mathcal{S}_{m}}$ of them.
-To isolate the relevant columns of $\mathbf{\bar{G}}_{m}$, we define the selection matrix $\mathbf{E}_{\mathcal{S}_{m}}{=}[\mathbf{e}_{r_{m,1}},{\dots},\mathbf{e}_{r_{m,L_{\mathcal{S}_{m}}}}]{\in}\mathbb{C}^{L_{p}\times L_{\mathcal{S}_{m}}}$, where $\mathbf{e}_{r}$ denotes the $r$-th column of the identity matrix $\mathbf{I}_{L_{p}}$.
-The resulting product $\mathbf{\bar{G}}_{m}\mathbf{E}_{\mathcal{S}_{m}}$ yields a full-rank matrix containing only $L_{\mathcal{S}_{m}}$ linearly independent columns corresponding to the strong pilots.
+First, before establishes the relationship between c-LSFD and d-LSFD, we first propose Lemma 1.
+Lemma 1.
+For any $q\geq 0$, define
+$F(x,q)\triangleq\frac{|1+x|^{2}}{|1+x|^{2}+(q-|x|^{2})},\quad x\in\mathbb{C}.\ \ \text{Then}$
+$\inf_{|x|^{2}\leq q}F(x,q)=\begin{cases}1-q,&0\leq q<1,\\[5.69054pt] 0,&q\geq 1.\end{cases}$
+Moreover, for $0\leq q<1$, the minimum is attained at x=qx=-q, while for $q\geq 1$ the infimum 0 is approached as $x\to-1$.
+Proof: See Appendix -B.
+The following theorem establishes the relationship between c-LSFd and d-LSFD weights and quantifies the performance loss incurred by decentralization.
+Theorem 2.
+: The c-LSFD weights and the d-LSFD weights that maximize the SE lower bound of UE $t$ are given by $\mathbf{a}_{t}^{\mathrm{o}}=p^{u}_{t}\mathbf{J}_{t}^{-1}\mathbf{d}_{t}$ and $\mathbf{a}_{t}^{\mathrm{d}}=p^{u}_{t}\mathbf{D}_{t}^{-1}\mathbf{d}_{t}$ respectively, where,
+$\mathbf{J}_{t}=\sum\nolimits_{k=1}^{T}p^{u}_{k}\mathbb{E}\{\textbf{u}_{t}\textbf{u}^{H}_{t}\}-p^{u}_{t}\textbf{d}_{t}\textbf{d}^{H}_{t}+\textbf{F}_{t}+\mathbf{\tilde{D}}_{t},$
+$\text{where},$
+$\textbf{d}_{t}{=}\!\Big[\!\mathbb{E}\{\!\mathbf{v}_{1{t}}^{H}\mathbf{g}_{1t}\!\},{\ldots},\!\mathbb{E}\{\!\mathbf{v}_{M{t}}^{H}\mathbf{g}_{Mt}\!\}\!\Big]^{T}\!\!\!\!,\ \!\textbf{u}_{tk}{=}\Big[\!\mathbf{v}_{1{t}}^{H}\mathbf{g}_{1k},{\ldots},\!\mathbf{v}_{M{t}}^{H}\mathbf{g}_{Mk}\!\Big]^{T}$
+$\textbf{F}_{t}=\operatorname{diag}\!\Big(\Big[\mathbb{E}\{|\mathbf{v}_{1{t}}^{H}\mathbf{n}_{1}|^{2}\},\;\ldots,\;\mathbb{E}\{|\mathbf{v}_{M{t}}^{H}\mathbf{n}_{M}|^{2}\}\Big]^{T}\Big).$
+Here $\mathbf{D}_{t}{=}\mathrm{diag}(\mathbf{J}_{t})$ and $\mathbf{\tilde{D}}_{t}\in\mathbb{R}^{M\times M}$ is the diagonal matrix with $\mathbf{\tilde{D}}_{t}]_{m,m}=1$, if $m\notin\mathcal{M}_{t}$ and zero otherwise.
+Define $\rho_{t}\triangleq\frac{\|\mathbf{a}_{t}^{\mathrm{d}}-\mathbf{a}_{t}^{\mathrm{o}}\|_{2}}{\|\mathbf{a}_{t}^{\mathrm{o}}\|_{2}}$, and the condition number of $\mathbf{J}_{t}$ as $\kappa_{t}\triangleq\frac{\lambda_{\max}(\mathbf{J}_{t})}{\lambda_{\min}(\mathbf{J}_{t})}$.
+Then the achievable SINR of d-LSFD satisfies
+$\frac{\mathrm{SINR}_{t}(\mathbf{a}_{t}^{\mathrm{d}})}{\mathrm{SINR}_{t}(\mathbf{a}_{t}^{\mathrm{o}})}\;\geq\;\max\!\left\{0,\;1-\kappa_{t}\rho_{t}^{2}\right\}.$ (8)
+Proof: See Appendix -B.
+Remark 2: The bound in (8) shows that the performance gap between d-LSFD and c-LSFD depends on the relative weight mismatch $\rho_{t}$ and the conditioning of $\mathbf{J}_{t}$.
+In practice, the dominant factor is the mismatch $\rho_{t}$, which arises due to local nature of d-LSFD weights.
+Consequently, d-LSFD approximates the centralized solution, and when $\rho_{t}$ is small the resulting SINR loss becomes negligible.
+The following proposition highlights a special case in which the d-LSFD becomes identical to c-LSFD.
+Proposition 1 (Exactness of the bound).
+: If the local combining scheme at each AP is either MR or ZF and each UE is assigned an orthogonal pilot sequence ($L_{p}=T$), then the interference covariance matrix $J_{t}$ becomes diagonal.
+Consequently, the d-LSFD weights coincide with the c-LSFD weights, i.e., $\mathbf{a}_{t}^{d}=\mathbf{a}_{t}^{o}$, and the SINR bound in Theorem 2 holds with equality.
+Proof: See Appendix -C.
+Remark 3: For ZF/MR combining, the approximation of d-LSFD depends on pilot contamination as non-diagonal term of $\mathbf{J}_{t}$ depends on pilot contamination.
+This implies that lesser the pilot contamination better the approximation.
+III Adaptive Combining Analysis
+In this section, we focus on G-PFZF and G-PWPFZF schemes that overcome the limitations of threshold-based UE partitioning.
+For these combining schemes, we derive closed-form expressions for the uplink SE and the corresponding d-LSFD weights.
+III-A Generalized Partial Full-Pilot Zero-Forcing Combining Scheme
+In this work, we propose the G-PFZF combining scheme, in which each AP independently designs combining vectors at the pilot level.
+Since APs cannot distinguish between individual UEs sharing the same pilot, all UEs associated with the same pilot employ a common pilot-level combining vector.
+The pilot-level combining design is governed by a local sum spectral efficiency optimization that determines whether each pilot is treated as strong or weak, thereby explicitly controlling the allocation of spatial degrees of freedom between interference suppression and desired signal enhancement.
+Unlike conventional PFZF, the proposed G-PFZF does not enforce that at least one pilot must be treated as strong, thereby allowing each AP to flexibly operate in MR, PFZF, or full ZF mode depending on local channel and interference condition.
+For AP $m$, let $\mathcal{S}_{m}$ and $\mathcal{W}_{m}$ denote the sets of strong and weak pilots, respectively, and let $L_{\mathcal{S}_{m}}$ be the number of strong pilots.
+Define the set of strong pilot indices as $\mathcal{R}_{\mathcal{S}_{m}}{=}\{r_{m,1},\dots,r_{m,L_{\mathcal{S}_{m}}}\}$.
+Let $\bar{\mathbf{G}}_{m}{\in}\mathbb{C}^{A\times L_{p}}$ collect the estimated channel vectors corresponding to all $L_{p}$ orthogonal pilots, and define the selection matrix $\mathbf{E}_{\mathcal{S}_{m}}{=}[\mathbf{e}_{r_{m,1}},{\dots},\mathbf{e}_{r_{m,L_{\mathcal{S}_{m}}}}]$.
 For a given pilot $i{\in}\mathcal{S}_{m}$, let $j_{mi}{\in}\{1,2,{\dots},L_{\mathcal{S}_{m}}\}$ denote the index of pilot $i$ within $\mathcal{R}_{\mathcal{S}_{m}}$, and define the vector $\bm{\varepsilon}_{j_{mi}}{\in}\mathbb{C}^{L_{\mathcal{S}_{m}}}$ as the $j_{mi}$-th column of the identity matrix $\mathbf{I}_{L_{\mathcal{S}_{m}}}$, such that $\mathbf{E}_{\mathcal{S}_{m}}\bm{\varepsilon}_{j_{mi}}{=}\mathbf{e}_{i}$.
-Following the zero-forcing principle, the pseudo-inverse of this matrix can be expressed as $(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}})^{-1}$.
+The pseudo-inverse of matrix for ZF structure can be expressed as $(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}})^{-1}$.
 To obtain the combining vector associated with a specific pilot, this pseudo-inverse term is multiplied by $\bm{\varepsilon}_{j_{mi}}$ and subsequently normalized.
-Thus, the local ZF combining vector for pilot $i{\in}\mathcal{S}_{m}$ at AP $m$ is given by:
+Thus, the local ZF vector for pilot $i{\in}\mathcal{S}_{m}$ at AP $m$ is given by:
 $\mathbf{v}_{mi}^{\text{LZF}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\bm{\varepsilon}_{j_{mi}}}{\sqrt{\mathbb{E}\big\{\big\|\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\bm{\varepsilon}_{j_{mi}}\big\|^{2}\big\}}},$ (9)
@@
 $\mathbf{v}_{mi}^{\text{MR}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{e}_{i}}{\sqrt{\mathbb{E}\big\{\big\|\mathbf{\bar{G}}_{m}\mathbf{e}_{i}\big\|^{2}\big\}}}=\frac{\mathbf{\bar{G}}_{m}\mathbf{e}_{i}}{\sqrt{A\theta_{mi}}}.$ (10)
-Following [19, 20, 21], the normalization term in (7), for $L_{\mathcal{S}_{m}}\times L_{\mathcal{S}_{m}}$ complex Wishart matrix ($\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}$) with degree of freedom satisfying $A\geq L_{\mathcal{S}_{m}}+1$, can be rewritten as
+Following [18, 19, 20], the normalization term in (9), for $L_{\mathcal{S}_{m}}\times L_{\mathcal{S}_{m}}$ complex Wishart matrix ($\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}$) with DoF satisfying $A\geq L_{\mathcal{S}_{m}}+1$, is
 $\!\!{\mathbb{E}\Big\{\bm{\varepsilon}^{H}_{j_{mi}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}\!)^{-1}\bm{\varepsilon}_{j_{mi}}\Big\}}=\frac{1}{{(A-L_{\mathcal{S}_{m}})\theta_{mi}}},$ (11)
@@
 $(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}=\begin{cases}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t},\end{cases}$ (12)
-$\begin{split}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}})\gamma_{mk}\ \ \text{if }i_{k}=i_{t},\\ 0,\ \ \text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},\ \ \text{if }i_{k}\notin\mathcal{S}_{m}\end{cases}\end{split}$ (11)
+$\begin{split}&\mathbb{E}\Big\{\!\Big|(\!\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{g}}_{mk}\!\Big|^{2}\!\Big\}=\!\!\begin{cases}(A-L_{\mathcal{S}_{m}})\gamma_{mk}\ \ \text{if }i_{k}=i_{t},\\ 0,\ \ \text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},\ \ \text{if }i_{k}\notin\mathcal{S}_{m}\end{cases}\end{split}$ (13)
 Similarly, for any UE $t$ such that $i_{t}\in\mathcal{W}_{m}$
@@
 $\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A+1)\gamma_{mk}&\text{if }i_{k}=i_{t},\\ \gamma_{mk},&\text{if }i_{k}\neq i_{t},\end{cases}$ (15)
-Note 1: From (11) and (13), it is evident that local ZF achieves superior interference suppression compared to MR, since each AP can eliminate interference among UEs with pilots grouped as strong.
-However, this interference mitigation comes at the cost of $L_{\mathcal{S}_{m}}$ degrees of freedom, which reduces the array gain available for desired signal enhancement.
-By using the combining vectors in (7) and (8), the closed-form expression of achievable SE for G-PFZF combining can be computed using Theorem 1 with the corresponding SINR given by
+By using the combining vectors in (9) and (10), the closed-form expression of achievable SE for G-PFZF combining can be computed using Theorem 1 with the corresponding SINR given by (16),
 […]
@@
 $\delta_{mi_{t}}=\begin{cases}1,&\text{if }\ i_{t}\in\mathcal{S}_{m},\\ 0,&\text{otherwise}.\end{cases}$ (17)
-The derivation of the SINR expression is given in Appendix -A.
-Local Weight Design for G-PFZF Scheme: A fundamental limitation of the conventional LSFD architecture is its reliance on centralized optimization [13].
-The optimal LSFD weights are calculated at the CPU using knowledge of global channel statistics, which creates a significant fronthaul signaling overhead and computational burden, thus undermining the scalability of the network.
-In this work, we break this dependency by introducing a novel decentralized decoding architecture where each AP $m$ independently computes a local interference-suppressing weight $a_{mt}$ for each UE $t$, based solely on its local channel information.
-Since these weights are derived from the LSFCs, they evolve much slowly than the small-scale fading components, ensuring long-term stability and minimal operational overhead.
-The primary objectives of this local weight design are twofold: (1) to maximize the contribution of the local estimate to the final SINR at the CPU, and (2) to actively mitigate the dominant interference components, specifically, pilot contamination and inter-user interference, at their source, prior to transmission.
-This approach eliminates the need for the APs to send the channel estimates to the CPU and the CPU to perform any complex optimization, reducing its role to that of a simple aggregator that sums the pre-weighted estimates from all APs.
-The main idea behind designing decoding weights is that it should amplify the data signal received from the APs with strong local desired signal component and suppress those dominated by interference.
-In LSFD design, weights are computed through coordination among the APs, specifically the coherent interference terms.
-In contrast, the proposed decentralized framework eliminates inter-AP coordination, requiring each AP to determine its weight independently.
-Accordingly, an AP with a higher local SINR is given a larger weight than the one with a lower SINR in order to enhance the global SINR.
-For simplicity, the local weight is chosen to be equal to the local SINR.
-The closed-form expression of the local SINR for UE $t$ at AP $m$ can be obtained from (2) by expanding it in the same manner as in (4) and following the derivation procedure outlined in Appendix -A.
-Accordingly, the closed-form of the local SINR or local weight with respect to AP $m$ of UE $t$ is given by:
-$a_{mt}=\text{SINR}^{\text{G-PFZF}}_{mt}=\frac{p^{{u}}_{t}(b_{tt}^{m})^{2}}{\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{{u}}_{k}(b_{kt}^{m})^{2}+W_{mt}}$ (16)
-where
+The closed-form expression of d-LSFD weights for UE $t$ at $m$ for G-PFZF combining scheme following the Theorem 2 for SINR expression in (16) is given by:
+$a_{mt}=\frac{\sqrt{p^{{u}}_{t}}(b_{tt}^{m})}{\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{{u}}_{k}(b_{kt}^{m})^{2}+W_{mt}},$ (18)
+where,
 $b_{kt}^{m}=\sqrt{(A-\delta_{{m}i_{t}}L_{\mathcal{S}_{m}})\gamma_{{m}k}},$
-${W}_{{m}t}=\sum\limits_{k=1}^{T}p^{{u}}_{k}(\beta_{{m}k}-\delta_{{m}i_{t}}\delta_{{m}i_{k}}\gamma_{{m}k})+1.$
-The numerator represents the desired signal strength for the desired UE $t$ at AP $m$, capturing the gain achieved through the chosen combining strategy.
-The denominator is designed to penalize the weight by two main sources of impairment: the coherent interference from pilot contamination and the combined power of non-coherent interference due to channel estimation error and noise.
-By applying these weight to its local estimate, each AP independently produces a soft decision that is already pre-optimized for local interference suppression.
-This design ensures that the signals forwarded to the CPU require no further large-scale fading decoding, thereby completely eliminating the associated fronthaul overhead for channel statistics and the computational cost of centralized weight calculation.
-The result is a radically simplified and highly scalable uplink processing architecture that retains the performance benefits of distributed APs design.
-Adaptive Pilot Grouping for G-PFZF Combining: This subsection details the local optimization framework that each AP employs to independently partition its set of pilots into strong and weak groups.
-This classification dictates the appropriate combining strategy, G-PFZF for strong pilots or MR for weak ones, and crucially, generalizes to pure FZF or pure MR should the optimization deem them to be optimal.
-The basis of this optimization is the local SINR expression given in (16).
-As $L_{\mathcal{S}m}{=}\sum_{i=1}^{L_{p}}\delta_{mi}$ and substituting $L_{\mathcal{S}m}$ into (16) makes the dependency explicit as given in (17).
+${W}_{{m}t}=\sum\nolimits_{k=1}^{T}p^{{u}}_{k}(\beta_{{m}k}-\delta_{{m}i_{t}}\delta_{{m}i_{k}}\gamma_{{m}k})+1.$
+The derivation of the closed-form SINR and d-LSFD expressions are given in Appendix -D.
+Remark 4: In conventional PFZF, UEs are classified as strong or weak using LSFCs thresholds.
+However, because multiple UEs may share the same pilot, UE-level binary grouping does not yield a unique or explicit mapping to the number of spatial degrees of freedom required for interference suppression.
+In contrast, pilot-level binary grouping establishes a explicit mapping, allowing the number of required spatial degrees of freedom to be written explicitly as $L_{\mathcal{S}_{m}}=\sum_{i=1}^{L_{p}}\delta_{mi}$, which is essential for formulating a tractable optimization problem.
+Adaptive Pilot Grouping for G-PFZF Combining: This subsection details the local optimization framework that each AP employs to independently partition its set of pilots into strong and weak.
+The optimization is based on a local SINR metric derived from the global SINR expression in (16).
+Specifically, the local SINR of UE $t$ at AP $m$ is obtained by assuming that only AP $m$ contributes to the reception of UE $t$.
+Moreover, by substituting $L_{\mathcal{S}_{m}}{=}\sum_{i=1}^{L_{p}}\delta_{mi}$, the dependence of the SINR on the pilot-level grouping decision at AP $m$ becomes explicit.
+This explicit representation is essential, as the pilot grouping variables $\{\delta_{mi}\}$ constitute the optimization variables in the proposed framework.
+The resulting local SINR expression is given in (19).
 […]
-This expression reveals a fundamental trade-off governed by the size of the strong pilot set, $L_{\mathcal{S}m}$.
-Allocating a pilot to the strong group ( $\delta_{mi}{=}1$) employs ZF to suppress the non-coherent interference from that pilot’s UEs as well as other strong pilots UEs.
-However, this comes at the cost of a reduced array gain, $(A-L_{\mathcal{S}_{m}})$, for all UEs served by the AP, which diminishes the desired signal power.
-Conversely, classifying a pilot as weak preserves the full array gain but forgoes any interference suppression for its UEs.
-This inherent trade-off between maximizing signal strength and minimizing interference makes the grouping problem non-trivial and renders a fixed, network-wide threshold highly suboptimal.
-To navigate the above mentioned trade-off, AP $m$ independently solves a local optimization problem, which is formulated as
+Allocating a pilot to the strong group ($\delta_{mi}{=}1$) assigns a ZF combining vector to that pilot, thereby suppressing coherent interference from UEs sharing the same pilot as well as from UEs associated with other strong pilots.
+This interference suppression is achieved by consuming one spatial DoF per strong pilot, which reduces the available array gain to $(A{-}L_{\mathcal{S}_{m}})$ for all UEs served by AP $m$.
+Conversely, classifying a pilot as weak preserves the full array gain but provides no interference suppression for the corresponding UEs.
+This fundamental trade-off between array gain and interference mitigation makes the pilot grouping problem non-trivial and renders fixed, network-wide threshold-based grouping highly suboptimal.
+To navigate this trade-off, AP $m$ independently solves a local optimization problem, which is formulated as
 $\underset{\bm{\delta}_{m}}{\max~}$ $\ \sum\limits_{t=1}^{T}\log_{2}\left(1+\frac{S_{mt}}{I_{mt}}\right),$ (20a)
@@
 Also, intuitively, this would otherwise lead to invalid negative values for the desired signal power and coherent interference in the SINR expression.
-The optimization problem in (18) is a Mixed-Integer Non-Linear Program (MINLP), a class of problems known for its exponential computational cost.
+The problem in (20) is a Mixed-Integer Non-Linear Program (MINLP), a class of problems known for its exponential computational cost.
 To develop a more tractable formulation, we begin by relaxing the binary constraint (20b) into a continuous one:
 $0\leq\delta_{mi}\leq 1,\ \forall i.$ (21)
-To ensure that the solutions of this new NLP satisfy the original binary requirement ($\delta_{mi}\in{0,1}$), we introduce the following complementary constraint:
+To ensure original binary requirement ($\delta_{mi}\in{0,1}$), we introduce the complementary constraint:
 $\delta_{mi}-\delta^{2}_{mi}\leq 0,\ \forall i.$ (22)
@@
 […]
-This problem can be solved using non-convex solvers or the successive convex approximation optimization method.
-Both of these are computationally expensive for large network deployment as non-convex solvers.
+Although the problem can be solved using generic non-convex optimization solvers, such approaches become computationally expensive for large networks.
 Therefore, to solve this problem with low computation cost, we employ gradient based proximal gradient ascent (PGA) method [21, 22].
-We first reformulate the problem to be unconstrained except for the box constraint (19).
-The reformulated problem is:
+We first reformulate the problem as:
 $\underset{\bm{\delta}_{m}}{\max~}$ $f(\bm{\delta}_{m}),$ (24a)
 […]
+$\begin{split}f(\bm{\delta}_{m})&=\!\!\sum\limits_{t=1}^{T}\log_{2}(1+\frac{S_{mt}}{I_{mt}})-\chi\Bigg(\!\lambda_{1}\!\sum_{i=1}^{L_{p}}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)^{2}\\ &+\lambda_{2}\max\!\Big(0,\sum_{i=1}^{L_{p}}\delta_{mi}-A+1\Big)^{2}\Bigg),\end{split}$
+where $\chi$ is a penalty parameter, and $\lambda_{1}$ and $\lambda_{2}$ are weights.
+Lemma 2.
+: For a sufficiently large penalty $\chi^{*}$, the solution to Problem (24) is equivalent to the solution of Problem (23).
+Proof: As $\chi{\rightarrow}+\infty$, the penalty terms $\chi\lambda_{1}$ and $\chi\lambda_{2}$ enforce that any constraint violation is driven to zero for the solution to remain finite.
+Given that the original problem’s feasible set is bounded (as argued in [23, Proposition 1]), there exists a finite $\chi^{*}$ such that the solutions coincide.
+For practical consideration, the penalty function should be sufficiently small (${\leq}10^{-3}$), for some value of $\chi$.
+In our numerical simulations, we initialize $\chi{=}1$ with $\lambda_{1}{=}5$ and $\lambda_{2}{=}10$ to ensure sufficiently small penalty function [22].
+To solve Problem (24) using the PGA method at each AP independently, we follow the steps outlined in Algorithm 1.
+We initialize $\bm{\delta}_{m}$ and update it along with the gradient ascent direction of $f(\bm{{\delta}}^{(j)}_{m})$ with step size $\alpha$, followed by projection onto $[0,1]$ to satisfy the box constraint (21):
+$\begin{split}\bm{{\delta}}^{(j+1)}_{m}&=\text{proj}_{[0,1]}\!\left(\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right).\end{split}$ (25)
+The components of gradient $\nabla f(\bm{\delta}_{m})$ can be written as
+$\begin{split}\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}=&\sum_{t=1}^{T}\frac{I_{mt}}{I_{mt}+S_{mt}}\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}\\ &-\chi\Bigg(\!2\lambda_{1}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)\Big(1-2\delta_{mi}\Big)\\ &+2\lambda_{2}\max\!\Big(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\Big)\!\Bigg),\end{split}$ (26)
 where
-$\begin{split}f(\bm{\delta}_{m})&=\!\!\sum\limits_{t=1}^{T}\log_{2}(1+\frac{S_{mt}}{I_{mt}})-\chi\Bigg(\!\lambda_{1}\!\sum_{i=1}^{L_{p}}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)^{2}+\lambda_{2}\max\!\Big(0,\sum_{i=1}^{L_{p}}\delta_{mi}-A+1\Big)^{2}\Bigg),\end{split}$
-where $\chi$ is a penalty parameter, and $\lambda_{1}$ and $\lambda_{2}$ are penalty weights.
-Lemma 1.
-: For a sufficiently large penalty parameter $\chi^{*}$, the solution to Problem (22) is equivalent to the solution of Problem (21).
-Proof: As $\chi\rightarrow+\infty$, the penalty terms $\chi\lambda_{1}$ and $\chi\lambda_{2}$ enforce that any constraint violation is driven to zero for the solution to remain finite.
-Given that the original problem’s feasible set is bounded (as argued in [24, Proposition 1]), there exists a finite $\chi^{*}$ such that the solutions coincide.
-To solve Problem (22) using the PGA method at each AP independently, we follow the steps outlined in Algorithm 1.
-Since the objective function (22a) is smooth and differentiable, the proximal step reduces to a projection.
-We initialize $\bm{\delta}_{m}$ and update it along with the gradient ascent direction of $f(\bm{{\delta}}^{(j)}_{m})$ with step size $\alpha$, followed by projection onto $[0,1]$ to satisfy the box constraint (19):
-$\begin{split}\bm{{\delta}}^{(j+1)}_{m}&=\text{proj}_{[0,1]}\!\left(\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)=\min\!\left(1,\max\!\left(0,\bm{{\delta}}_{m}^{(j)}+\alpha\nabla f(\bm{{\delta}}_{m}^{(j)})\right)\right).\end{split}$ (23)
-The components of gradient $\nabla f(\bm{\delta}_{m})$ can be written as
-$\begin{split}\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}=&\sum_{t=1}^{T}\frac{I_{mt}}{I_{mt}+S_{mt}}\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}-\chi\Bigg(\!2\lambda_{1}\max\!\Big(0,\delta_{mi}-\delta^{2}_{mi}\Big)\Big(1-2\delta_{mi}\Big)\\ &+2\lambda_{2}\max\!\Big(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\Big)\!\Bigg),\end{split}$ (24)
-where
-$\frac{\partial S_{mt}}{\partial\delta_{mi}}=\begin{cases}-p^{u}_{t}\Big(1+\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\delta_{mj}\Big)\gamma_{mt},&\text{if }\ i_{t}=i,\\ -p^{u}_{t}\delta_{mi_{t}}\gamma_{mt},&\text{if }\ i_{t}\neq i,\end{cases}$ (25)
-$\frac{\partial I_{mt}}{\partial\delta_{mi}}=\begin{cases}-\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!p^{u}_{k}\ \Big(1+\!\!\!\!\!\!\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\!\!\!\!\!\!\!\delta_{mj}\Big)\gamma_{m}-\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}}\!\!\!\!p^{u}_{k}\gamma_{mk},\ \ \text{if }i_{t}=i,\\ \vskip 5.69054pt-\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!p^{u}_{k}\delta_{mi_{t}}\gamma_{mk}-\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}}\!\!\!\!p^{u}_{k}\delta_{mi_{t}}\gamma_{mk},\ \ \text{if }i_{t}\neq i,\end{cases}$ (26)
+$\frac{\partial S_{mt}}{\partial\delta_{mi}}=\begin{cases}-p^{u}_{t}\Big(1+\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\delta_{mj}\Big)\gamma_{mt},&\text{if }\ i_{t}=i,\\ -p^{u}_{t}\delta_{mi_{t}}\gamma_{mt},&\text{if }\ i_{t}\neq i,\end{cases}$
+$\frac{\partial I_{mt}}{\partial\delta_{mi}}=\begin{cases}-\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!p^{u}_{k}\ \Big(1+\!\!\!\!\!\!\sum\limits_{j=1,j\neq i_{t}}^{L_{p}}\!\!\!\!\!\!\!\delta_{mj}\Big)\gamma_{m}-\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}}\!\!\!\!p^{u}_{k}\gamma_{mk},\ \ \text{if }i_{t}=i,\\ \vskip 5.69054pt-\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!p^{u}_{k}\delta_{mi_{t}}\gamma_{mk}-\!\!\!\!\!\sum\limits_{k\in\mathcal{P}_{i_{t}}}\!\!\!\!p^{u}_{k}\delta_{mi_{t}}\gamma_{mk},\ \ \text{if }i_{t}\neq i,\end{cases}$
 1:for all $m\in\mathcal{M}$ independently do
@@
 7:   until $\Big|\frac{f(\bm{{\delta}}^{(j)}_{m})-f(\bm{{\delta}}^{(j-1)}_{m})}{f(\bm{{\delta}}^{(j-1)}_{m})}\leq\epsilon\Big|$
-8:   Update $\chi=\chi\times\Delta$
-9:   Update $k=k+1$
-10:   Update $s^{(k)}=f(\bm{{\delta}}^{(j)}_{m})$
-11:  until $\big|\frac{s^{(k)}-s^{(k-1)}}{s^{(k-1)}}\leq\epsilon\big|$
-12:end for
+8:   Update $\chi=\chi\times\Delta$, $k=k+1$ and $s^{(k)}=f(\bm{{\delta}}^{(j)}_{m})$
+9:  until $\big|\frac{s^{(k)}-s^{(k-1)}}{s^{(k-1)}}\leq\epsilon\big|$
+10:end for
 Algorithm 1 Proximal Gradient Ascent (PGA) Method
-Convergence and Computation Cost Analysis: Since the feasible set defined by the box constraint (19) is bounded and the gradient function $\nabla f(\bm{\delta}_{m})$ is Lipschitz continuous (proved in Appendix -B) with constant $L>0$, one should choose $\alpha\in(0,1/L]$ to ensure the convergence of the PGA method [22, 23].
-In numerical simulations, small step size parameter $\alpha$ has been observed to achieve convergence for Algorithm 1.
-At each AP, the computation cost of Algorithm 1 in each iteration depends upon the gradient computation step.
-The computation cost of calculating $f(\bm{\delta}_{m})$ is $\mathcal{O}(T^{2})$.
-Thus, the cost of calculating $\nabla f(\bm{{\delta}}_{m})$ is also $\mathcal{O}(T^{2})$.
-Since we have only $L_{p}$ variables, it converges in a few iterations only.
-Note 2: To reduce the computational cost, we use pilot-based grouping instead of UE-based grouping as in [19, 13].
-In UE-based grouping, each AP would have $T$ binary variables (one per UE), leading to $TM$ total variables across the network.
-In the proposed pilot-based grouping, each AP has $L_{p}$ binary variables, leading to $ML_{p}$ total variables.
-Since $L_{p}\ll T$ in dense deployments, this significantly reduces the computational cost.
-III-B Generalized Protected Weak Partial Full-Pilot Zero Forcing Combining Scheme
-In the G-PFZF scheme, the combiner for the UEs assigned to a strong pilot is designed only to suppress interference from other UEs that are also on strong pilots.
-This intra-group suppression leaves UEs on weak pilots vulnerable to dominant interference from the strong UEs.
-The proposed G-PWPFZF scheme introduces a more comprehensive, protective strategy.
+III-B Generalized Protected Weak Partial Full-Pilot Zero-Forcing Combining Scheme
 In this framework, the combiners for all UEs, whether assigned to a strong or weak pilot, are designed to actively suppress interference from all UEs that have a strong pilot.
@@
 $\mathbf{B}_{m}=\mathbf{I}_{A}-\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}}(\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m}\mathbf{\bar{G}}_{m}\mathbf{E}_{{\mathcal{S}_{m}}})^{-1}\mathbf{E}^{H}_{{\mathcal{S}_{m}}}\mathbf{\bar{G}}^{H}_{m},$
-This projection ensures that the combiner for a weak pilot is orthogonal to the channel estimates of all strong-pilot UEs, thereby nullifying the dominant interference from that group.
 The expected value of the effective channel gain for UE $t$ such that $i_{t}\in\mathcal{W}_{m}$
 $\mathbb{E}\{(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}{\mathbf{\hat{g}}_{mk}}\}=\begin{cases}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}&\text{if }i_{k}=i_{t},\\ 0,&\text{if }i_{k}\neq i_{t}.\end{cases}$ (28)
-Furthermore
-$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}}+1)\gamma_{mk}\ \ \text{if }i_{k}=i_{t},\\ 0,\ \ \text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},\ \ \text{if }i_{k}\neq i_{t},\end{cases}$ (29)
-These equations confirm that the projected MR (PMR) combiner successfully nullifies interference from all UEs sharing strong pilots $i_{k}\in\mathcal{S}_{m}$.
-This design effectively mitigates the most significant intra-network interference, enhancing the performance of the UEs with weaker channel conditions.
-By using the combining vector (7) for strong pilot and (27), the closed-form expression for achievable SE for G-PWPFZF combining, as in Theorem 1, can be obtained with the SINR given by
+$\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{g}}_{mk}\Big|^{2}\Big\}=\begin{cases}(A-L_{\mathcal{S}_{m}}+1)\gamma_{mk}\ \ \text{if }i_{k}=i_{t},\\ 0,\ \ \text{if }i_{k}\neq i_{t}\ \text{and}\ i_{k}\in\mathcal{S}_{m},\\ \gamma_{mk},\ \ \text{if }i_{k}\neq i_{t},\end{cases}$
+These equations confirm that the PMR combiner successfully nullifies interference from all UEs sharing strong pilots $i_{k}\in\mathcal{S}_{m}$.
+By using the combining vector (9) for strong pilot and (27), the closed-form expression for achievable SE for G-PWPFZF combining, as in Theorem 1, can be obtained with the SINR given by (29).
 […]
-The derivation of the SINR expression is given in Appendix -C.
-Local Weights Design for G-PWPFZF Scheme: The design of local weights for the G-PWPFZF scheme follows a similar design to that of the G-PFZF scheme, aiming to maximize the local contribution to the final SINR while suppressing interference in a fully decentralized manner.
-Here too the local weights are the same as the local SINR.
-The closed-form expression of the local SINR for UE $t$ at AP $m$ can be obtained from (2) by expanding it in the same manner as in (4) and following the derivation procedure outlined in Appendix -C.
-Accordingly, the closed-form of the local SINR or local weight is given by:
-$a_{mt}=\text{SINR}^{\text{G-PWPFZF}}_{mt}=\frac{p^{{u}}_{t}(b_{tt}^{m})^{2}}{\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{{u}}_{k}(b_{kt}^{m})^{2}+W_{mt}}$ (31)
-where
+The corresponding closed-form expression of d-LSFD weights for UE $t$ at $m$ for G-PWPFZF combining is:
+$a_{mt}=\frac{\sqrt{p^{{u}}_{t}}b_{tt}^{m}}{\sum\limits_{k\in\mathcal{P}_{i_{t}}\backslash\{t\}}p^{{u}}_{k}(b_{kt}^{m})^{2}+W_{mt}},\ \ \ \text{where},$ (30)
 $b_{kt}^{m}=\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{{m}k}},$
-${W}_{{m}t}=\sum\limits_{k=1}^{T}p^{{u}}_{k}(\beta_{{m}k}-\delta_{{m}i_{k}}\gamma_{{m}k})+1.$
-By employing these locally computed weights, each AP independently scales its data estimates to prioritize signals with high desired gain and low interference.
-This process effectively embeds interference suppression functionality at the source, thereby completely eliminating the need for centralized LSFD at the CPU and the associated fronthaul overhead for channel statistics.
+${W}_{{m}t}=\sum\nolimits_{k=1}^{T}p^{{u}}_{k}(\beta_{{m}k}-\delta_{{m}i_{k}}\gamma_{{m}k})+1.$
+The derivation of the closed-form SINR and d-LSFD expressions are given in Appendix -E.
 Adaptive Pilot Grouping for G-PWPFZF Combining: This subsection discusses the local optimization framework for pilot grouping in the G-PWPFZF scheme.
@@
 The sole distinction lies in the expression for the local SINR, which forms the foundation of the utility function $f(\bm{\delta}_{m})$ for pilot-grouping.
-The local SINR for UE $t$ at AP $m$ for G-PWPFZF scheme is given in (31).
-After substituting $L_{\mathcal{S}m}=\sum_{i=1}^{L_{p}}\delta_{mi}$ in (31):
+The local SINR for UE $t$ at AP $m$ for G-PWPFZF scheme is $\text{SINR}^{\text{G-PWPFZF}}_{mt}=\frac{S_{mt}}{I_{mt}}$, where
 $S_{mt}=p^{u}_{t}\Big(A-\sum\limits_{i=1}^{L_{p}}\delta_{mi}\Big)\gamma_{mt}$
@@
 These expressions replace their G-PFZF counterparts in the gradient calculation step of Algorithm 1. All other steps of the algorithm proceed identically.
+Convergence and Computation Cost Analysis of PGA: Since the feasible set defined by the box constraint (21) is bounded and the gradient function $\nabla f(\bm{\delta}_{m})$ is Lipschitz continuous (proved in Appendix -F) with constant L>0L{>}0, one should choose $\alpha{\in}(0,1/L]$ to ensure the convergence of the PGA method [21, 22].
+Step size $\alpha$ is selected using backtracking line-search to achieve convergence for Algorithm 1.
+At each AP, the computation cost of Algorithm 1 in each iteration depends upon the gradient computation step and is given by $\mathcal{O}(TL_{p})$.
+For both proposed schemes, the average number of PGA iterations per AP is approximately 7, regardless of whether the number of UEs is 50 or 100.
+This indicates that the convergence speed of PGA is largely insensitive to UE density.
+The average runtime per AP on a standard laptop equipped with an Intel Core Ultra 7 processor and 16 GB RAM is approximately 0.01 s for 50 UEs and increases to about 0.02–0.03 s for 100 UEs.
+This increase is mainly due to the higher cost of objective evaluation rather than slower convergence.
+These results confirm that the proposed PGA-based pilot grouping is computationally efficient and scalable to dense UE deployments.
 IV Performance Evaluation
-TABLE I: Baseline Simulation Parameters
-Parameter Value
-Number of APs ($M$) 100
-Antennas per AP ($A$) 8
-Number of UEs ($T$) 100
-Channel bandwidth ($B$) 20 MHz
-Uplink pilot symbols ($L_{p}$) 7
-Coherence block length ($L_{c}$) 200 symbols
-Maximum UE transmit power ($p_{\text{max}}$) 100 mW
-Shadow fading standard deviation 8 dB
-This section demonstrates the gains of the our proposed decentralized decoding and combining schemes in terms of the sum SE, per user SE and computation cost.
-We consider a network with $M$ APs and $T$ UEs, uniformly distributed in a square area of size $1\times 1$ km2.
-The relevant system parameters are mentioned in Table I.
-The LSFCs are generated according to the model in [3], incorporating shadow fading.
-Pilots are assigned using the scalable pilot assignment strategy from [3], ensuring efficient reuse under limited pilot resources.
-All APs are assumed to operate at full transmit power during both pilot training and data transmission.
-For the baseline PFZF and PWPFZF schemes, UE grouping is performed using a $90\%$ of total LSFCs received at the AP, are identified using numerical simulation, to provide the best performance.
+We consider a network with $M=100$ APs with $A=8$ antennas each and $T=100$ UEs, uniformly distributed in a square area of size $1\times 1$ km2.
+The LSFCs are generated according to the model in [3], incorporating shadow fading with standard deviation of 8 dB.
+Pilots are assigned using the scalable pilot assignment strategy from [3].
+All UEs are assumed to operate at full transmit power of 100 mW.
+For the baseline PFZF and PWPFZF schemes, UE grouping is performed using $90\%$ of total LSFCs received at the AP, as identified via numerical simulation, to provide the best performance.
+Also $\mathcal{K}_{t}$ denotes the set of UEs that have atleast one serving AP in common with UE $t$.
 To ensure statistical reliability, all reported results are averaged over 500 independent network realizations.
 IV-A Computation cost and Fronthaul cost
-In this subsection, we analyze the computational and fronthaul signaling costs of our proposed decentralized decoding and combining schemes.
-The costs are computed following the approach outlined in [16, 19, 15, 10], where only multiplication and division operations are counted per coherence block, as these dominate the computational load.
-TABLE II: Decoding Computation and Fronthaul Overhead per UE per Coherence Block
-Decoding Type Fronthaul Cost Computation Cost
-LSFD [16, 15] $L_{u}M+\frac{3MT+M}{2}$ $\left(\frac{M^{2}+M}{2}\right)T+\frac{M^{3}-M}{3}+M^{2}$
-Proposed Decoding $L_{u}M$ $\left(\frac{M^{2}+M}{2}\right)T+M+M^{2}$
-Table II compares the fronthaul and computational costs for optimal LSFD and the proposed decentralized decoding architecture.
-In conventional LSFD, each AP must transmit both channel estimates and local data estimates to the CPU, resulting in substantial fronthaul overhead of $L_{u}M+\frac{3MT+M}{2}$ complex scalars per UE per coherence block.
-Furthermore, LSFD requires computationally expensive matrix operations such as matrix inversion at the CPU that scale as $\mathcal{O}(TM^{2}+M^{3})$ per UE, making it impractical for large-scale network deployments.
-The proposed decentralized decoding architecture eliminates these bottlenecks by avoiding centralized LSFD entirely.
-Each AP independently computes and applies interference-suppressing local weights, reducing the fronthaul load to only $L_{u}M$ complex scalars.
-The computational cost is similarly streamlined to $\mathcal{O}(TM^{2})$ by eliminating the cubic $M^{3}$ term associated with large matrix inversions required in the LSFD architecture due to centralized AP coordination.
-This combination of reduced fronthaul overhead and scalable computation enables practical deployment in dense network scenarios.
-Figure 2: Comparison of computational and fronthaul costs for different decoding and combining schemes.
+We analyze the computational and fronthaul signaling costs of our proposed d-LSFD and combining schemes.
+The costs are computed following the approach outlined in [17, 18, 12, 8], where only multiplication and division operations are counted, as these dominate the computational load.
+TABLE I: Decoding Computation and Fronthaul Overhead per Coherence Block
+Decoding Fronthaul Computation Cost
+c-LSFD [17] $L_{u}\sum_{m=1}^{M}|\mathcal{T}_{m}|+2TM$ $\sum_{t=1}^{T}|\mathcal{M}_{t}|^{2}\frac{|\mathcal{P}_{i_{t}}|+2}{2}+\sum_{t=1}^{T}|\mathcal{M}_{t}|T+\sum_{t=1}^{T}\frac{|\mathcal{M}_{t}|^{3}-|\mathcal{M}_{t}|}{3}$
+p-LSFD [17] $L_{u}\sum_{m=1}^{M}|\mathcal{T}_{m}|+2TM$ $\sum_{t=1}^{T}|\mathcal{M}_{t}|^{2}\frac{|\mathcal{P}_{i_{t}}|+2}{2}+\sum_{t=1}^{T}|\mathcal{M}_{t}||\mathcal{K}_{t}|+\sum_{t=1}^{T}\frac{|\mathcal{M}_{t}|^{3}-|\mathcal{M}_{t}|}{3}$
+d-LSFD $L_{u}\sum_{m=1}^{M}|\mathcal{T}_{m}|$ $\sum_{t=1}^{T}|\mathcal{M}_{t}|T+\sum_{t=1}^{T}|\mathcal{M}_{t}|+\sum_{t=1}^{T}|\mathcal{M}_{t}|^{2}$
+Table I compares the fronthaul and computational costs for c-LSFD, p-LSFD and proposed d-LSFD.
+In c-LSFD and p-LSFD, each AP must transmit both channel estimates and local data estimates to the CPU, resulting in substantial fronthaul overhead.
+Furthermore, c-LSFD and p-LSFD require computationally expensive matrix operations such as matrix inversion at the CPU.
+The d-LSFD eliminates these bottlenecks by avoiding centralized decoding entirely, reducing the fronthaul load for sending data estimates only.
+The computational cost is similarly streamlined by eliminating the matrix inversions required in c-LSFD.
+Figure 2: Comparison of computational and fronthaul costs for different decoding schemes.
 As demonstrated in Fig.
-2a, the proposed decentralized decoding weights computation provides substantial computation cost reduction compared to LSFD, particularly as the number of UEs increases.
+2a, the proposed d-LSFD weights computation provides substantial computation cost reduction compared to c-LSFD, particularly as the number of UEs increases.
 Similarly, Fig.
 2b shows that eliminating channel estimate transmission reduces fronthaul overhead.
-TABLE III: Computational cost per AP per Coherence Block
+TABLE II: Computational cost per AP per Coherence Block
 Scheme Combining Vector Computation
-PFZF[19] $\frac{3L_{\mathcal{S}_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}AT$
-G-PFZF $\frac{3L_{\mathcal{S}_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}AL_{p}$
-PWPFZF[19] $\frac{3L_{S_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}2(L_{p}{-}L_{\mathcal{S}_{l}})L_{\mathcal{S}_{l}}A{+}AT$
-G-PWPFZF $\frac{3L_{S_{l}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{l}}A}{2}{+}\frac{L_{\mathcal{S}_{l}}^{3}{-}L_{\mathcal{S}_{l}}}{3}{+}2(L_{p}{-}L_{\mathcal{S}_{l}})L_{\mathcal{S}_{l}}A{+}AL_{p}$
-The computational cost per AP per coherence block for calculating the combining vectors is detailed in Table III.
-The primary distinction between the traditional and proposed generalized schemes lies in their scaling behavior relative to the number of users.
-The traditional PFZF and PWPFZF schemes exhibit computational costs that scale linearly with the total number of user equipments, $T$, as in the $AT$ term for PFZF and PWPFZF.
-This dependence creates a significant computational bottleneck in densed networks with massive connectivity.
-In contrast, the proposed G-PFZF and G-PWPFZF schemes achieve a fundamental scalability advantage by scaling with the number of pilots, $L_{p}$, rather than the number of UEs.
-This is reflected in the $AL_{p}$ term for G-PFZF and G-PWPFZF, which is a fixed system parameter independent of the user population.
+PFZF[18] $\underbrace{\frac{3L_{\mathcal{S}_{m}}^{2}A}{2}{+}\frac{L_{\mathcal{S}_{m}}A}{2}{+}\frac{L_{\mathcal{S}_{m}}^{3}{-}L_{\mathcal{S}_{m}}}{3}}_{C_{m}}{+}A|\mathcal{T}_{m}|$
+G-PFZF $C_{m}{+}AL_{p}$
+PWPFZF[18] $C_{m}{+}2(L_{p}{-}L_{\mathcal{S}_{m}})L_{\mathcal{S}_{m}}A{+}A|\mathcal{T}_{m}|$
+G-PWPFZF $C_{m}{+}2(L_{p}{-}L_{\mathcal{S}_{m}})L_{\mathcal{S}_{m}}A{+}AL_{p}$
+The PFZF and PWPFZF schemes exhibit costs that scale linearly with the total number of serving UEs, $|\mathcal{T}_{m}|$, as shown in Table II.
+In contrast, the proposed G-PFZF and G-PWPFZF schemes cost scales only with the number of pilots, $L_{p}$.
+IV-B Numerical Simulations
+Figure 3: Empirical distribution versus SINR lower bound in Theorem 2 across different network realizations.
+First, we empirically evaluate the tightness of the bound presented in Theorem 2.
+We also numerically verify Proposition 1 and demonstrate that, for ZF/MR-type combining, the bound is tight and depends on the level of pilot contamination.
 Fig.
-2c further validates the scalability benefits of the generalized schemes, where computational cost saturates once the number of UEs exceeds the available pilot sequences, unlike traditional schemes that continue to scale linearly with user density.
-IV-B Performance Evaluation
+3 shows the empirical distribution of the bound $\max(0,1-\kappa_{t}\rho_{t}^{2})$ for pilot lengths $L_{p}\in{7,15,25,T}$ and for different local combining schemes.
+For ZF and MR combining, increasing the pilot length shifts the distribution toward the interval 0.80.811, indicating that the d-LSFD solution closely approximates the c-LSFD in most realizations.
+When $L_{p}=T$, the bound becomes exactly one for both ZF and MR combining, consistent with Proposition 1.
+For these schemes, the off-diagonal entries of $\mathbf{J}_{t}$ are dependent on pilot contamination, which decreases with increase in pilot length, allowing the d-LSFD weights to remain well aligned with the centralized solution.
+In contrast, for the MMSE scheme, the bound is concentrated near the interval 00.20.2 even when pilot contamination is removed.
+This behavior occurs because the MMSE combining vector depends on the channel estimates of all UEs, regardless of whether they share the same pilot sequence.
+As a result, the off-diagonal entries of the interference matrix $\mathbf{J}_{t}$ remain significant, and discarding them in the d-LSFD design leads to large deviations from the centralized solution.
+These results confirm that the proposed d-LSFD scheme is highly effective for ZF/MR-type combining schemes but unsuitable for combiners with strong global coupling, such as MMSE.
 Figure 4: Distribution of strong pilot decisions.
 Fig.
-3 shows the distribution of strong pilot decisions ($L_{\mathcal{S}_{m}}$) across all APs.
-The baseline PFZF and PWPFZF schemes exhibit a pronounced peak at $L_{\mathcal{S}_{m}}{=}6$-77, revealing their rigid threshold-based strategy that sacrifices excessive degrees of freedom for interference suppression regardless of local conditions.
-In striking contrast, the proposed G-PFZF and G-PWPFZF schemes show fundamentally different behavior: G-PFZF peaks at 2-4 strong pilots while G-PWPFZF shows an even more conservative distribution peaking at 1-3.
+4 shows that the baseline PFZF and PWPFZF schemes exhibit a pronounced peak at $L_{\mathcal{S}_{m}}{=}6$-77, revealing their rigid threshold-based strategy that sacrifices excessive degrees of freedom for interference suppression regardless of local conditions.
+In striking contrast, the proposed G-PFZF and G-PWPFZF schemes show fundamentally different behavior: G-PFZF and G-PWPFZF peaks at 1-4 strong pilots.
 This efficient allocation preserves spatial resources for signal enhancement.
 Most notably, the significant non-zero count at $L_{\mathcal{S}_{m}}{=}0$ demonstrates our algorithm’s ability to adaptively default to simple MR combining when it is the optimal strategy for a given AP’s local conditions.
-This demonstrates a seamless, adaptive switching between local zero-forcing and local MR processing that is impossible for the threshold-bound baselines.
+This demonstrates a seamless, adaptive switching between ZF and MR combining.
+Figure 5: The uplink sum SE comparison for various decoding schemes.
+Figure 6: The uplink sum SE comparison for various decoding schemes.
+Figs.
+5 and 6 illustrate the sum SE versus the number of UEs for the proposed G-PFZF and G-PWPFZF combining schemes, respectively, comparing c-LSFD, the proposed d-LSFD, and h-LSFD [13].
+For both G-PFZF and G-PWPFZF, the performance gap between d-LSFD and c-LSFD remains consistently within 1–2% across all UE densities.
+This confirms that the proposed locally computable LSFD weights achieve near-optimal performance within the class of decentralized linear decoders, while completely eliminating the need for network-wide coordination and centralized matrix inversions.
+In contrast, h-LSFD exhibits a substantially larger performance loss of approximately 10–12%, highlighting the benefit of the proposed principled weight design.
+Fig.
+6 additionally includes LP-MMSE combining for comparison.
+While c-LSFD achieves the highest performance for LP-MMSE, the gap between c-LSFD and d-LSFD is very significant.
+This behavior is consistent with the theoretical analysis in Fig.
+3, which shows that the SINR guarantee for d-LSFD becomes vacuous when the $\rho_{t}$ is large, as is typical for MMSE-based local combining due to strong global interference coupling.
+These results further validate that the proposed d-LSFD scheme is particularly effective for ZF-type combining schemes.
+Figure 7: The uplink sum SE comparison for various decoding schemes.
+Fig.
+7 illustrates the sum SE versus the number of pilots ($L_{p}$), comparing decoding schemes for G-PFZF and G-PWPFZF combining.
+For G-PFZF and G-PWPFZF combining, the performance gap between d-LSFD and c-LSFD is marginal (less than 3%), when $L_{p}=5$, as high pilot contamination makes c-LSFD better equipped to handle interference.
+However, as the number of pilots increases to moderate values, the gap reduces to negligible.
+This trend occurs because with sufficient orthogonal pilots, both approaches have similar capability to mitigate pilot contamination.
+Notably, even in the worst-case scenario with limited number of orthogonal pilots, the performance penalty remains modest, validating that the substantial fronthaul and computation reductions of our d-LSFD approach come with acceptable performance trade-offs across all pilot sequences.
+Furthermore, the proposed d-LSFD outperform the h-LSFD with a significant margin across all pilot densities for both the combining schemes.
+These results further show that the proposed d-LSFD retains near-centralized performance, for all pilot densities.
 Figure 8: The uplink sum SE comparison for various combining schemes.
@@
 8 depicts the sum SE versus the number of UEs ($T$) for various combining schemes.
-The proposed decentralized optimization framework delivers significant and scalable performance gains, with G-PFZF outperforming its baseline PFZF by 6.5% to 9.5% and G-PWPFZF surpassing PWPFZF by 5.5% to 10%, with the gaps widening as the number of UEs increases.
-Better performance of the proposed schemes highlights the critical limitation of the baseline PFZF and PWPFZF schemes, which rely on a hard, network-wide threshold using LSFCs to rigidly classify users as strong or weak at every AP.
-In contrast, the proposed framework introduces a decentralized optimization algorithm, executed independently by each AP, to determine a better pilot-based grouping strategy.
-This allows an AP to adaptively select none, a subset, or all of the pilots as strong, based on its local channel conditions and specific geometric configuration.
-This per-AP flexibility enables a more efficient and dynamic trade-off between utilizing degrees of freedom for interference suppression and desired signal enhancement.
+Compared with the baseline PFZF and PWPFZF schemes, when using either c-LSFD or d-LSFD, the proposed G-PFZF and G-PWPFZF with d-LSFD achieve gains of 6%-9.5% and 5%-10%, respectively, with the gaps widening as the number of UEs increases.
+This performance improvement comes from the adaptive per-AP pilot classification, which replaces the fixed threshold-based UE grouping of baselines.
+This enables a more efficient and dynamic trade-off between utilizing degrees of freedom for interference suppression and desired signal enhancement.
 Consequently, the proposed G-PWPFZF scheme achieves better performance over the G-PFZF scheme by suppressing interference of the weak UEs from the strong UEs at the APs.
+By comparing PFZF and PWPFZF under both c-LSFD and d-LSFD with their proposed generalized counterparts, the observed performance gains persist irrespective of the decoding strategy.
+This confirms that the improvements primarily stem from the proposed adaptive pilot-based combining enabled by local optimization.
 Figure 9: The uplink sum SE comparison for various combining schemes.
 Fig.
-5 illustrates the sum SE versus the number of orthogonal pilots $L_{p}$, demonstrating the improved performance of the proposed G-PFZF and G-PWPFZF schemes over their respective baseline counterparts.
-The proposed schemes achieve significant performance improvements of 6-14% for G-PFZF and 6-13% for G-PWPFZF compared to their respective baselines across different pilot lengths, with the performance gap widening as the number of pilots increases.
-This expanding gap highlights the advantage of the proposed adaptive user grouping strategy over the rigid network-wide threshold approach used in baseline schemes.
-The baseline schemes, must rigidly limit the number of strong pilots to avoid violating the fundamental condition for zero-forcing ($L_{\mathcal{S}_{m}}{>}A$), which leads to inefficient interference management.
-In contrast, our proposed decentralized framework empowers each AP to independently and optimally select its set of strong pilots, inherently ensuring that the selection is both locally optimal and always feasible ($L_{\mathcal{S}_{m}}{<}A$).
-This inherent adaptability allows our methods to leverage larger pilot sets more effectively, maximizing array gain and suppressing interference, which explains the observed widening of the performance gap with the number of pilots.
-Figure 6: The uplink sum SE comparison for various combining schemes.
-Fig.
-6 illustrates the sum SE versus the number of antennas per AP ($A$) for various combining schemes, revealing a characteristic performance pattern that highlights the adaptability of our proposed framework.
-The G-PFZF and G-PWPFZF schemes demonstrate consistent superiority across all antenna configurations, with the performance gap exhibiting a distinctive trajectory that reflects adaptive resource utilization.
-At the minimal antenna configuration (2 antennas), where spatial degrees of freedom $A-L_{\mathcal{S}_{m}}$ are severely constrained, both proposed and baseline schemes face fundamental challenges in interference suppression.
-Nevertheless, our adaptive grouping strategy achieves measurable gains of 2.5% for G-PFZF and 6% for G-PWPFZF through an adaptive selection of combining schemes, dynamically switching between the G-PFZF/G-PWPFZF, and the MR at each AP based on local conditions.
-In the mid-range antenna configuration (4-8 antennas), where careful trade-off between interference suppression and signal enhancement becomes crucial, our proposed method demonstrates performance gain of 10-14% for G-PFZF and 10-18% for G-PWPFZF, showcasing their superior grouping capabilities.
-As antenna numbers increase further (10-16 antennas), providing abundant spatial resources that benefit all schemes, the performance gap narrows to 4.5-6.5% for G-PFZF and 3.5-5.5% for G-PWPFZF, though our approach maintains consistent superiority.
-Figure 7: The uplink sum SE comparison for various combining schemes.
-Fig.
-7 illustrates the sum SE versus the number of UEs ($T$), comparing local and LSFD-based combining schemes.
-For our G-PFZF and G-PWPFZF schemes, the performance gap between local weights and optimal LSFD remains minimal at just 3% across all UE densities.
-This marginal performance difference is dramatically outweighed by the substantial system benefits: our local weight approach eliminates the need for complex network-wide coordination, reducing computational cost at the CPU by avoiding large-scale matrix inversions and reducing fronthaul overhead by eliminating the need to share channel state information across the network.
-Despite this streamlined architecture, our locally-weighted G-PFZF and G-PWPFZF schemes still outperform the baseline PFZF and PWPFZF schemes with LSFD by margins of 4.5-9% and 3.5-7.5% respectively.
-This demonstrates that the gains from our adaptive per-AP grouping strategy are so substantial that they outperform the benefits of optimal coordination applied to suboptimal threshold-based grouping.
-These results confirm that our fully decentralized approach achieves a better balance between performance and practicality.
-Figure 8: The uplink sum SE comparison for various combining schemes.
-Fig.
-8 illustrates the sum SE versus the number of pilots ($L_{p}$), comparing local and LSFD-based combining schemes.
-The performance gap between local and LSFD-based weights is less than 5.5% for sparse pilot configurations, where high pilot contamination makes centralized LSFD better equipped to handle interference.
-However, as the number of pilots increases to moderate and high values, the gap reduces to less than 1%, demonstrating near-identical performance.
-This trend occurs because with sufficient orthogonal pilots, both approaches have similar capability to mitigate pilot contamination.
-Notably, even in the worst-case scenario with limited pilots, the performance penalty remains modest, validating that the substantial fronthaul and complexity reductions of our decentralized decoding approach come with acceptable performance trade-offs across all pilot sequences.
-Furthermore, the proposed G-PFZF and G-PWPFZF schemes with local weights consistently outperform PFZF and PWPFZF with LSFD across moderate to high pilot lengths, confirming the advantage of adaptive per-AP combining strategies.
+9 illustrates the sum SE versus the number of orthogonal pilots $L_{p}$, demonstrating the improved performance of the proposed G-PFZF and G-PWPFZF with d-LSFD schemes over their respective baseline counterparts with c-LSFD and d-LSFD.
+The proposed schemes achieve significant performance improvements of 6-14% for G-PFZF and 5-13% for G-PWPFZF compared to their respective baselines across different pilot lengths, with the performance gap widening as the number of pilots increases.
+This expanding gap highlights the advantage of the proposed adaptive pilot grouping strategy over the rigid network-wide threshold approach used in baseline schemes.
+The baseline schemes must rigidly limit the number of strong pilots to avoid violating the fundamental condition for zero-forcing ($L_{\mathcal{S}_{m}}{>}A$), which leads to inefficient interference management.
+In contrast, our proposed decentralized optimization, inherently ensures $L_{\mathcal{S}_{m}}{<}A$, while maximizing array gain and suppressing interference.
 Figure 10: The uplink SE per user comparison for various combining schemes.
 Fig.
-9 presents the 90%-likely per-user SE across different combining schemes, demonstrating significant improvements in user fairness through our decentralized optimization framework.
-The proposed G-PFZF scheme achieves 45% improvement in 90%-likely SE compared to conventional PFZF, while G-PWPFZF shows a 34% gain over its baseline PWPFZF.
+10 presents the 90%-likely per-user SE across different combining schemes, demonstrating significant improvements in per user SE through our decentralized optimization framework.
+The proposed G-PFZF with d-LSFD scheme achieves 40-45% improvement in 90%-likely SE compared to conventional PFZF with both decoding scheme, while G-PWPFZF with d-LSFD shows a 28-34% gain over its baseline PWPFZF across both decoding scheme.
 These gains confirm that our adaptive per-AP grouping strategy not only enhances the overall system capacity but fundamentally transforms user experience across the network.
-The gains in 90%-likely SE particularly highlight our framework’s effectiveness in improving the performance of weak users.
+The gains in 90%-likely SE, particularly, highlight our optimization framework’s effectiveness in improving the performance of weak users.
 Figure 11: The uplink SE per user comparison for various decoding schemes.
 Fig.
-10 presents the 90%-likely per-user SE comparing local versus LSFD processing, revealing crucial insights about the performance-complexity trade-off in our decentralized framework.
-While our G-PFZF and G-PWPFZF schemes with local weights experience a modest performance reduction of approximately 10-10.7% compared to their LSFD-based counterparts, the 10% performance difference between local and LSFD processing represents a reasonable trade-off for achieving fully decentralized architecture, substantially reduced fronthaul overhead, and lower computational cost.
-Noticeably, our locally-weighted G-PFZF scheme outperforms the baseline PFZF with optimal LSFD by 32%, and G-PWPFZF with local weights surpasses the LSFD-enhanced PWPFZF baseline by 22.5%.
-These results demonstrate that the performance gains from our adaptive per-AP grouping strategy are so significant that they outweigh the benefits of optimal LSFD coordination applied to suboptimal threshold-based grouping.
+11 presents the 90%-likely per-user SE.
+While our proposed d-LSFD experience a modest performance reduction of approximately 5.5% and 7% compared to the c-LSFD for G-PFZF and G-PWPFZF scheme respectively.
+The less than 7% performance difference between d-LSFD and c-LSFD represents a reasonable trade-off for achieving fully decentralized architecture, substantially reduced fronthaul overhead, and lower computational cost.
+Noticeably, our proposed d-LSFD outperforms the h-LSFD by 100% and 65% for G-PFZF and G-PWPFZF combining respectively, thus outperforming the h-LSFD with significant margin.
+Overall, the results confirm that near-optimal decentralized decoding and adaptive combining address two orthogonal limitations of existing D-mMIMO systems and jointly enable scalable, high-performance uplink processing.
 V Conclusion and Future Work
-This paper has addressed the critical bottlenecks of sub-optimal performance, high fronthaul load, and computational cost in D-mMIMO networks by introducing a novel, decentralized decoding uplink architecture that fundamentally departs from conventional LSFD-based designs.
-Through the proposed local ZF framework, we have demonstrated that enabling each AP to independently determine its combining strategy via local optimization, classifying pilots as strong or weak without fixed thresholds and dynamically switching among PFZF/PWPFZF, FZF, or MR, yields significant performance improvements.
-The resulting generalized schemes, G-PFZF and G-PWPFZF, consistently outperform their fixed-threshold counterparts across all evaluated scenarios, achieving substantial gains in both sum and per user spectral efficiency.
-Moreover, the introduction of pilot-dependent combining vectors and interference-suppressing local weights applied distributively at each AP eliminates the need for centralized LSFD, thereby drastically reducing fronthaul overhead and computational cost.
-Remarkably, this decentralized approach incurs only a small performance penalty compared to idealized centralized coordination, while even outperforming conventional threshold-based schemes using optimal LSFD.
-These findings collectively establish that proposed decentralized decoding in the network, not centralized decoding, is the key to scalability and efficiency in future D-mMIMO systems without compromising on quality-of-service.
-In future, we plan to investigate the robustness of the proposed decentralized architecture and generalized combining schemes under more practical and challenging conditions, such as hardware impairments.
--A Derivation of closed-form SINR expression for G-PFZF combining
-In order to derive the closed-form expression, we need few expectations computation or expansions
-$\begin{split}\mathbb{E}\Big\{\Big|\!\!\sum_{m}\!\!\mathbf{v}_{mi}^{H}\mathbf{g}_{mi}\!\Big|^{2}\!\Big\}\!\!=&\!\sum_{m}\!\mathbb{E}\Big\{\!\Big|\mathbf{v}_{mi}^{H}\mathbf{g}_{mi}\!\Big|^{2}\!\Big\}+\!\Big|\!\sum_{m}\!\mathbb{E}\Big\{\mathbf{v}_{mi}^{H}\mathbf{g}_{m}\Big\}\Big|^{2}-\sum_{m}\Big|\mathbb{E}\Big\{\mathbf{v}_{mi}^{H}\mathbf{g}_{mi}\Big\}\Big|^{2},\end{split}$ (A1)
-$\mathbb{E}\Big\{\Big|\mathbf{v}_{mi_{t}}^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Big\}=$ $\beta_{mk}-\gamma_{mk}$ (A2)
-$\mathbb{E}\Big\{\Big|\mathbf{v}_{mi_{t}}^{H}{\mathbf{n}}_{m}\Big|^{2}\Big\}=$ $1.$ (A3)
-The desired signal term in (6) can be written as
+This paper establishes a principled framework for uplink processing in D-mMIMO systems by addressing both local combining and decentralized decoding.
+We propose two adaptive pilot-aware local combining schemes, G-PFZF and G-PWPFZF, where pilots are partitioned using local optimization.
+Furthermore, we establish a universal SINR lower bound that rigorously characterizes the performance loss incurred by decentralization of LSFD.
+A key insight from our theoretical analysis is that the effectiveness of decentralization hinges on the structure of the local combiner.
+While the proposed d-LSFD is not universally optimal for all combiners, we demonstrate that for the class of pilot-aware ZF-based scheme, it attains near-centralized performance.
+Numerical results confirmed that, for G-PFZF and G-PWPFZF combining, d-LSFD attains near-optimal performance.
+In contrast, for LP-MMSE combining, d-LSFD exhibits a larger performance gap due to the presence of strong global interference coupling.
+Overall, the proposed framework provides a scalable and theoretically grounded alternative to c-LSFD.
+In future, we plan to investigate the robustness of generalized combining schemes and the decentralized decoding under hardware impairments.
+-A Proof of Lemma 1
+On the domain where $F(x,q)$ is well defined, the denominator is strictly positive. Hence, minimizing $F(x,q)$ over $|x|^{2}\leq q$ is equivalent to maximizing
+$G(x)\triangleq\frac{q-|x|^{2}}{|1+x|^{2}}.$
+Write $x=r(\cos\theta+i\sin\theta)$ with $0\leq r\leq\sqrt{q}$. Then
+$G(r)=\frac{q-r^{2}}{1+2r\cos\theta+r^{2}}.$
+For fixed $r$, the $G(r)$ is minimized when $\cos\theta=-1$, yielding
+$G(r)=\frac{q-r^{2}}{(1-r)^{2}}.$
+For $0{\leq}q{<}1$, the stationary point r=qr{=}q lies in the admissible interval $0,\sqrt{q}]$ and gives $G(q){=}\frac{q}{1-q}$. Substituting into F=1/(1+G)F{=}1/(1+G) gives
+$\min_{|x|^{2}\leq q}F(x,q)=1-q.$
+For $q\geq 1$, $G(r)$ can be made arbitrarily large, and therefore
+$\inf_{|x|^{2}\leq q}F(x,q)=0.$
+-B Proof of Theorem 2
+$\text{We know,}\ \ \mathrm{SINR}_{t}(\mathbf{a}_{t}^{\mathrm{o}})=\frac{|(\mathbf{a}_{t}^{\mathrm{o}})^{H}\mathbf{d}_{t}|^{2}}{(\mathbf{a}_{t}^{\mathrm{o}})^{H}\mathbf{J}_{t}\mathbf{a}_{t}^{\mathrm{o}}}=(\mathbf{a}_{t}^{\mathrm{o}})^{H}\mathbf{J}_{t}\mathbf{a}_{t}^{\mathrm{o}}>0.$
+Define the error vector $\bm{\delta}_{t}\triangleq\mathbf{a}_{t}^{\mathrm{d}}-\mathbf{a}_{t}^{\mathrm{o}}$, so that $\mathbf{a}_{t}^{\mathrm{d}}=\mathbf{a}_{t}^{\mathrm{o}}+\bm{\delta}_{t}$ and also $\|\bm{\delta}_{t}\|_{2}=\rho_{t}\|\mathbf{a}_{t}^{\mathrm{o}}\|_{2}$ Using $\mathbf{d}_{t}=\mathbf{J}_{t}\mathbf{a}_{t}^{\mathrm{o}}$, define the normalized quantities $x\triangleq\frac{\bm{\delta}_{t}^{H}\mathbf{d}_{t}}{(\mathbf{a}_{t}^{\mathrm{o}})^{H}\mathbf{J}_{t}\mathbf{a}_{t}^{\mathrm{o}}}$, and $q\triangleq\frac{\bm{\delta}_{t}^{H}\mathbf{J}_{t}\bm{\delta}_{t}}{(\mathbf{a}_{t}^{\mathrm{o}})^{H}\mathbf{J}_{t}\mathbf{a}_{t}^{\mathrm{o}}}$.
+A direct expansion yields
+$\frac{\mathrm{SINR}_{t}(\mathbf{a}_{t}^{\mathrm{d}})}{\mathrm{SINR}_{t}(\mathbf{a}_{t}^{\mathrm{o}})}=\frac{|1+x|^{2}}{1+2\Re(x)+q}.$ (T1)
+By the Rayleigh quotient bounds,
+$q\leq\frac{\lambda_{\max}(\mathbf{J}_{t})\|\bm{\delta}_{t}\|_{2}^{2}}{\lambda_{\min}(\mathbf{J}_{t})\|\mathbf{a}_{t}^{\mathrm{o}}\|_{2}^{2}}=\kappa_{t}\rho_{t}^{2},$
+and by the Cauchy-Schwarz inequality in the $\mathbf{J}_{t}$-inner product, $|x|^{2}\leq q$. Using Lemma 1 for fixed $q$, we yields
+$\min_{|x|^{2}\leq q}\frac{|1+x|^{2}}{1+2\Re(x)+q}=\max\{0,\;1-q\}.$
+Substituting $q\leq\kappa_{t}\rho_{t}^{2}$ completes the proof.
+-C Proof of Proposition 1
+Consider an off-diagonal entry of $\mathbf{J}_{t}$ for $m\neq n$:
+$[\mathbf{J}_{t}]_{mn}$ $=\sum_{k=1,k\neq t}^{T}p_{k}^{u}\mathbb{E}\!\Big\{\mathbf{v}_{mt}^{H}\hat{\mathbf{g}_{mk}}\Big\}\mathbb{E}\!\Big\{\mathbf{v}_{nt}^{H}\hat{\mathbf{g}}_{nk}\Big\}.$
+The local MR and ZF combining vectors are linearly dependent only on the channel estimates of UE $t$ and the UEs sharing the same pilot as $t$, since pilot-sharing UEs have linearly dependent channel estimates as in (1).
+Hence, for any $k\neq t$ that does not share pilot with $t$, $\mathbb{E}\!\left\{\mathbf{v}_{mt}^{H}\hat{\mathbf{g}}_{mk}\right\}=0$, which implies $\mathbf{J}_{t}]_{mn}=0$ for all $m\neq n$.
+Therefore, $\mathbf{J}_{t}$ becomes a diagonal matrix.
+-D Closed-form SINR and d-LSFD expressions for G-PFZF combining
+Pilot-sharing UEs use $\mathbf{v}_{mi_{t}}^{\mathrm{LZF}}$ for strong pilots and $\mathbf{v}_{mi_{t}}^{\mathrm{MR}}$ for weak pilots. The desired signal term in (7) can be written as
 $|\text{DS}_{t}|^{2}=\Big|\mathbb{E}\Big\{\!\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{{g}}}_{mt}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{{g}}}_{mt}\Big\}\Big|^{2},$
-where $\mathcal{Z}_{i_{t}}$ and $\mathcal{Y}_{i_{t}}$ denote the subsets of APs where pilot $i_{t}$ is classified as strong and weak, respectively.
+where $\mathcal{Z}_{i_{t}}\subseteq\mathcal{M}_{t}$ and $\mathcal{Y}_{i_{t}}\subseteq\mathcal{M}_{t}$ denote the subsets of APs where pilot $i_{t}$ is classified as strong and weak, respectively.
 Using (12) and (14), the $|\text{DS}_{t}|^{2}$ can be written as:
-$\begin{split}|\text{DS}_{t}|^{2}&=\!\!\Big|\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mk}}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mk}}\Big|^{2}\\ &=\Big|\sum_{m=1}^{M}\!a_{mt}\sqrt{(A-\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2},\end{split}$ (A4)
+$\begin{split}|\text{DS}_{t}|^{2}&=\!\!\Big|\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mt}}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mt}}\Big|^{2}\\ &=\Big|\sum_{m=1}^{M}\!a_{mt}\sqrt{(A-\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mt}}\Big|^{2},\end{split}$ (A1)
 The interference term $\mathbb{E}\big\{|\text{UI}_{tk}|^{2}\big\}$ in (7) can be expanded as:
-$\begin{split}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mk}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}&=\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\!\Bigg\}\!+\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\!\Bigg\}\\ &+2\mathbb{E}\Bigg\{\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}.\end{split}$ (A5)
-Using (A1), the first term ($A_{1}$) of (A5) can be rewritten as:
-$\begin{split}A_{1}&=\mathbb{E}\Bigg\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!(\hat{\mathbf{g}}_{mk}+\tilde{\mathbf{g}}_{mk})\Big|^{2}\!\Bigg\}=\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}\!+\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}\\ &=\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\Big|\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\!\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}\\ &+\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|(a_{mt}\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}+\!\Big|\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2},\end{split}$
-Using (10), (11) and (A2), we can rewrite $A_{1}$ as
-$\begin{split}A_{1}&=\Big|\!\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (A6)
-Using (A1), the second term ($A_{2}$) of (A5) is expanded as:
-$\begin{split}A_{2}&=\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\Big|\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\!\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}\\ &+\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|(a_{mt}\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}+\!\Big|\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2},\end{split}$
-Using the (12), (13) and (A2), we can rewrite $A_{2}$ as
+$\begin{split}&\mathbb{E}\Big\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mk}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mk}\Big|^{2}\Big\}\\ &=\!\mathbb{E}\Big\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\Big\}\!+\mathbb{E}\Big\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\Big\}\\ &+2\mathbb{E}\Big\{\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nk}\!\Big\}.\end{split}$ (A2)
+The first term ($A_{1}$) of (A2) can be rewritten as:
+$\begin{split}&A_{1}=\mathbb{E}\Big\{\!\Big|\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!(\hat{\mathbf{g}}_{mk}+\tilde{\mathbf{g}}_{mk})\Big|^{2}\!\Big\}\\ &=\!\mathbb{E}\Big\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Bigg\}\\ &=\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\Big|\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\!\Big|^{2}\\ &-\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}+\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|(a_{mt}\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\\ &+\!\Big|\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2},\end{split}$
+Using (12) and (13), we can rewrite $A_{1}$ as
+$\begin{split}A_{1}&=\Big|\!\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}\\ &+\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (A3)
+The second term ($A_{2}$) of (A2) is expanded as:
+$\begin{split}&A_{2}=\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\!+\!\Big|\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\!\Big|^{2}\\ &-\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\hat{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}+\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\mathbb{E}\Big\{\!\Big|(a_{mt}\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\Big|^{2}\!\Big\}\\ &+\!\Big|\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2}-\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\Big|a_{mt}\mathbb{E}\Big\{\!(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\tilde{\mathbf{g}}_{mk}\!\Big\}\Big|^{2},\end{split}$
+Using the (14) and (15), we can rewrite $A_{2}$ as
 $\begin{split}A_{2}&=\Big|\!\!\!\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mk}}\Big|^{2}\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}\!\!+\!\!\sum\limits_{m\in\mathcal{Z}_{i_{t}}}\!\!\!|a_{mt}|^{2}\beta_{mk}.\end{split}$ (A4)
@@
 Using (A3), (A4) and (A5)
-$\begin{split}\mathbb{E}\Big\{\!|\text{UI}_{tk}|^{2}\!\Big\}&=\!\!\Big|\!\!\sum\limits_{m=1}^{M}\!\!\!a_{mt}\sqrt{(A{-}\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\!\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\!\right|^{2}+\sum\limits_{m=1}^{M}\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (A9)
-Using (A3), the term $\mathbb{E}\big\{|\text{GN}_{t}|^{2}\big\}$ of interference in (6) can be written as:
-$\begin{split}&\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!|a_{mt}|^{2}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Big\}+\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!\!a_{mt}|^{2}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Big\}=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (A10)
-Using the (A4), (A9) and (A10) in (6) gives the SINR closed-form expression in (14).
--B Proof of Lipschitz Continuity of $\nabla f(\bm{\delta}_{m})$
+$\begin{split}\mathbb{E}\Big\{\!|\text{UI}_{tk}|^{2}\!\Big\}&=\!\!\Big|\!\!\sum\limits_{m=1}^{M}\!\!\!a_{mt}\sqrt{(A{-}\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\!\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\!\right|^{2}\\ &+\sum\limits_{m=1}^{M}\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (A6)
+The term $\mathbb{E}\big\{|\text{GN}_{t}|^{2}\big\}$ of interference in (7) can be written as:
+$\begin{split}&\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!|a_{mt}|^{2}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Big\}+\!\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!\!a_{mt}|^{2}\mathbb{E}\Big\{\Big|(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{n}}_{m}\Big|^{2}\Big\}\\ &=\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (A7)
+Substituting (A1), (A6) and (A7) in (7) gives the (16).
+From Theorem 2, the d-LSFD weight for UE $t$ is $\mathbf{a}_{t}^{\mathrm{d}}{=}p^{u}_{t}\mathbf{D}_{t}^{-1}\mathbf{d}_{t}$.
+Since, $\mathbf{D}_{t}$ is a diagonal matrix, therefore, the d-LSFD for UE $t$ at AP $m$ can be written as $a_{mt}=p^{u}_{t}\frac{[\mathbf{d}_{t}]_{m}}{[\mathbf{D}_{t}]_{mm}}$.
+Here, the element of $\mathbf{d}_{t}$ for G-PFZF combining is given by
+$\mathbf{d}_{t}]_{m}=\mathbb{E}\{\mathbf{v}_{m{i_{t}}}^{H}\mathbf{g}_{mt}\}=\sqrt{(A-\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mt}}.$
+The diagonal element $\mathbf{D}_{t}]_{mm}$ is given by
+$\sum\limits_{k=1}^{T}p^{u}_{k}\mathbb{E}\{|\mathbf{v}_{m{i_{t}}}^{H}\mathbf{g}_{mk}|^{2}\}{-}p^{u}_{t}|\mathbb{E}\{\mathbf{v}_{m{i_{t}}}^{H}\mathbf{g}_{mt}\}|^{2}+\mathbb{E}\{|\mathbf{v}_{m{i_{t}}}^{H}\mathbf{n}_{m}|^{2}\}$
+$=\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\sum\limits_{\ \ \ \ \ k\in\mathcal{P}_{i_{t}}\backslash\{t\}}\!\!\!\!\!\!\!\!\!\!\!\!\!p^{{u}}_{k}\sqrt{\!(\!A{-}\delta_{mi_{t}}L_{\mathcal{S}_{m}}\!)\gamma_{mk}}{+}\!\sum\limits_{k=1}^{T}\!p^{{u}}_{k}\!(\beta_{{m}k}{-}\delta_{{m}i_{t}}\delta_{{m}i_{k}}\gamma_{{m}k}\!){+}1,$
+where $\mathbf{v}_{m{i_{t}}}=\mathbf{v}_{m{i_{t}}}^{\text{LZF}}$, if $i_{t}{\in}\mathcal{S}_{m}$, and $\mathbf{v}_{m{i_{t}}}=\mathbf{v}_{m{i_{t}}}^{\text{MR}}$, if $i_{t}{\notin}\mathcal{S}_{m}$.
+These expressions give the d-LSFD weights given in (18).
+-E Closed-form SINR and d-LSFD expressions for G-PWPFZF combining
+The derivation follows the same steps as in Appendix -D, the only difference is that instead of ${\mathbf{v}}^{\text{MR}}_{mi_{t}}$, we have ${\mathbf{v}}^{\text{PMR}}_{mi_{t}}$ combining vector.
+The desired signal term in (7)
+$|\text{DS}_{t}|^{2}=\Big|\mathbb{E}\Big\{\!\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{{g}}}_{mt}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{{g}}}_{mt}\Big\}\Big|^{2}.$
+Using (12) and (28), then
+$|\text{DS}_{t}|^{2}=p^{u}_{t}\Bigg|\sum_{m=1}^{M}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2},$ (B1)
+The first term in (7) can be written as
+$\begin{split}\mathbb{E}\Big\{\!|\text{UI}_{tk}|^{2}\!\Big\}&=\!\!\Big|\!\!\sum\limits_{m=1}^{M}\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\!\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}\\ &+\sum\limits_{m=1}^{M}\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (B2)
+The last term of interference in (7) can be written as:
+$\begin{split}\mathbb{E}\Big\{|\text{GN}_{t}|^{2}\Big\}=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (B3)
+Substituting (B1), (B2) and (B3) in (7) gives the (29).
+The d-LSFD for UE $t$ at AP $m$ can be written as $a_{mt}=p^{u}_{t}\frac{[\mathbf{d}_{t}]_{m}}{[\mathbf{D}_{t}]_{mm}}$, where
+$[\mathbf{d}_{t}]_{m}=\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mt}},$
+[…]
+Here, the weak pilot uses PMR combining vector insetad of MR combining vector. The above expression gives the d-LSFD weights in (30).
+-F Proof of Lipschitz Continuity of $\nabla f(\bm{\delta}_{m})$
 To establish the Lipschitz continuity of $\nabla f(\bm{\delta}_{m})$, we analyze its component functions and their boundedness over the compact domain $\bm{\delta}_{m}\in[0,1]^{L_{p}}$.
 The gradient component in (26) is $g(\delta_{mi})=\frac{\partial f(\bm{\delta}_{m})}{\partial\delta_{mi}}$ bounded as
-  • $0<S_{mt},I_{mt}\leq C_{1}$ (always bounded)
-  • $\frac{\partial S_{mt}}{\partial\delta_{mi}},\frac{\partial I_{mt}}{\partial\delta_{mi}}\leq C_{2}$ (bounded from (25) and (26))
+  • $0<S_{mt},I_{mt},\frac{\partial S_{mt}}{\partial\delta_{mi}},\frac{\partial I_{mt}}{\partial\delta_{mi}}\leq C_{1}$ (always bounded)
   • $\max(0,\delta_{mi}-\delta^{2}_{mi})\in[0,\frac{1}{4}]$ (bounded on $[0,1]$)
-  • $\max\left(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\right)\leq L_{p}$
-  • $\frac{I_{mt}}{I_{mt}+S_{mt}}\in(0,1),\qquad C_{3}\leq\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}\leq C_{4}$,
-where $C_{1}$, $C_{3}$, $C_{3}$ and $C_{4}$ are finite constants.
-The partial derivatives of $g(\delta_{mi})$ is given by
-$\begin{split}\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}&=\!\!\sum_{t=1}^{T}\!\!\Bigg[\frac{S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}-I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}}{\big(I_{mt}+S_{mt}\big)^{2}}\frac{I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}+\frac{I_{mt}}{I_{mt}+S_{mt}}\Bigg(\frac{I_{mt}\,\frac{\partial^{2}S_{mt}}{\partial\delta_{mi}^{2}}-S_{mt}\,\frac{\partial^{2}I_{mt}}{\partial\delta_{mi}^{2}}}{I_{mt}^{2}}\\ &-\frac{2\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big(I_{mt}\,\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\,\frac{\partial I_{mt}}{\partial\delta_{mi}}\big)}{I_{mt}^{3}}\Bigg)\Bigg]-\chi\Bigg[2\lambda_{1}\,\mathbf{1}_{\{\delta_{mi}-\delta_{mi}^{2}>0\}}\Big((1-2\delta_{mi})^{2}\\ &-2(\delta_{mi}-\delta_{mi}^{2})\Big)+2\lambda_{2}\,\mathbf{1}_{\{2-A>0\}}\Bigg].\end{split}$
+  • $\max\Big(0,\sum_{j=1}^{L_{p}}\delta_{mj}-A+1\Big)\leq L_{p}$
+  • $\frac{I_{mt}}{I_{mt}+S_{mt}}\in(0,1),\qquad C_{2}\leq\frac{I_{mt}\frac{\partial S_{mt}}{\partial\delta_{mi}}-S_{mt}\frac{\partial I_{mt}}{\partial\delta_{mi}}}{I_{mt}^{2}}\leq C_{3}$,
+where $C_{1}$, $C_{2}$ and $C_{3}$ are finite constants. The partial derivatives of $g(\delta_{mi})$ is given by
+$\begin{split}&\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}{=}\sum_{t=1}^{T}\Bigg[\frac{\big(S_{mt}\partial_{\delta_{mi}}I_{mt}{-}I_{mt}\partial_{\delta_{mi}}S_{mt}\big)\big(I_{mt}\partial_{\delta_{mi}}S_{mt}{-}S_{mt}\partial_{\delta_{mi}}I_{mt}\big)}{(I_{mt}{+}S_{mt})^{2}I_{mt}^{2}}+\frac{I_{mt}}{I_{mt}{+}S_{mt}}\Bigg(\frac{I_{mt}\partial^{2}_{\delta_{mi}}S_{mt}-S_{mt}\partial^{2}_{\delta_{mi}}I_{mt}}{I_{mt}^{2}}\\ &-\frac{2\,\partial_{\delta_{mi}}I_{mt}\big(I_{mt}\partial_{\delta_{mi}}S_{mt}{-}S_{mt}\partial_{\delta_{mi}}I_{mt}\big)}{I_{mt}^{3}}\Bigg)\Bigg]{-}\chi\Bigg[2\lambda_{1}\,\mathbf{1}_{\{\delta_{mi}{-}\delta_{mi}^{2}>0\}}\big((1{-}2\delta_{mi})^{2}{-}2(\delta_{mi}{-}\delta_{mi}^{2})\big){+}2\lambda_{2}\,\mathbf{1}_{\{2{-}A{>}0\}}\Bigg].\end{split}$ (C1)
 For all $\delta_{mi}\in[0,1]$, since all component functions $\frac{\partial g(\delta_{mi})}{\partial\delta_{mi}}$ are bounded and continuously differentiable on the compact domain $0,1]^{L_{p}}$, the Hessian matrix $\nabla^{2}f(\bm{\delta}_{m})=\nabla g(\bm{\delta}_{m})$ exists and is continuous.
@@
 This establishes the Lipschitz continuity of $\nabla f(\bm{\delta}_{m})$, ensuring convergence of the proximal gradient algorithm.
--C Derivation of closed-form SINR expression for G-PWPFZF combining
-The derivation of G-PWPFZF SINR closed-form expression follows the same step as the derivation of G-PFZF SINR closed-form expression, the only difference is that instead of ${\mathbf{v}}^{\text{MR}}_{mi_{t}}$, we have ${\mathbf{v}}^{\text{PMR}}_{mi_{t}}$ combining vector.
-The desired signal term in (6)
-$|\text{DS}_{t}|^{2}=\Big|\mathbb{E}\Big\{\!\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{{g}}}_{mt}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{PMR}})^{H}\hat{\mathbf{{g}}}_{mt}\Big\}\Big|^{2}.$
-Using (10) and (28), then
-$|\text{DS}_{t}|^{2}=p^{u}_{t}\Bigg|\sum_{m=1}^{M}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2},$ (C1)
-Following the steps in Appendix -A, the first term in (6) can be written as
-$\begin{split}\mathbb{E}\Big\{\!|\text{UI}_{tk}|^{2}\!\Big\}&=\!\!\Big|\!\!\sum\limits_{m=1}^{M}\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2}\!\!\left|\bm{\psi}^{H}_{i_{t}}\bm{\psi}_{i_{k}}\right|^{2}+\sum\limits_{m=1}^{M}\!|a_{mt}|^{2}(\beta_{mk}-\delta_{mi_{t}}\delta_{mi_{k}}\gamma_{mk}).\end{split}$ (C2)
-Similarly, the last term of interference in (6) can be written as:
-$\begin{split}\mathbb{E}\Big\{|\text{GN}_{t}|^{2}\Big\}=p^{u}_{k}\sum_{m=1}^{M}|a_{mt}|^{2}\end{split}$ (C3)
-Using the (C1), (C2) and (C3) in (6) gives the SINR closed-form expression in (30).